A Bibliography of Publications by, and about, I. J. Good

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

19 January 2026 Version 1.01

Abstract

This bibliography records publications by, and about, I. J. Good.

Title word cross-reference

#24636 [Goo81-30]. #6957 [Goo82q].

−2 log(likelihoodratio) [Goo77a]. 2 [Goo63d]. 2^{n} [Goo78l]. 2 × 2 [Mag81]. 2 × s [GL81, GL83]. ^{n} [Goo86v]. χ^{2} [GGM70, GGM71, GC78, GH83a, Goo90l]. g [Goo81w]. H_{N}/H_{S+N} [OG74a]. I(R_{T}: S_{T}) [OG74a]. J_{T} (N, S + N) [OG74a]. k [Goo77c, Goo80o]. n [Goo75h, Goo80d, Goo85s]. P [Goo87q, Goo93c, Goo80k, Goo83c, Goo86q, Goo89l, Goo89s, Goo90k, Goo90c, Goo92c]. R^{2} [AG81]. ρ^{|r|^{c}} [Goo80r]. s [GH82a, GH82b]. \sqrt{2} [GG68a, GG67, GG68b]. t [Goo86u, GS86b, Goo86c, LS86]. W_{T} [OG74a]. x! [Goo85f]. x [Goo85f]. X^{2} [Goo77a, Goo79w, Goo79x, Goo79i, HG86].

-Darwinism [Goo86v]. -fold [GH82a, GH82b]. -frames [Goo80o]. -Statistics [Goo77c]. -test [Goo86u, LS86]. -tuples [Goo85s]. -value [Goo80k, Goo83c, Goo90c]. -values [Goo86q, Goo89l, Goo89s, Goo90k, Goo92c, Goo93c].

11 [Goo99]. 13 [Goo01]. 142 [Mar90]. 15 [GM89]. 1950 [Ano53]. 1968 [ABH^{+}69]. 1970 [Goo70h].

2 [Goo63a]. 2003a [Goo01]. 22 [Mar90]. 228 [Goo89p]. 232 [Ano81]. 24 [Goo78e]. 29 [Goo74b]. 293 [Ano81].

3 [Goo89w]. 301 [CG96]. 305 [Ano90]. 307 [Ano90]. 31 [Goo93g]. 310 [Ano90]. 311 [Ano90]. 32 [CG77a]. 334 [Ano90]. 354 [Goo81-30]. 355 [Goo56b]. 37 [Goo89p, Goo89w]. 3d [Goo86k].

43 [FOO01]. 45 [Goo78a].

52 [Goo82q]. 58 [Goo81-30].

66 [Goo92g].

7 [McC96]. 75 [Goo89w]. 78 [Goo82q].

88f [GM89].

93k [Goo93g]. 96c [McC96]. 98e [FOO01]. 99g [Goo99].

A. [Bri92, Goo79-29]. Abelian [Goo51c, Goo52b]. Abramson [Goo64]. Abstract [Goo90a]. absurdum [Goo81b]. Acceleration [Goo61n]. acceptable [Goo87d]. Acceptance [Cox60, Doy94, GM84, Wel94]. accumulation [Goo03d]. accuracy [Goo85f, Mil75a, Mil75b]. achieve [Goo84l]. achievement [Goo84v]. ad [Goo81b]. Addenda [Goo89p]. Addendum [Goo60d, Goo86x, GM89, Goo82n, Goo86a, Goo89o]. additional [OG74b]. additions [FOO01, GD59]. Additive [PG86, GG71c, GH82a, GS86a]. Adenine [Goo78r]. adhockery [Goo83d]. adjustment [Goo79g]. Advanced [And74, AF75, BKM^{+}67, BvL68, Ber73, BK75, CGI78, Ehr74, EB75, ES76, ECG^{+}77, EP78, Gil74, GBMW75, GK75, GRM76, GSA75, Goo73c, Goo75i, GR78, GB80, Goo81-31, GC68b, HK75, KC75, KS68, MC75, MTD68, MR78, MW80, MM75, Oso78, PJR^{+}68, RC74, SB74, SH74, SAG68, SG68, VM68, WF68]. advice [Goo78g]. After [GM96, Goo90f]. again [Goo86g].

Against [CG96, Goo75a, Goo81i, Goo86s]. Agassi [Goo75d, JL95]. Age [SWT^{+}81, Goo86o]. Alan [vdV09]. alarm [Goo61o]. aléatoires [Goo66b]. Algorithm [Goo58e, Goo60d, Goo67a, Goo78q, Goo83l, GS86a, CG96]. Algorithms [CG96]. Allais [Goo56a]. Alleged [Goo78a]. allowed [Goo80d]. allowing [Goo85s]. also [Goo88a]. Alternating [Goo94o]. Alternatives [Goo72f, Goo81i, GT81]. always [Goo95d]. Amalgamation [GM87, Goo86h, GM89]. amazing [Goo72h]. Amendment [Goo76a]. Americaine [Goo56a]. Among [Goo72f]. Amount [Goo61a]. analogous [GS86a, Goo92f]. analogue [Goo61g, Goo82r, Goo81i]. Analogues [Goo62a]. analogy [And89]. analyses [Goo80a, Goo81s]. Analysis [Ano81, Goo58e, Goo60d, Goo61b, Goo83-33, Zel80, Goo78n, Goo80i, Goo80p, Goo83-31, Goo90p, Goo62e, SWT^{+}81]. analytic [Goo58g, Goo88i]. ’ancillarity [Goo87h, Goo84o]. Anecdotes [SWT^{+}81]. Ann [GM89]. Annenda [Goo81-29]. Anova [Goo86l]. Answer [Sme94]. Answers [RG69]. Anthologie [MA66]. Anthology [GMM62, MGM62]. anticipation [Goo00b]. Antisurprise [Goo83i]. any [Goo69b]. apology [Goo89r, Goo95b]. Appl [CG77a]. applicable [Goo82d]. Application [CG80, CG81, Goo76d, KG77, Goo79j, Goo85p]. Applications [BG76b, BG76a, DE86, Goo69e, Goo73b, Goo83-35, Goo86y, Goo86z, Hey86, Sin86, Sol86, Goo60c, Goo77b, Goo80w, Goo81-30, Goo84x, Goo09, Goo72b]. applied [Goo84p, HG85]. Appreciation [Fie08]. Approach [Goo84a, RN95, RN98, RN10, RNC^{+}21, Goo81w, Goo86h, Goo86l]. approaches [Goo83-28, JK89]. Approximate [Goo69f, Goo40, Goo72h]. Approximating [PG76]. approximation [Goo80a, Goo81s]. approximations [Goo86r, HG86]. April [ABH^{+}69]. arabinoside [Goo78r]. arc [Goo87n]. arc-sine [Goo87n]. area [Goo82p, Goo84g, Goo84m, Goo84n, Goo85c, GS86b]. Areas [PG76, OG74b]. Argument [Gil86, Goo87o, Goo90w, Goo79m, Goo87k, Goo89e, Goo89b, Goo90j, Goo94b]. Arguments [Goo62d, Goo59a]. Arnold [Goo75b, Goo76a]. arrays [GC77]. Article [Goo92g]. Artificial [Nor12, RN95, RN98, RN10, RNC^{+}21]. assertions [Goo89t]. assessment [Goo88d, Goo81m]. Associated [JG81]. Association [SC74, Goo81d, GL83, Goo85q]. Assuming [Goo75a, GLGH79]. Assumptions [Goo95h, Sto93]. astronomers [Goo98a]. Asymptotic [Goo61c]. atomic [Goo88d]. atoms [Goo98a]. attack [GHDB89]. attacks [Goo89j]. attribute [GT76]. authorship [Goo81l]. Autobiographic [Oyv52]. Autobiography [Oyv26]. autocorrelation [Goo80r]. avoiding [Goo78e]. Axiomatic [JK89]. Axiomes [Goo56a]. Axioms [Goo86-27].

B [FOO01, Goo74b, Goo78t, Goo84b, Goo84c, Goo84-27, Goo91c, Goo03c]. Background [Gol94]. Baked [GMM62, MGM62]. balanced [Goo87p, Goo91d]. ball [Goo81m, Goo88a]. Banquet [SWT^{+}81]. Barnard [Goo90t]. Barndorff [Goo83w]. Barnsley [Goo90b, Goo91c, Goo90r]. Based [Goo86b, Hal60, Goo41c, Goo86f]. Basis [BGHK94, Doy94, Gol94, Goo94m, Kyb94, She94, Sme94]. Bayes [Cox60, CG96, GC74, Goo75a, Goo79m, Goo81x, Goo83-30, Goo83-29, Goo83x, Goo84s, Goo85c, Goo86i, Goo86s, Goo86f, Goo88a, Goo92a, Goo94e, Goo00b]. Bayes/non [Goo83-30, Goo86i, Goo86f, GC74, Goo92a]. Bayes/non-Bayes [Goo83-30, Goo86i, Goo86f, GC74, Goo92a].

Bayesian [Goo75b, Goo76a, Goo65d, Goo67b, GHJT69, Goo74b, Goo76b, GG78, Goo78m, Goo79l, Goo80a, Goo80v, Goo80l, Goo80u, Goo81s, Goo82i, Goo82g, Goo83s, Goo84a, Goo84o, Goo84n, Goo86p, GC87, Goo89v, Wet60, Zel80]. Bayesianism [Goo78a, Goo86f]. Be [Goo74c, BH96, Goo60b, Goo78h, Goo81a, Goo81o, Goo81f, Goo84g, Goo88c, Goo93a, Goo96a, Kha78]. beer [Goo90l, Goo92c]. before [Goo81o]. Behrens [Goo86l]. Believing [Kyb94, Sme94, BGHK94, Doy94, Gol94, Goo94m, She94]. bending [And89]. Benoît [Goo78t, Goo84b, Goo84c, Goo84-27]. Berger [Jos96a]. Bernoulli [Goo70c, Goo83g]. Bessel [Goo90j]. Bessel-function [Goo90j]. best [Goo84k]. Between [Goo71e, Goo88h, Goo88j, OG74l, GG80a, Goo42, Goo80f, Goo80e, Goo81e, Goo82l, Goo83e, GL83, Goo85t, GL86, Goo87p, Goo89z, Goo90n, Goo91d, OG74a, OG74q, OG74r]. Bhattacharyya [GS85a]. bias [GS84]. Bienayme [GG90]. Biennial [SC74]. billiard [Goo79m, Goo88a]. billiard-table [Goo79m]. Binary [Goo63d, GG68a, Goo73b, GG67, GG68b, OG76]. binomial [Goo62g]. Biometrika [Goo92g]. Birkbeck [Sav62]. Birnbaum [Jos96a]. bisection [Goo85e]. bivariate [GHDB89]. Blackwell [Goo56b]. Bletchley [Goo79z, Goo79y, Goo80t, GMT^{+}12, Gre22, HS93, RDF12, Shi25, vdV09]. blurred [Goo95b]. Bocharova [FOO01]. Bode [Goo69f, Goo71a]. Book [Dav51, Goo51a, Goo51b, Goo52a, Goo55a, Goo56b, Goo56a, Goo57a, Goo58a, Goo61d, Goo62c, Goo62b, Goo63a, Goo64, Goo66a, Goo66b, Goo70a, Goo72c, Goo72b, Goo73a, Goo75b, Goo78b, Goo78c, Goo78d, Goo79a, Goo79b, Goo79c, Goo80b, Goo80c, Goo82a, Goo83a, Goo84b, Goo84c, Goo85a, Goo86b, Goo87a, Goo90b, Goo91a, GBS08, MA66]. Boolean [Goo81q, Goo79q]. both [Goo80d]. Bounded [GR48, Goo50a]. Box [Goo83-30].

Brainerd [WG82]. Brains [Goo65a]. Branching [Goo72b, Goo75f, Goo89m]. Breaking [GM96, GMT^{+}12, RDF12]. Brief [Goo92a, Goo95g, Goo85w]. Bruno [Goo69b]. Bump [GG80b, GG80c]. Bump-Hunting [GG80b, GG80c]. bumpy [Goo81i]. butterfly [Goo94d].

C [FOO01, Goo78a, Goo78e]. C10 [Goo78f]. C100 [Goo81a]. C101 [Goo81b]. C102 [Goo81c]. C103 [Goo81d]. C104 [Goo81e]. C105 [Goo81f]. C106 [Goo81g]. C107 [Goo81h]. C108 [Goo81i]. C109 [Goo81j]. C11 [Goo78g]. C110 [Goo81k]. C111 [Mag81]. C114 [Goo81l]. C115 [Goo81m]. C116 [Goo81n]. C117 [Goo81o]. C118 [AG81]. C119 [Goo81p]. C12 [Goo78h]. C121 [Goo82b]. C122 [Goo82c]. C123 [Goo82d]. C124 [Goo82e]. C126 [Goo82f]. C127 [Goo82g]. C129 [Goo82h]. C130 [Goo82i]. C131 [GJ82]. C133 [Goo82j]. C134 [Goo82k]. C135 [Goo82l]. C136 [Goo82m]. C138 [Goo82n]. C139 [Goo82o]. C14 [Goo78i]. C140 [Goo82p]. C141 [GH82a]. C142 [GH82b]. C143 [Goo83b]. C144 [Goo83c]. C145 [Goo83d]. C146 [Goo83e]. C147 [GL83]. C148 [Goo83f]. C149 [Goo83g]. C15 [Goo78j]. C150 [Goo83h]. C151 [Goo83i]. C152 [Goo83j]. C154 [Goo83k]. C155 [GS83]. C157 [Goo83l]. C158 [Goo83m]. C16 [Goo78k]. C160 [Goo83n]. C161 [Goo83o]. C162 [Goo83p]. C163 [Goo83q]. C164 [Goo83r]. C165 [Goo83s]. C166 [Goo83t]. C167 [Goo83u]. C168 [Goo83v, Goo89n]. C169 [Goo83w]. C17 [Goo78l]. C170 [Goo83x]. C171 [Goo83y]. C172 [Goo83z, Goo89o]. C173 [Goo83-27]. C174 [Goo83-31, Goo83-29, Goo83-28]. C175 [Goo83-29]. C176 [Goo83-30]. C177 [Goo83-31]. C178 [Goo83-32]. C179 [GH83a]. C18 [Goo78m]. C180 [GH83b]. C184 [GS84]. C185 [GL84]. C186 [Goo84d]. C187 [Goo84e]. C189 [Goo84f].

C19 [GG78]. C191 [Goo84g]. C192 [Goo84h]. C193 [Goo84i]. C194 [Goo84j]. C195 [Goo84j]. C197 [Goo84k]. C198 [Goo84l]. C199 [Goo84m]. C2 [Goo77a]. C200 [Goo84n]. C201 [Goo84o]. C202 [Ano90, Goo84p]. C203 [Goo84q]. C204 [Goo84r]. C205 [Goo84s]. C208 [Goo84t]. C209 [Goo84u]. C210 [Goo84v]. C212 [GBSS84]. C213 [Goo84w]. C214 [Goo84x]. C215 [Goo84y]. C216 [Goo85b]. C217 [Goo85c]. C218 [Goo85d]. C22 [Goo78n]. C220 [Goo85e]. C221 [GT85]. C223 [Goo85f]. C224 [Goo85g]. C225 [Goo85h]. C226 [Goo85i]. C227 [Goo85j]. C228 [Goo85k]. C229 [Goo85l]. C23 [Goo78o]. C232 [Goo85m]. C233 [Goo85n]. C234 [Goo85o]. C235 [Goo85p]. C236 [Goo85q]. C238 [Goo85r]. C240 [GS85b, Mar90]. C241 [Goo85s]. C242 [HG85]. C244 [Goo85t, Goo86d]. C246 [Goo86c]. C247 [GS86b]. C248 [Goo86d]. C25 [Goo79d]. C250 [Goo86e]. C257 [Goo86u, LS86]. C258 [Goo86u]. C26 [Goo79e]. C260 [Goo86f]. C261 [Goo86g]. C264 [Goo86h]. C265 [Goo86i]. C266 [Goo86j]. C267 [Goo86k]. C268 [Goo86l]. C269 [Goo86m]. C270 [Goo86n]. C271 [Goo86o]. C272 [Goo86p]. C273 [Goo86q]. C274 [Goo86r]. C275 [Goo86s]. C276 [Goo86t]. C277 [Goo87b]. C278 [Goo87c]. C28 [Goo79f]. C280 [GS87]. C283 [Goo87d]. C284 [Goo87e]. C285 [Goo87f]. C286 [Goo87g]. C288 [Goo87h]. C289 [Goo87i]. C29 [Goo79g]. C294 [Goo87j]. C297 [Goo88a]. C299 [Goo88b]. C300 [Goo88c]. C302 [Goo88d]. C303 [Goo88e]. C304 [GL88]. C305 [Goo88f]. C307 [And89]. C310 [JK89]. C311 [Goo89a]. C312 [Goo89b]. C313 [Goo89c]. C314 [Goo89d]. C315 [Goo89e]. C316 [Goo89f]. C317 [Goo89g]. C318 [Goo89h]. C319 [Goo89i]. C320 [Goo89j]. C321 [Goo89k]. C322 [Goo89l]. C323 [Goo89m]. C324 [Goo89n]. C325 [Goo89o]. C327 [Goo89p]. C328 [GHDB89].

C329 [GF89]. C330 [Goo89q]. C331 [Goo89r]. C332 [Goo89s]. C333 [Goo89t]. C334 [Goo89u]. C335 [Goo89v]. C340 [GC90]. C342 [Goo90c, Joh90]. C342.5 [Goo90c]. C344 [Goo90d]. C345 [GJH90]. C346 [GG90]. C347 [Goo90e]. C350 [Ano90]. C356 [Goo90f]. C357 [Goo90g]. C359 [Goo90h]. C361 [Goo90i]. C363 [Goo90j]. C367 [Goo90k]. C368 [Goo90l]. C369 [Goo90m]. C370 [Goo90n]. C371 [Mar90]. C372 [Goo90o]. C377 [Goo90p]. C378 [Goo90q]. C38 [Goo79p, Goo79h]. C383 [Goo91b]. C386 [Goo92b]. C387 [Goo92c]. C388 [Goo92d]. C389 [Goo92e]. C390 [Goo92f]. C394 [Goo93a]. C397 [Goo93b]. C398 [Goo93c]. C399 [Goo93d]. C4 [GG77]. C40 [Goo79i]. C400 [Goo93e]. C407 [Goo94a]. C41 [Goo79k, Goo79j]. C415 [Goo94b]. C416 [Goo94c]. C417 [Goo94d]. C418 [Goo94e]. C419 [Goo94f]. C42 [Goo79k]. C420 [Goo94g]. C421 [Goo94h]. C423 [Goo94i]. C424 [Goo94j]. C426 [Goo94k]. C43 [Goo79l]. C435 [Goo95a]. C436 [Goo95b]. C437 [Goo95c]. C438 [Goo95d]. C439 [Goo95e]. C44 [Goo79m]. C443 [Goo96a]. C444 [BH96, Goo96b]. C445 [Goo96b]. C446 [Goo98a]. C45 [Goo79n]. C46 [Goo79o]. C47 [Goo79p]. C473 [Goo03a]. C49 [Goo79q, Goo80h, Goo81q]. C5 [Goo78p]. C50 [Goo79r]. C52 [Goo79s]. C53 [Goo79t]. C54 [Goo79u]. C56 [Goo79v]. C57 [Goo79w]. C58 [Goo79x]. C59 [Goo80j, Goo80d]. C6 [Goo78q]. C62 [GG80a]. C64 [Goo80e]. C67 [Goo80f]. C68 [Goo80g]. C69 [MG80]. C7 [Goo78r]. C70 [Goo80h]. C71 [Goo80i]. C72 [Goo80j]. C73 [Goo80k]. C74 [Goo80l, Goo83-30]. C77 [Goo80m]. C78 [Goo80n]. C79 [Goo80o]. C80 [Goo80p]. C81 [Goo80q]. C84 [Goo81q]. C87 [Goo81r]. C88 [Goo81s]. C89 [Goo81t]. C9 [Goo78s]. C90 [Goo80r]. C91 [Goo81u]. C92 [Goo81v]. C93 [GT81]. C94 [Goo81w]. C95 [Goo81x]. C96 [Goo81y]. C97 [Goo81z]. C98 [Goo81-27]. C99 [Goo81-28]. Calculating [PG83, Goo60a, Goo61h, Goo51a, Goo51b]. calculation [Goo81p, Goo82n, Goo83l, Goo87d]. Calculus [Goo61e, Goo61f, Goo80q]. Cambridge [Goo82q, Goo89p, Goo89w]. Can [Goo58d, Goo74c, Goo60b, Goo81a, Goo96a]. Candidates [Goo75h, Goo80d]. Capsule [SWT^{+}81]. carbon [Goo98a]. Carlitz [Goo76e]. Carlo [Goo60e, Goo81x, Goo81o, Goo81j, Goo86q]. carpet [Goo76b]. cascade [Goo49, Goo55d]. case [Goo89u]. cases [Goo81f, GS84]. categorical [Goo83f]. Categorized [CG70, CG71]. Causal [Goo61e, Goo61f, Goo94l, Goo95h, Sto93, Goo80q, Goo94b]. Causality [Goo59d, Goo61o, Goo71d]. causation [Goo80i, Goo85d, Goo93e]. causes [Goo93a]. cells [Goo81p, Goo82n]. centroid [GG70, GG71a]. Centrosymmetric [Goo70d]. Century [MHR80]. certain [Goo42, OG74q]. Chain [Goo61k, Goo63d, Goo97a]. Chains [Goo55e, Goo57c, Goo73b, Goo82m]. Chance [Goo78t, Goo57a]. Chang [FOO01]. channel [Goo89c]. Chaology [Goo89y]. Chaotic [Goo61n]. Chaotics [Goo89d, Goo89y]. Chapman [Goo56b]. Characteristic [Goo61n, Goo68b, Goo68a, Goo81j, GJ82]. Charles [Goo72b]. Chebyshev [Goo61g]. Checks [Goo67a, Goo87b]. chess [Goo55g]. chess-players [Goo55g]. Chi [Goo61l, Goo69a, Goo70b, Goo85u, Kha78, GG80a, Goo78s, Goo81i, Goo81p, Goo81j, Goo81n, Goo82r, Goo82n, Goo83l, GS84, HG86, Mag81]. Chi-Square [Goo85u]. Chi-Squared [Goo61l, Goo69a, Goo70b, Kha78, Goo78s, Goo81i, Goo81p, Goo81j, Goo81n, Goo82r, Goo82n, Goo83l, GS84, HG86, Mag81]. chi-squareds [GG80a]. Choice [Goo57a, Goo72f, GM72, Goo79h, Goo83n, GL86]. Choix [Goo56a]. choosing [GL86]. chronic [Aga75a]. Ciphers [GMT^{+}12, RDF12]. Circulant [Goo50b]. Circulants [Goo86w, Goo86x, Goo86a]. classes [Goo78l, Goo84x]. classical [Goo41b]. Classification [Goo55b, Goo62d, Goo59a]. Claude [SW93]. clinical [Goo81g]. closely [vdV09]. clustering [Goo79s, Goo86c]. clusters [Goo82h]. codebreaker [Ano09f, vdV09]. Codebreakers [Gre22, HS93, Smi19]. Codes [GM96, FOO01]. Coding [Goo64, Goo62b]. coefficient [Goo89f]. coefficients [Goo83m, Goo92f]. colleague [Goo61g]. Collected [Anoxx, Bri92, SW93]. collecting [Goo81t]. Collection [MHR80]. collective [GT76]. Colleges [Sav62]. Colossus [WG82]. columns [Goo81d]. Combination [Goo55h, Goo79j, Goo89a]. combinatorial [Goo83y, Goo85k]. combining [Goo84w, JK89]. coming [Nor12]. Comment [Aga75b, Bar90, DE86, Goo71a, Goo72f, Goo82s, Goo82u, Goo84z, Goo85v, Goo86-27, Goo87q, Hey86, Ros75, Sin86, Sol86, Goo79g, Goo79p, Goo79d, Goo81q, Goo83n, Goo84j, Goo85j, Goo86q, GG90, Goo90c, Goo90t, Goo91b, Goo94a, Goo94i, Goo95a, Goo96e, Goo79k, Goo80h, Goo80j, Goo86d]. Commentary [Isr94]. Comments [Doy94, FOO01, Goo75c, Goo75d, Goo75e, Goo86u, GM96, Goo98b, Goo02, Goo04, Goo06a, Goo06b, GBS08, Har75, Wet75, Goo81h, Goo90m, Goo92c, Goo93b, Goo90q]. commonsense [Goo84f]. communication [Goo89c]. companion [Goo61g]. compared [Goo83-28, HG86]. comparing [Goo86p]. Comparison [OG74a, Goo60a, Goo61h, Goo79e, GS86a]. comparisons [Goo84f, Goo84g, Goo84i]. compelling [Goo89i]. Compenis [GG77]. competitors [Goo80g]. compilers [Goo68h]. complete [Goo79-27]. Complex [Goo93f, Goo93g]. Complexity [Goo74a, Goo78g]. component [GS85b]. Components [CG70, CG71]. Comportant [Goo56a]. Compound [Hal60, LR77]. compounds [Goo98a]. Compromise [GC74, Goo92a, Goo83-30, Goo86i, Goo90n, Wet75]. compromises [Goo86f]. Computation [CG96, Goo71c, Goo81x, Goo81j]. Computer [Lee95, Anoxx, Goo60a, Goo61h, Goo84l, Goo84-28, Goo86i]. computer-intensive [Goo86i]. Computers [Goo79y, Goo80t, Goo79z]. Computing [MHR80, Goo87f]. concentration [Goo87c]. Concept [Goo82s, Goo41c]. Concerning [Goo74a, Goo85v, Goo90a, Goo55c, GD58, GD59, Goo65e, Goo79j, Goo80h, Goo80j, Goo81c, Goo83s, GH83b, Goo84z, Goo85d, Goo85j, Goo86q, Goo87h, Goo90q, Goo90i, GG90, Goo90u, Goo91b, Goo92c, Goo93b, Goo94i, Goo95a]. Conclusions [Goo06b, GBS08, OG74b, Goo86u, Goo98b, Goo02, Goo04, Goo06a]. Conditions [Goo69a, Goo70b, Kha78]. Conference [SWT^{+}81]. Confidence [Goo87e]. Confirmation [Goo60f, Goo61m, Goo85v, Goo89h]. congruential [GG71c]. Conjecture [Goo70e, Goo83k, Goo90i, Goo94i, LR77]. conjectured [Goo84z, Goo87b]. Conjectures [Goo55c, Goo06b, GBS08, Goo83r, Goo86u, Goo98b, Goo02, Goo04, Goo06a, GH83b]. connected [Goo81k]. connection [Goo00b]. connections [Ked12]. consequence [Goo83g]. consistency [Goo79f, Goo87b, Goo94n, Goo97b, Goo97c, Goo98c, Goo99, Goo03b, Goo03c, McC96, Joh90]. constant [GM72]. Contingency [CG80, CG81, CG96, Goo56c, Goo63b, Goo76d, GM87, GC87, PG86, CG82, GG80a, GC77, Goo79e, Goo79x, Goo79i, Goo80a, Goo81s, Goo81x, Goo81w, GL81, Goo83f, Goo83-34, Goo85q, Goo86s, GM89, GF89].

Continued [Goo93f, Goo41a, Goo58g, Goo83v, Goo83u, Goo83z, Goo89p, Goo89n, Goo89w, Goo93g]. continuity [Mag81]. Continuous [Goo61k]. continuum [GS85b]. contour [Goo61b, Goo62e]. contradict [Goo81b]. Contribution [Goo94m]. contributions [Goo80v]. conversation [Ban96]. convolutional [FOO01]. correct [Goo88f]. Correcting [Per00]. Correction [Goo68c, Goo74a, Goo83s, Goo85d, Goo87h, Goo00a, Goo01, Mag81, Mar90]. Corrections [Ano81, CG81, Goo57b, Goo61i, Goo61j, Goo70b, GD59]. Correlation [Goo72d, Goo72a, And89, Goo78n, Goo79t, Goo87p, Goo91d, Goo92e, Goo92f, Goo83e, Goo87g]. correlations [Goo81e]. Correspondence [Goo68d]. Corrigenda [CG71, Goo52b, Goo57c, Goo62e, Goo62f, Goo66c, Goo66e, GGM71, GC78, Goo82q, Goo89w]. Corrigendum [GG68a, Goo68e, Goo74b]. Corroborated [Goo82t]. Corroboration [Goo60g, Goo68f, Goo68e, Goo75k, Goo89g]. cost [Goo81t]. Costs [Hal60]. cotangents [Goo93d]. Could [Goo58b, Goo58c]. Count [Goo61k]. Counter [PBB^{+}71, Goo87n]. Counter-examples [PBB^{+}71]. counter-intuitive [Goo87n]. course [Goo95d]. covariance [GG80a, Goo60a, Goo61h, Goo63e, GS85a, OG74j]. Coverage [GT56, Goo82o]. covert [Goo86f]. Cox [BLH^{+}60]. credibility [Goo90n]. criminals [Goo78o]. criteria [GL81]. Critical [JL95]. Critique [Goo56a]. cryptanalysis [Goo00b]. cubes [GT78]. cumulants [Goo71b, Goo72e, Goo75g, Goo79v, Goo82r, Goo82q, GH83b, Goo83r, Goo83y, Goo83p]. cumulative [Goo61b, Goo62e]. curiosity [Goo85k]. current [Goo83-30]. Cuthbert [Goo79g].

D [FOO01, Goo51a, Goo51b, Goo56b, RDF12].

D. [BG70]. D.-J [FOO01]. Dangerous [Goo74c]. Daniel [Goo79g, Goo83g]. Darwinism [Goo86v]. Data [GG80b, GG80c, Goo83-33, Goo82f, Goo83f, Goo90p]. David [Aga75b, Goo75c, Ros75]. Debate [GM96, Per00]. deciding [Goo61a]. decimals [Goo46a]. Deciphered [Ano81, Rej81]. decision [FOO01]. decisionary [Goo61a]. Decisions [Goo52d, Goo56b, Goo63a]. decoding [FOO01]. Decomposition [Goo69e, GS86a]. decompositions [Goo80o]. defeat [Goo95c]. Defense [Gil86, Goo85o]. Degrees [Goo73d, Goo85d, Goo80i]. demarcation [Goo89g]. DeMorgan [BG76b]. Densities [GG71b, GG72, Goo78j, Goo82e]. Density [GG80b, GG80c, Goo82a, ST08, Goo79u, Goo81r, Goo82f, Goo84q, Goo85l, GHDB89, Goo95e, Goo83h]. dependence [Goo85q]. depends [Goo85r, Goo95d]. Derivation [Goo66d]. derivations [Goo00a, Goo01]. Derived [Cox60, Goo90j]. descriptions [Goo81m]. design [Goo84f]. designs [Goo87p, Goo90h, Goo91d]. details [GHDB89]. Detection [OG74l, Goo60b]. detectors [Goo90e]. determinant [Goo84z]. determinants [Goo47, Goo80s, Goo81u, Goo81z, GL83]. determination [Goo81l, Goo83s, Goo89f]. Determining [GJ82]. Devaney [Goo91c]. device [Goo83-31]. diagnosing [Goo89c]. Diagnosis [GHDB89]. Dickey [Goo84u]. Dietmar [Goo91a, Goo91c]. Differences [Goo61c]. differential [Goo92f]. digits [Goo88f]. dimension [Goo41c, Goo78t]. dimensional [Goo41a, Goo89p, Goo89w]. dimensionality [Goo90f]. dimensions [Goo83v, Goo86s, Goo89n]. diminishing [Goo80k, Goo83c]. Diophantine [Goo84p]. direction [Goo63e]. Dirichlet [CG81, CG80, Goo52c, Goo75a, Goo76d, Goo81w, Goo83-29, Goo84d, Goo86r, GC87]. dirty [Goo84q]. Disclaimer [GM96]. discovered [Goo85h, Goo86m]. discrepancy [GL85]. Discrete [CG77a, CG77b, Goo79-28, Goo83c]. discriminant [Goo82e]. discriminiation [Goo79f]. Discussion [ABH^{+}69, BLH^{+}60, CMU^{+}53, CMU^{+}93, Goo69b, Goo85u, Goo87k, Goo94m, SBC^{+}62, Sav62, She94, Goo67b, GL81, Goo90v, Goo19]. disease [Aga75a]. distance [Goo90m]. distances [Goo85c]. distinct [Goo80m]. Distributed [Kha78]. Distribution [Goo56e, Goo57b, Goo57h, Goo61i, Goo61l, Goo69a, Goo70b, GGM70, GC74, Hal60, KG77, GG77, Goo55d, Goo69g, Goo78p, Goo79j, Goo80e, Goo81p, Goo81-27, Goo82n, GH83b, Goo83r, GH83a, Goo83y, Goo83l, Goo84d, Goo85s, Goo85t, GS86b, Goo86c, Goo90l, GC90, Goo96a, GGM71, GC78]. Distributions [CG80, CG81, Goo67b, GGM70, GGM71, Goo75f, Goo76d, GC78, Hal60, JG81, JG83, Goo61j, Goo74b, Goo78m, Goo80m, Goo81e, Goo83-27, Goo87f]. Divergence [OG74c, Goo84s, OG74f]. divergent [Goo41b, Goo90d]. Diversity [Goo82s, GS83]. Does [Goo71a, Goo81b, Goo77a]. dogs [Goo98a]. dragged [Goo78h]. dropped [Goo81f]. Dual [Goo95e]. Duke [Ked12]. d’une [Goo56a]. duplicated [Goo83j]. dynamic [Goo87j, Goo81b]. Dyson [Goo70e].

E1915 [TLM^{+}66, TQ68]. E1916 [LB68]. E1917 [MH70]. E1918 [WTR68]. E1919 [RF68]. E1920 [CB68]. E1921 [CGGR68]. E1922 [WV68]. E1923 [GT68]. E1924 [HG68, TLM^{+}66]. E2884 [HSO^{+}81, HS82]. E2885 [SS84]. E2886 [OK84]. E2887 [AC84]. E2888 [MS82]. E2889 [GB82, HSO^{+}81]. earlier [Goo90k]. earliest [Goo86o]. Early [Goo79z, Goo79y, GS83, Goo84r, Goo84t, Goo88d]. East [SC74]. École [Goo56a]. Econometrics [Zel80, Goo75b, Goo76a]. economic [Goo87c]. Eddington [Goo98a]. Edition [RDF12]. Editor [BRG^{+}91, WTA^{+}90, BDR^{+}72, BBC^{+}89, CTL^{+}77, DTB^{+}70, DMK^{+}70, DGK^{+}75, GGF^{+}01, GFG^{+}65, Goo65c, Goo68h, GBC69, GFM^{+}77, GN85, HPG^{+}71, HLT^{+}87, JEE^{+}83, KBE^{+}73, KsLG^{+}74, KLC^{+}88, LFB^{+}72, LOG^{+}69, PCM^{+}74, SKR^{+}71, VKT^{+}69, WSG^{+}73]. editorial [Goo90c, Goo84j, Goo86u, Goo90i]. eds [Goo91a]. effect [Goo81d, Goo84v, Goo86k, Goo94d]. Effective [Goo60b]. effort [Goo61a]. Einstein [Goo00a, Goo01]. electron [Goo89v]. electronic [Goo60a, Goo61h]. Electronics [SWT^{+}81]. Elementary [AC84, CB68, CGGR68, GT68, GB82, HSO^{+}81, HS82, HG68, LB68, MH70, MS82, OK84, RF68, SS84, TLM^{+}66, TQ68, WTR68, WV68, GJH90]. Ellipsoidal [JG83]. Elwood [SW93]. Emergence [McS70, Goo70h]. Emphasis [GMT^{+}12, RDF12]. empirical [Goo00b]. English [Goo55b]. ENIAC [WG82]. Enigma [Ano81, Goo00b, Goo93h, Rej81]. Entropy [Goo63b, OG74d, Goo61d]. enumeration [Goo65b, Goo68c, GC77]. enveloping [Goo89x]. enzyme [GHDB89]. equal [OG74j]. Equations [Goo57f, Goo84p]. Equiprobability [Goo75a]. Equiprobable [GGM70, GGM71, GC78, GG77, GH83a, GS84, HG86]. Errata [CG77a, Goo62f, Goo72e, Goo81-29, Ked12, Ano90]. Erratum [Goo81-30, Goo93g, Goo99, Goo01, Goo89o, Goo89n]. error [Goo79q, Goo81q, Goo81c, Goo82o, Goo84u, Goo87d]. errors [Goo80f, Goo83k, GT85, Goo90o]. Especially [Goo63b, GS85a]. Essay [Goo65d, GHJT69]. Essays [Goo99, JL95, MHR80, Zel80, Goo01, McC96]. estimate [Goo84e]. estimated [Goo82o]. estimates [Goo78f, Goo79e]. Estimating [PG86, Goo63e, Goo78m, Goo78j, Goo81-27]. Estimation [Goo53a, Goo56c, Goo65d, GLGH79, GG80b, GG80c, ST08, GHJT69, GG72, Goo79l, Goo79j, Goo79r, Goo81r, GL81, Goo82f, Goo82g, Goo83f, Goo84q, Goo85l, GL85, GL86, Goo86q, Goo87i, GL88, GHDB89, Goo82a]. etc [Goo83-29]. Ethical [Goo78k]. Euler [Goo70c]. Eulerian [GT78]. Evaluation [Goo69f, Goo89v, Goo94a]. evaluations [Goo89q]. even [Goo98a]. Event [Goo82t, Goo81y, Goo81-32]. events [OG74k]. Everywhere [Goo90r, Goo90b]. Evidence [BGHK94, Doy94, Gol94, Goo50d, Goo60g, Goo67d, Goo68e, Goo87q, Goo94m, Kyb94, OG74l, She94, Sme94, Goo61o, Goo71d, Goo80f, Goo81c, Goo81y, Goo81-32, Goo83p, Goo83t, Goo84k, Goo85m, Goo85w, Goo89a, Goo89i, Goo89b, Goo03d, JK89, OG74i, OG74f, OG74h, OG74k, OG74p, Dav51]. Evolving [Goo68f]. ’exact [Goo84r, DG57, GGM70, Goo81p, Goo82n, GH83a, Goo83l, Goo85s, Goo90l, GGM71, GC78]. examinations [Goo79h, Goo83n]. example [GS86a, Goo87j, Goo89c]. examples [PBB^{+}71]. exceed [Goo77a]. Excerpts [SWT^{+}81]. excess [Goo98a]. excluded [Goo85o]. Exemplified [GG80b, GG80c]. existence [Goo94g]. Expansion [GG68a, Goo76e, Goo60c, Goo65b, GG67, Goo68c, GG68b]. Expected [OG74e, OG74f, OG74g, OG74h, Goo83o, OG74i, OG74o, OG74p]. experiment [Goo86o]. Experiments [Goo60g, Goo68e]. Explanation [Goo68f, Goo82t, Goo71d, Goo77b, Goo80w, Goo81-30]. Explanatory [Goo60g, Goo68e, Goo85r]. Explication [Goo66d, Goo71d, Goo94b]. Explicativity [Goo75k, Goo77b, Goo80w, GM84, Goo81-30]. explicatum [Goo84k, Goo89b]. Exploratory [Goo83-33, Goo90p]. Explosion [Yud13]. exponential [Goo83q]. exposition [Goo87g]. Expressing [OG74i]. Expressions [Goo63c, Goo66e, Goo66c, Goo71b, Goo72e, GJ82, OG74o]. extended [FOO01, Goo79m, GF89]. extensions [Goo97a]. extraction [Goo83m]. eye [Goo81m]. eye-ball [Goo81m].

F [Goo63a, Goo91c]. F20 [Goo89x]. F21 [Goo89y]. F28 [Jos96a]. F44 [Goo96c, Jos96a]. F44A [Goo96c, Jos96b]. F44B [Jos96b]. F46 [Goo96d]. F49 [Goo96e]. F49A [Goo96e]. F6 [Goo87l]. F7 [Goo87m]. F8 [Goo87n]. Factor [CG96, Goo75a, Goo84s, Goo86s]. Factorial [Goo90h]. factorization [Goo71b, Goo72e, Goo93d]. factors [Goo81x, Goo85c]. Fallacious [Goo62d, Goo59a]. fallacy [Goo96d, Goo98a, Jos96a]. false [Goo61o]. false-alarm [Goo61o]. fast [Goo81p, Goo82n, Goo83l, Goo71e, Goo81p, Goo82n, Goo87f, HJB85]. fault [Goo89c]. Feature [Goo90w, Goo89e]. Features [GC68a]. fees [Goo79p, Goo79h]. Feynman [Goo80p]. Fibonacci [BG76b, BG76a, Goo74d, Goo93f, Goo93d, Goo93g, Goo94c]. Fifty [Goo90o]. Finetti [Goo69b, Goo88e, Goo89r]. fingerprints [Goo83j]. Fingertips [Goo58d]. finite [Goo51c, Goo52b, Goo55f]. first [Goo65e, Goo80f, Goo83-30, Goo83k, Goo83-31, Goo83-29]. Fish [Goo93h]. Fisher [GF89, Goo91c, Goo84r, Goo84t, Goo85n, Goo86l]. fisherman [Goo98a]. fit [Goo78s]. fixed [Goo83c]. flat [Ked12]. flaw [Per00]. flexible [Goo86p, Goo81-27]. Flowers [WG82]. flybutter [Goo94d]. fold [GH82a, GH82b]. Follow [Goo71a]. Fondements [Goo56a]. Form [Goo69a, Goo70b, Goo78t, Kha78, Goo86i]. formal [Ked12]. Forms [KG77, MG65, Goo86t]. Formula [Goo61c, Goo62a, Goo76e, Goo77c, Goo70c, Goo75g, Goo82q, Goo83g, Goo84r, Goo84t, Goo85f]. formulas [JK89]. Formulation [Goo63b]. forum [Goo90v, Goo19]. Foundations [Bir62, Goo55a, Goo83-35, Sav54, SBC^{+}62, Sav62, Sav72, Goo86n, Goo09]. Fourier [CG77a, Goo71e, Goo81p, Goo82n, Goo87f, HJB85, CG77b, Goo58e, Goo60d]. fourth [Goo94b]. Foutz [Goo81h]. Fractal [Goo84b, Goo84c, Goo84-27, Goo91a, Goo91c]. Fractals [Goo90r, Goo78t, Goo90b]. fractional [Goo41a, Goo89p, Goo89n, Goo89w, Goo79v, Goo83v]. Fractions [Goo93f, Goo41a, Goo58g, Goo83v, Goo83u, Goo83z, Goo89p, Goo89n, Goo89w, Goo93g]. frames [Goo80o]. Fraser [BG70]. Free [Goo71c, Goo83o]. Freedom [Goo73d]. Frequencies [Goo53a, Goo56c, Goo63d]. Frequency [Goo61k]. Function [Goo73b, Goo81j, GJ82, Goo83w, Goo84q, Goo87d, Goo88c, Goo88i, Goo90j, HG85]. functionals [Goo85k]. Functions [Goo61n, Goo72a, PG86, Goo40, Goo60a, Goo61h, Goo68a, Goo68b, Goo72d, Goo78n, Goo80r, Goo82e, Goo83e, Goo95e, LR77, OG74j, Goo80m]. further [Goo83n, Goo84t, GL88, Goo92d, Goo79w, Goo83y, Goo92c]. Future [Goo90a, Goo90u]. fuzzy [Goo78m, Goo94e].

G [Goo52a, Goo57a, Goo73a, RDF12, WG82]. G. [Goo90t]. Galton [GG90]. Games [Goo56b, Goo90h]. gamma [Goo79u]. gaps [Goo80e, Goo85t]. Gauss [Goo79-28, HJB85]. Gaussian [Goo60b, Goo63e, OG74a, OG74i, OG74j, OG74s]. Geiringer [Goo58a, Goo66a]. General [GMT^{+}12, Goo46b, GS87, RDF12]. generalisation [Goo52c]. Generalization [BG76b, BG76a, FG53, Goo65b, Goo68c, Goo70c, GC77, GL83, GS85a, Goo88i, GC90]. Generalizations [Goo60c, Goo85q, Goo86s, Goo79x]. Generalized [GG68a, Goo76e, Goo80s, GG67, GG68b, Goo79u, Goo81u, Goo81z, GL83, GS85a, Goo86l, OG74i, Goo81u, Goo81z]. generallzations [Goo79w]. Generating [Goo73b, MG80]. Generation [Goo73a, GG71c, GL84]. generations [Goo55d]. genetic [Goo81k]. geometric [GT77]. geometrical [Goo86h]. Geometry [Goo83-27, Goo84b, Goo84c, Goo84-27, GM87, GM89]. get [AG81]. Giere [Goo75e]. Gill [Goo70h]. Gillies [Goo87k, Goo87o]. Girschick [Goo56b]. given [Goo78m, Goo84d, MG80]. gives [Goo83t]. Global [GG72]. God [Goo85h]. Godambe [Goo80n]. Gödel [Goo68d, Goo68g, Goo69c]. Golden [Goo93f, Goo93g]. Goldstine [Ano81, SWT^{+}81]. Good [Ano09e, RDF12, Sin08, FOO01, Ano09a, Ano09b, Ano09c, Ano09d, Ano09f, Ano09g, Ano13, Ano21, Anoxx, Ban96, Dav51, Fie08, GE73, GBS08, Har75, Ked12, Lee95, LR77, Lin10, Mar90, McC96, MA66, OR09, Per00, Ree46, ST08, Wet75, Wil13, Zab23, vdV09]. goodness [Goo78s]. goodness-of-fit [Goo78s]. grading [Goo80g]. Graig [Goo87l]. Graphs [DG57]. group [Goo51c, Goo52b]. Groups [Goo63a].

H

[Ano81, Goo52a, Goo78a, SWT^{+}81, WG82]. H. [SWT^{+}81]. hadron [Goo90q]. hadrons [Goo90s]. Hakvunic [Goo84i]. Hakvunic-mean [Goo84i]. Hald [BLH^{+}60]. Hall [Goo56b]. Hamaker [Goo78a]. Haphazardness [Goo95c]. harmonic [Goo84x, Goo84h, Goo84w, Goo92b]. harmonic-mean [Goo84x, Goo84h, Goo84w, Goo92b]. Harold [Goo62c, Zel80]. Hartree [Goo51a, Goo51b]. having [Goo80m]. heart [GHDB89]. Heaviside [HG85]. Hedged [Wel94]. Hedges [Gol94]. Heinz [Goo91a, Goo91c]. Heinz-Otto [Goo91a, Goo91c]. Hellinger [GS85a]. Henri [Goo41c]. Henry [Goo94m]. Herdan [Goo57a]. Herring [Goo67e, Goo68i, Goo69c, Goo88a]. heteroscedasticity [Goo86l].

Hettmansperger [Goo86b]. hierarchical [Goo80u, Goo83-34, Goo94e, Goo83-29]. higher [Goo86s]. Hilda [Goo58a, Goo66a]. Histograms [Goo78l]. Historical [Goo85v, Goo85j, Goo85g, Goo86q]. History [Ano81, MHR80, SWT^{+}81, Goo79-29, Goo80u, Goo83k, GS83, Goo84r, Goo84t, HJB85]. Hitler [Goo95c]. homogeneity [Goo81n, Goo86j]. honest [Goo82b]. Honor [Zel80]. hot [Goo98a]. Hotel [SC74]. Human [Goo67c, Goo78a]. Hunting [GG80b, GG80c]. Hwang [FOO01]. Hypergeometric [Hal60]. hyperons [Goo57e]. Hypotheses [GM84, Goo75k, Goo83t, Goo94g, Goo95b]. Hypothesis [Goo63b, Goo87q, Goo83s, Goo85c, Goo90q, Goo90s, GC68a]. hypothyroidism [Goo94i].

I. [Ano21, ABH^{+}69, Ban96, Fie08, GE73, GBS08, Har75, LR77, McC96, Per00, RDF12, Ree46, ST08, Wet75, Zab23]. idea [Goo96a]. Ideas [GMM62, MGM62, Goo41c, Goo89q]. Ideen [MA66]. identical [Goo80m]. identities [Goo62g, Goo83y]. IEEE [FOO01]. IF [Goo83-31]. II [CG81, Goo92g, CG80, Goo50a, Goo58c, Goo61f, Goo61m, Goo79-29, Goo85t, Goo91d, Goo92e, Goo97b]. III [Goo97c]. illness [Aga75a]. Images [Goo91a, Goo91c]. imaginary [GT81, Goo83-31]. Impact [ST08]. Imperial [Sav62]. implications [GG71c]. importance [Goo79f]. important [Goo94j]. impossibility [Goo79-27]. impracticable [Goo81o]. imprecise [Goo03d]. improved [Goo83l, Goo86r]. improvement [Goo78s, Goo82k]. improvements [Goo87m]. improving [Goo87b]. including [Goo83r, OG74j, OG74s]. income [Goo81t]. Increase [GT56]. Increased [GT56]. increases [Goo80k, Goo83c]. Independence [CG96, Goo63c, Goo66e, Goo85u, GC87, Goo66c, Goo86s]. independent [Goo94f]. Index [Goo56e, Goo57b, Goo79n, Goo82h, GS83, Goo85j, GS85a]. indexes [Goo82l, Goo89s]. Indices [Goo87c]. individual [GT76]. individuals [Goo49]. Induction [Goo52a, Goo87o, Goo81a, Goo83b, Goo83-28, Goo85b, Goo87k, Goo69b]. inequality [Goo50c, Goo88b, Goo88g, Goo90j]. inevitable [Goo85b]. inexactification [Goo93b]. infantile [Aga75a]. Inference [Bir62, Doy94, Goo75b, Goo76a, Goo84a, Goo86b, SBC^{+}62, Sav62, BG70]. Inferences [Goo95h, Sto93]. infinite [Goo42]. influence [Goo76b, Shi25]. influentiality [Goo79n]. Inform [FOO01]. informal [Goo86o]. Information [Ano53, Goo60g, Goo62b, Goo66d, Goo68e, OG74k, OG74l, Goo56d, GD58, GD59, Goo71d, OG74e, OG74f, OG74g, OG74o, Goo64]. informational [GL83, Goo81e]. Informative [Goo55b]. Initial [Goo69b]. Innovation [SWT^{+}81]. Inspection [Hal60, Wet60]. Instruments [Goo51a, Goo51b]. integer [Goo72h]. Integral [Goo57f, GR48, Goo50a, Goo52c, OG74i]. integrals [Goo80p]. Integration [GT78, Goo61b, Goo62e, GG70, GG71a]. Intelligence [RN95, RN98, RN10, RNC^{+}21, Yud13, Goo70f, Nor12]. intensive [Goo86i]. Interaction [Goo58e, Goo60d]. Interactions [Goo82l]. interests [Goo81f]. Interface [Goo88h, Goo88j, Goo89z]. intermediate [Goo83t]. Internat [Goo78a]. International [Goo81g]. Interpolation [Goo78p]. interpretation [Goo83s, Goo84o, Goo84m, Goo84n, Goo89f]. Interpretations [Goo62d, Goo85u, Goo59a]. interquantile [Goo03a]. interview [GE73]. Introduction [Goo75b, Goo76a, OG74m]. Introductory [Goo92g]. intuitive [Goo87n]. Invariant [JG81, Goo78f, Goo79l, Goo81r, Goo90m]. Inverse [Goo70d]. Inversion [Goo50b, Goo79-28]. Irreconcilability [Goo87q]. irregular [Goo82m]. irrelevance [Goo88e]. Irving [MA66, Ano09a, Ano09e, Ano09d, Ano13, Anoxx, Lee95, Lin10, OR09, Wil13]. Irwin [Goo84r, Goo84t]. Isolikelihoods [Goo90g]. IV [Goo98c, Goo99].

J [Ano13, Ano21, ABH^{+}69, Ban96, CG77a, Dav51, Fie08, FOO01, Goo55a, Goo62b, Goo72b, GE73, Goo74b, GBS08, Har75, Ked12, LR77, McC96, Per00, RDF12, Ree46, Sav62, ST08, Wet75, Zab23]. J. [SWT^{+}81, WG82]. J.-J [FOO01]. Jack [Ano09f, Ano09g, Lee95, Wil13, Ano09b, Ano09c, Sin08, vdV09]. Jacobians [Goo81z]. James [Goo82a]. Jeffreys [Goo62c, Goo80v, Goo83-28, Zel80]. Jewels [Oyv26, Oyv52]. John [Ano09a, Ano09e, Ano09d, Anoxx, Goo56b, Lee95, Lin10, MA66, OR09, Wil13]. Joint [Goo73b, Goo55d, Sav62]. Joseph [Goo75d, JL95]. Jr. [Goo94m]. JSCS [Mar90]. judgement [Goo82b, Goo89t]. judgements [Goo79j]. judging [Goo90s]. judgments [Goo58b, Goo58c, Goo87b]. judicial [GT85]. Jupiter [Goo84e, Goo85g]. just [Goo84v].

Kempthorne [Goo84j]. key [Goo94f]. Kinds [Goo59b, Goo80f, Goo80l, Goo83k]. kinematics [Goo94n, Goo97b, Goo97c, Goo98c, Goo99, Goo03b, Goo03c, McC96]. Knowledge [Gol94, Goo61d]. Known [PG86]. Koheleth [Goo85h]. Kolmogorov [Goo78i, PG76, PG83]. Kudryashov [FOO01]. Kyburg [Goo94m, Doy94, Gol94, Isr94, Lou94, She94, Sme94]. Kyburgian [Wel94].

L [Goo55a, Goo73a, Goo91c, SWT^{+}81]. L. [Sav62]. lady [Goo90l, Goo92c]. Lagrange [Goo60c, Goo65b, Goo68c, Goo75f, Goo76e]. Landsberg [Goo61d]. Language [Goo57a].

Languages [SWT^{+}81]. Lansing [SC74]. Laplace [Goo84e]. lapses [Goo87g]. large [AG81, Goo83f, Goo89f]. Lattice [Goo59c]. Law [Goo69f, Goo71a, Goo81-28, Goo85o, Goo85h, Goo87n, Goo93e]. Laws [JG81]. leads [Per00]. Learning [Goo74c]. least [Goo79f, Goo83w, Goo84y]. Lebesgue [Goo41c]. Lecture [WG82]. Lecturer [Goo57d]. Lee [SWT^{+}81]. Legendre [Goo55f, Goo58f]. Legends [Ano21]. lemma [Goo80o]. Lengths [Goo73b]. Letter [Goo65c, Goo68h]. Letters [BDR^{+}72, BBC^{+}89, BRG^{+}91, CTL^{+}77, DTB^{+}70, DMK^{+}70, DGK^{+}75, GGF^{+}01, GFG^{+}65, GBC69, GFM^{+}77, GN85, HPG^{+}71, HLT^{+}87, JEE^{+}83, KBE^{+}73, KsLG^{+}74, KLC^{+}88, LFB^{+}72, LOG^{+}69, PCM^{+}74, SKR^{+}71, VKT^{+}69, WSG^{+}73, WTA^{+}90]. Liapounoff [Goo50c]. lie [Goo90e]. lie-detectors [Goo90e]. life [Zab23]. Likelihood [Goo55e, Goo57c, GGM70, GGM71, GC78, GG80b, GG80c, GG77, Goo78f, Goo79i, Goo82i, Goo82g, Goo83-29, Goo83f, Goo86l, Goo86r, Goo87i, LR77, PBB^{+}71]. Likelihood-Ratio [GGM70, GGM71, GC78, Goo86l]. likelihood/minimum [GL85]. likelihoods [JK89]. Lin [FOO01]. Lindley [Goo82u]. line [OG76]. Linear [DG57]. linearity [Goo87g]. List [OG74n]. literature [Goo87m]. Little [Goo74c, Goo94h]. local [Goo40]. logarithmic [Goo83g]. Logic [Goo67c, Goo85o]. lognormal [GH83b, Goo83r, Goo83y]. London [Ano53, Goo56b, Goo81-30, Sav62]. Look [Goo71a]. Lorentz [Goo95f, Goo00a, Goo01]. loss [GM72, Goo88c]. Lower [PG76]. Lucas [Goo93f, Goo93d, Goo93g, Goo94c]. luck [Goo93c].

M [Bri92, FOO01, Goo56b, Goo56a, Goo79-29, Goo91c, Goo92g]. M.-C [FOO01]. Machine [Goo67c, Goo58b, Goo58c, Goo65e, Goo70f]. Machines [Goo65a, Goo51a, Goo51b]. MacMahon [Goo62g, Goo62h]. MacMillan [Goo70h]. mainly [KG89]. major [OG74s]. make [Goo58b, Goo58c]. Mandelbrot [Goo91c, Goo78t, Goo84b, Goo84c, Goo84-27]. Marginal [PG86, Goo81f]. margins [GG80a]. marking [Goo55g]. Markoff [Goo57c, Goo55e]. Markov [DG57, Goo61k, Goo63d, Goo73b, Goo89m, Goo97a]. Markov-Chain [Goo63d]. Markovity [Goo83v, Goo89n]. mass [Goo84e, Goo85g, Goo89v]. masses [Goo57e, Goo89u, GJH90]. Master [Goo62g, Goo62h]. matched [Goo85s]. Math [CG77a, Ked12]. Math. [Goo82q]. mathematical [Goo77b, Goo79f, Goo80w, Goo81-30, Goo66a]. Mathematically [Goo85q]. Mathematician [Ano09f]. Mathematicians [Ano81, Rej81]. Mathematics [Bri92, Goo65a, Goo95g, Shi25]. Matrices [Goo50b, Goo71b, Goo72e, Goo86t, Goo95f, KG89]. Matrix [Goo69e, Goo70d, JG81, Goo61g, Goo88b, Goo88g]. maximizes [Goo86r]. Maximum [Goo63b, Goo78f, Goo82g, Goo83f, GL85, Goo86s, Goo87i]. maximum-likelihood [Goo78f, GL85]. maximum-likelihood/minimum-discrepancy [GL85]. McCausland [Per00]. McGuire [Goo91c]. McShane [Goo70h]. mean [Goo84x, Goo84h, GBSS84, Goo84i, Goo84w, Goo92b, MG80]. meaning [Goo79l]. meaningfully [Goo81a]. means [Goo62g, GHDB89]. measurable [Goo40]. measure [Goo62i, Goo83d, GL83, Goo85i, Goo93e]. measureable [Goo62i]. Measurement [Goo82s, Goo82j]. measurements [GHDB89]. Measures [OG74l, Goo81d, GL83, Goo84s, Goo85q, GS87, LSG89, OG74q, OG74r]. mechanics [Goo82d]. Medialized [Goo89l, Goo90k]. Meeting [Sav62, SC74]. Meetings [WG82]. memory [Goo86o]. meromorphic [Ked12]. Mersenne [Goo55c, Goo80e, Goo85t]. Messrs. [BLH^{+}60]. Metaphysics [JL95]. metaprinciple [Goo89i]. Meteorite [GG80b, GG80c]. Method [Goo60e, Goo61l, GLGH79, GG80b, GG80c, Goo46b, GG70, GG71a, Goo78m, Goo78e, Goo81o, Goo82f, Goo84q, GL85, Goo87b]. methodology [Goo80u]. Methods [Goo57h, Goo61i, Goo65d, GMT^{+}12, RDF12, Goo42, Goo44, Goo60a, Goo61h, Goo61j, GHJT69, Goo70f, Goo80l, Goo82d, GS85b, GL86]. Michael [Goo90b, Goo90r]. Michie [RDF12]. Michigan [SC74]. Microeconomics [Yud13]. middle [Goo85o]. midrank [HG85]. Miller [Aga75b, Gil86, Goo70g, Goo75c, Goo89e, Goo90w, Ros75]. mineral [Goo79j]. Minicommunications [Goo76c, GSJ76, GSK76]. minimal [Goo78q]. minimal-vocabulary [Goo78q]. minimax [Goo94k]. minimum-discrepancy [GL85]. miscellaneous [Goo90m]. Mises [Goo58a, Goo66a]. misstatement [Goo83x]. Mixed [GC87, Goo81w, Goo83-29]. Mixed-Dirichlet [GC87]. Mixture [CG70, CG71]. Mixtures [CG80, CG81, Goo76d]. Möbius [GC68a]. Mode [Goo72b]. model [Goo60b, GM72, Goo83c, Goo83-34, Goo86p, Goo86e, Goo89m]. models [GS86a]. Modern [Goo65d, RN95, RN98, RN10, RNC^{+}21, GHJT69]. Modification [GM84]. moment [And89, Goo44]. moments [Goo79v, GH82b, Goo83q, Goo83t, Goo83z, Goo84s, Goo78i, Goo79u, Goo83p]. monocular [Goo86k]. Monotonic [Goo84s]. monotony [Goo40]. Monozygotic [Goo78o]. Monte [Goo60e, Goo81x, Goo81o, Goo81j, Goo86q]. Morgan [BG76a]. most [Goo72h]. motion [Goo51c, Goo52b, Goo58g]. MR [Goo81-30, Goo89w, Goo82q]. MR0888434 [GM89]. MR1202337 [Goo93g]. MR1324249 [McC96]. MR1476810 [FOO01]. MR1646418 [Goo99]. MR1867077 [Goo01]. MR2682186 [Ked12]. Much [Goo58d]. Multidimensional [Goo63b, GT76, Goo81r]. Multinomial [Goo57h, Goo61i, Goo61l, Goo67b, GGM70, GGM71, GC74, Goo75a, GC78, GG77, Goo61j, Goo74b, Goo78m, Goo81p, Goo82n, GH83a, Goo84d, GL85, Goo86r, LR77, Goo86r]. multinomials [CG82, Goo79m, Goo83-34, GL85, GL88, HG86]. multinormality [GS87]. multiple [Goo52c, Goo61b, Goo62e, Goo79h, Goo83n, Goo84f]. multiple-choice [Goo79h, Goo83n]. multiplications [Goo85e]. multiplicative [GG71c, GH82a, GS86a]. multiplying [Goo85e]. Multitype [Goo72b]. Multivariate [Goo56e, Goo57b, Goo61l, LSG89, Goo71b, Goo72e, Goo82f, GH83b, Goo90f]. Must [BH96]. mutual [Goo79-27, OG74e, OG74f, OG74g, OG74o]. My [Goo95g, Goo83s, Goo83-28].

N [SWT^{+}81]. natural [Goo85q]. Nature [Goo84b, Goo84c, Goo84-27]. Naval [Goo00b]. Necessarily [Goo82t]. Necessary [Kha78, Goo93a]. necessitivity [Goo94l]. need [Goo89k]. Needle [Goo57d]. neighbours [Goo87p, Goo91d]. Neo [Goo86v]. Neologisms [Goo81v]. nets [Goo81k]. neutron [Goo57e, Goo89u]. Newcomb [Goo86n]. newly [Goo86m]. Newman [Goo73a]. Neyman [Goo82c, Goo83s, Goo84u]. niche [LSG89]. Nielsen [Goo83w]. no [CG77a, FOO01, Goo81-30, Goo82q, GM89, Goo93g, Goo99, Goo01, McC96]. noise [OG74a, OG74j, OG74s]. Nombres [Goo66b]. Non [KG77, Goo62i, GH83a, GS84]. Non-Bayes [GC74, Goo92a, Goo83-30, Goo86i, Goo86f]. non-equiprobable [GH83a, GS84]. non-measureable [Goo62i]. Non-Normal [KG77]. noncentral [GJ82]. Nonparametric [GG71b, GG72, Goo84v, Goo84q, Goo82a]. nonrandom [OG74j]. Normal [Goo46a, Goo56e, Goo57b, KG77, Goo78p, Goo81-27, Goo87d, MG80]. normality [Goo96a]. normally [Goo87d]. Norman [Goo63a, Goo64]. normative [Goo89k]. notation [Goo56d]. Note [Goo41b, Ree46, Goo47, Goo83m, Goo84t, Goo85g, Goo86u, GL88, Goo92d, LS86, Goo67a, Goo72a]. Notes [BRCG55, Goo61i]. noticed [Goo84u]. Novel [Goo85v]. Null [Goo87q, Goo85c, Goo94g]. Number [GT56, Goo75h, Goo49, Goo68h, GG71c, Goo80d, Goo81p, Goo82n, GH82b, Goo83j, Goo85k, Goo85e, Goo87j, Goo88f, Goo89d]. Numbers [Goo71a, Goo74d, Goo86w, Goo86x, Goo93f, Goo53b, Goo55c, Goo69d, GT77, GT78, Goo79e, Goo86a, Goo93d, Goo93g, Goo94c, Goo98a]. Numerical [Ano81, Goo57f, Goo69f, SWT^{+}81, GG70, GG71a, Goo85e]. numerological [Goo89t, Goo89u]. numerology [Goo88f, Goo89v, Goo90s, Goo92d, Goo94a, Goo95a, GJH90]. nutrition [Goo78g].

O.K. [Goo82g]. Obituaries [Ano09d]. Obituaries/Passings/Irving [Ano09d]. Obituary [Ano09e, Ano09f, vdV09]. ’Objective [Goo69f, PG86]. Objectivity [Goo78a, Goo84v]. observation [Goo79-27]. observations [Goo83o, Goo89h]. Observed [Goo71a]. obverse [Goo94d]. occurring [Goo84z]. Octav [Goo66b]. October [SC74]. odd [Goo83u]. odds [Goo75k]. Odell [Goo73a]. oedemology [GG78]. Olds [SC74]. omissions [Goo80d]. One [PG76, PG83, Goo78i, GL83, Goo86l, Goo00a, Goo01, LS86, Goo84h]. One-Sample [PG76, PG83, Goo78i, LS86]. one-way [Goo86l]. Onicescu [Goo66b]. only [Goo84v]. only-just [Goo84v]. Opened [Sav62]. operators [OG74i]. optional [Goo91b]. Order [Goo63d]. ordering [GT76]. Orderings [Goo75h, Goo80d]. organic [Goo98a]. Oriented [DG57]. orthogonal [KG89]. Other [OG74o, Goo53b, Goo83f, Goo87m, Goo93b]. Otto [Goo91a, Goo91c]. outliers [Goo78h]. overflows [Goo78e]. overlap [LSG89]. Own [Gre22].

P [Goo61d, Goo73a, Goo86b]. Paired [Goo84i]. Pairs [CG77a, CG77b, GL84, GL88]. Paper [ABH^{+}69, Sme94, Goo69b, Ree46]. Papers [BRCG55, CMU^{+}53, CMU^{+}93, SW93, Ano13, BLH^{+}60, FOO01]. Paradox [Goo60f, Goo61m, Goo82u, MG65, GD58, GD59, Goo80n, Goo82c, Goo86h, Goo89h, Per00, Goo70g]. Paradoxes [Goo72f]. paradoxical [Goo94h]. paragraph [Goo83-30, Goo83-31, Goo83-29]. Parallel [Goo58i, Goo84w]. parameter [Goo86r]. parameterization [KG89]. Parameters [Goo53a]. parametric [Goo87i]. Park [Goo79z, GMT^{+}12, Gre22, HS93, RDF12, Shi25]. Part [CG81, CG80]. partial [Goo41c, Goo83u, Goo83z, Goo92f, Goo79t, Goo92e, Goo92f]. partially [Goo89v]. partition [Goo70c]. Partitioning [Goo79i]. Partly [GMM62, MGM62]. Partly-Baked [GMM62, MGM62]. Parts [Oyv52]. Passings [Ano09d]. path [Goo80p, Goo78n]. Patrick [Hum94]. Pearson [Goo81p, Goo81j, Goo82r, Goo82n, Goo82c, Goo83s, Goo83l, GS84, Goo90l, HG86]. Peirce [Goo81c, Goo83s]. Peitgen [Goo91a, Goo91c]. Penalized [GG80b, GG80c, Goo83-29, Goo83f, Goo87i, Goo84y]. Penalties [GG71b, Goo81r]. Perks [Goo71g]. permanents [GL83]. Permitted [Goo75h]. permutability [Goo86j]. permutation [Goo82h]. permutations [Goo81d]. Phantasie [MA66]. Philip [Goo70h]. Philos [Goo82q, Goo89p, Goo89w]. Philosopher [Hum94]. Philosophy [Goo83-33, Goo84a, Goo88h, Goo88j, Goo95g, SC74, Goo83-30, Goo89z, Goo93e, Goo94f]. Phys [Goo01, McC96]. Phys. [Goo99]. physical [Goo88f, Goo90s, Goo92d, Goo95a]. piece [Goo88f]. pieces [Goo89a]. Pink [Goo68i]. Pioneering [Goo80t]. Pioneers [Lee95]. Plans [Hal60]. plausibility [Goo83w, Joh90]. players [Goo55g]. Plaza [SC74]. pling [Goo86g]. poem [Goo86m]. Point [Goo87q, Goo57h, Goo61i, Goo82i, Goo83h, Goo86e]. Poisson [GF89, DE86, Goo62a, GBSS84, Goo86y, Goo86z, Hey86, HG86, Sin86, Sol86]. polarization [GS86a]. Polish [Ano81, Rej81]. Pólya [FG53]. polynomials [Goo55f, Goo58f, Goo85e]. polypeptide [Goo82m]. Popper [Gil86, Goo89e, Goo89g, Goo90w]. Popper/Miller [Goo89e, Goo90w]. Population [Goo53a, GT56, Goo75a, GLGH79, Goo84d]. Positive [Goo56a, Goo47, Goo81w, Goo83o]. possible [Goo68h, GS84, Goo84m]. possibly [Goo86m]. postulate [Goo82b]. Postulats [Goo56a]. potential [Goo79j, Goo83x]. Power [Goo60g, Goo68e, Goo72a, Goo72d, Goo78n, Goo79o, Goo83e, Goo83m, Goo85r, GS86b]. power-summing [Goo79o]. powerful [Goo81i]. Powers [Goo61c, CG82, Goo93d]. Pp [Goo56b, Mar90]. Practical [Goo58e, Goo60d]. Pragmatic [GM84, Goo88e]. Preceding [CMU^{+}53, CMU^{+}93]. Prediction [MG65, Goo86r]. predictions [Mil75a, Mil75b]. predictive [GL81, Goo83b, Goo79r]. preference [Goo89j]. preposterior [Goo83-31]. prerequisite [Goo69b]. primes [Goo80e, Goo85t]. principal [GS85b].

Principle [Goo67d, Goo83w]. Prior [Goo75a, Hal60]. priors [Goo81-27, Goo83-29]. Probabilistic [Goo66d, BH96, Goo71d, Goo85o, Goo93e, Goo85b]. Probabilities [DG57, Goo65d, PG86, Goo61o, Goo69b, GHJT69, Goo78m, Goo79r, Goo82p, Goo84g, Goo84m, Goo85r, Goo85c, GL85, Goo83z]. Probability [Goo50d, Goo66a, Goo66f, Goo68f, GG71b, Goo72f, Goo73b, GL81, Goo82a, Goo83-35, Goo86-27, OG74l, Goo58b, Goo58c, Goo59b, Goo62i, GG72, Goo78j, Goo79-29, Goo80e, Goo81b, Goo81-28, Goo81r, Goo83u, GL84, Goo84l, Goo84n, Goo85t, Goo86n, GS86b, Goo87j, Goo87b, GL88, GHDB89, Goo90n, Goo94e, Goo95d, Goo09, HG86, OG74q, OG74r, Goo82e, Goo83f, Goo84p, Goo85l, Goo52a, Goo62c, Dav51, Goo58a]. Problem [CG70, CG71, Goo79f, GBSS84, Goo86l, Goo86n, Goo86g]. Problems [AC84, And74, AF75, BKM^{+}67, BvL68, Ber73, BK75, CGI78, CB68, CGGR68, Ehr74, EB75, ES76, ECG^{+}77, EP78, Gil74, GBMW75, GK75, GRM76, GSA75, GT68, Goo72g, Goo73c, Goo75i, GR78, GB80, Goo81-31, GB82, GC68b, HK75, HSO^{+}81, HS82, HG68, KC75, KS68, LB68, MC75, MH70, MTD68, MR78, MW80, MS82, MM75, OK84, Oso78, PJR^{+}68, RF68, RC74, SB74, SS84, SH74, SAG68, SG68, TLM^{+}66, TQ68, VM68, WTR68, WF68, WV68, Goo79v, Goo90d]. Proc [Goo81-30, Goo82q, Goo89p]. Proc. [Goo89w]. procedure [Goo89x]. procedures [Goo84v, HG85]. proceedings [SC74, Ano53]. process [Goo49, Goo55d, Goo82k, Goo89m, GG90]. Processes [Goo72b, Goo75f, Goo60c, OG74i, OG74f, OG74j, OG74q, OG74s]. product [Goo88g]. Products [Goo94o, Goo88b]. profession [Goo78a]. Professor [Ano09f, Ano09g]. Programming [SWT^{+}81, Goo90h]. projection [Goo82f]. projection-tomographic [Goo82f]. Proof [Goo70e, Goo50c, Goo62h, Jos96a, LR77]. Proofs [Goo62g]. Proper [Goo79h, Goo79p]. properties [GH82a, Goo84s]. Property [Goo94o]. Proportional [GLGH79, Goo87d]. proposal [Goo79o, Goo82k]. PROTOM [Goo82f]. proton [Goo57e, Goo89u, Goo89v]. provided [Goo81y, Goo81-32]. PSA [SC74]. Pseudorandom [GC68a, GG71c]. publication [Goo88a]. Publications [Anoxx]. punctured [FOO01]. Pure [Bri92].

QH [Goo96a]. Qua [Goo68i]. Quadratic [Goo63c, Goo66e, Goo69a, Goo70b, Kha78, KG77, Goo66c, Goo71b, Goo72e, GJ82, GS83]. Quadratics [Goo63d]. qualitative [GS86a]. quantal [Goo90q, Goo90s]. quantiles [Goo79j]. Quart [Goo93g]. quasimodel [Goo86e]. Queries [GM96]. question [Goo83z, Goo96c]. questioned [Goo81a]. questionnaires [Goo90e]. Questions [RG69]. quick [Goo84q]. quotients [Goo83u, Goo83z].

R [Goo51a, Goo51b, Goo82a, Goo91c]. R. [SWT^{+}81]. Ramanujan [GBSS84]. Random [Goo51c, Goo58g, Goo72f, Goo73a, Goo74e, FG53, Goo57g, Goo58f, Goo69d, Goo79s, Goo80m, OG74f, OG74q, Goo52b, Goo90e]. Random-response [Goo90e]. random-walk [FG53]. randomization [Goo92c]. Randomly [Goo81k]. Randomness [Goo74e, McS70, Goo53b, Goo95c, Goo70h]. range [Goo85l, Goo03a]. rank [Goo85p, Goo94i]. Ranks [Goo86b]. Rate [OG74p, FOO01, GD58, GD59, GS83, Goo84d, Goo89c, OG74g]. rates [Goo60b]. Ratio [Goo55e, Goo57c, GGM70, GGM71, GC78, Goo93f, KG77, GG77, Goo79i, Goo86l, Goo89v, Goo93g]. Rational [FOO01, Goo52d]. Rationalism [JL95]. Rationality [Goo69f, Goo89k, Goo94k].

Razor [Goo68f]. re [Goo79r]. re-use [Goo79r]. reactor [Goo88d]. Real [Goo61n]. really [AG81, Goo94h]. Reciprocal [Goo74d]. Reciprocals [Goo94o]. Recognizability [Goo86j]. recurring [Goo46a]. recursive [Goo85e]. Red [Goo67e, Goo69c, Goo88a]. Redhead [Goo87k, Goo87o]. reductio [Goo81b]. reduction [Goo90f]. refereeing [Goo82k]. reference [Goo90k]. reflections [Per00]. Refutation [Goo93b]. regarding [Goo85g, Goo86u]. Regeneration [OG76]. Regenerative [Goo73b]. regression [Goo80i, Goo89x, Goo92b, Goo96d]. regret [Goo82j]. regularity [Goo44, Goo46b]. Reinstatement [Goo87o, Goo87k]. Rejection [Wel94, Goo93b]. Rejoinder [GG80c, Goo86z, Goo88j]. related [GC77, Goo78m, Goo83y, GL84, GL85, GL88, Goo95b]. relations [Goo42]. Relationship [Goo71e, Goo76e, Goo80f, GL83, Goo84s]. Relationships [OG74q, GG71c]. relative [Goo75k, Joh90]. relativity [Goo94n, Goo97b, Goo97c, Goo98c, Goo99, Goo03b, Goo03c, McC96, Per00]. relevance [GT81, GL83]. relevant [Goo93e]. reliability [BH96, Goo95d]. Remark [Kha78]. Remarks [Goo92g, Wet60]. Removal [GLGH79]. repeat [GS83, Goo84d]. repeats [Goo81p, Goo82n, GH82a, GH82b]. Replies [Goo75j]. Reply [Gol94, Goo78a, Goo90t, Goo96b, Jos96a, Jos96b, McC96, Mil75b]. Report [Ano53, GMT^{+}12, RDF12, WG82]. Representation [JG83]. research [Goo70f, OG74b]. residues [Goo83z]. Resolution [Goo70g]. resources [Goo79j]. respect [Goo84d]. Response [BGHK94, Goo87o, Goo87k, Goo90e]. Rest [Sto93, Goo95h]. restricted [Goo92c]. results [FOO01, Goo83-31, OG74s]. Retrospect [WG82]. Rev [Goo78a]. Review [Dav51, Goo51a, Goo51b, Goo52a, Goo55a, Goo56b, Goo56a, Goo57a, Goo58a, Goo61d, Goo62c, Goo62b, Goo63a, Goo64, Goo66a, Goo66b, Goo70a, Goo70h, Goo72c, Goo72b, Goo73a, Goo75b, Goo76a, Goo78b, Goo78c, Goo78d, Goo78t, Goo79a, Goo79b, Goo79c, Goo80b, Goo80c, Goo82a, Goo83a, Goo84b, Goo84c, Goo84-27, Goo85a, Goo86b, Goo87a, Goo90b, Goo91a, Goo91c, Goo92a, Goo95g, Goo95h, MA66, Goo94l, Goo95f, Goo97a, KG89]. Reviews [BG70, Goo58h, Goo71f, Goo76f, SWT^{+}81, Ano81]. Richard [Goo58a, Goo66a, Goo82a, Goo95h]. Ridge [Goo92b]. Riemann [GC68a]. risk [Goo88d]. Risque [Goo56a]. Rissanen [GC90]. River [Ano09e]. Robustness [GC87, Goo83-34, GS86a]. Ronald [Goo75e]. Root [Goo93f, Goo93g]. rotation [Goo81r]. Roughness [GG71b, Goo97a, Goo81r]. rows [Goo81d]. Roy [Goo74b, Goo81-30]. Rule [Goo71a, Goo84x, Goo84h, Goo84i, Goo84w, Goo89g, Goo92b]. Rules [Goo55b]. Run [Goo73b]. Run-Lengths [Goo73b]. runs [Goo79d].

S [BG70]. Saddle [Goo57h, Goo61i]. Saddle-point [Goo57h, Goo61i]. Saddlepoint [Goo61l, Goo61j]. salvaged [Goo60b]. sam [Goo86g]. Sample [GT56, PG76, PG83, Wet60, Goo78i, Goo79r, Goo80k, Goo82o, Goo83c, Goo84g, LS86, MG80]. Sampling [Cox60, Hal60, Goo53b, Goo60b, Goo83-32, Goo86o]. satisfied [GG71c]. Saturated [Goo86e]. Saupe [Goo91a, Goo91c]. Savage [Goo55a, Sav62]. Scattering [GG80b, GG80c]. Scheme [Wet60]. Schemes [Cox60]. Science [Goo58d, Goo88h, Goo88j, Goo91a, Goo91c, JL95, SC74, Goo89z, Goo94f, Goo94j]. Scientific [Hum94, Goo81a, Goo81-28, Goo83b, Goo83-28]. Scientist [GMM62, MGM62, Anoxx, Goo79n]. search [FOO01]. second [Goo80f, Goo83k]. secret [Zab23]. see [Goo83-30, Goo83-31, Goo83-29]. self [Goo94n, Goo97b, Goo97c, Goo98c, Goo99, Goo03b, Goo03c, McC96]. self-consistency [Goo94n, Goo97b, Goo97c, Goo98c, Goo99, Goo03b, Goo03c, McC96]. Seminar [Sav62]. Sensitivity [GC87]. separateness [Goo82l, Goo82h, Goo85j]. September [Ano53]. Sequence [GC68a, Goo71a]. sequences [Goo57g, GG71c]. Ser [Goo74b, Goo81-30]. Serial [Cox60, GG68a, Goo87p, Goo91d, Goo53b, Goo57g, GG67, GG68b, And89]. Series [BG76b, BG76a, Goo58i, Goo74d, Goo41b, Goo42, Goo55f, Goo83m, GS86b, Goo90d]. set [Goo62i, Goo71b, Goo72e]. several [Goo60c]. Sewall [Goo80p]. Shakespeare [Goo86m]. Shannon [SW93]. shape [Goo88c]. shapes [Goo82m]. sharp [Goo94g]. Sharpened [Goo68f]. sharpening [Goo84w]. Sherman [Goo81h]. Shoe [Goo67e, Goo68i]. Short [BRCG55, Goo70e, Goo62h]. should [Goo81f, Goo84g, Goo88c, Goo78h]. shrinking [Goo86r]. SIAM [CG77a]. sic [GJH90]. sign [Goo94i]. sign-rank [Goo94i]. Signal [OG74l, Goo60b, OG76]. signals [OG74a, OG74j, OG74s]. signed [Goo85p]. signed-rank [Goo85p]. Significance [Goo55h, Goo58i, Goo67b, Goo74b, Goo80k, Goo82h, Goo83c, Goo86c, Goo87h, Goo94j]. significant [Goo88f]. signs [Goo78j, Goo79d]. similarity [GS85a, GS87, LSG89]. Simple [Goo71a, Goo83g, Goo88i, GG90]. simplex [GT78]. Simplicity [Goo68f, Goo81f]. simplification [Goo80q]. sine [Goo87n]. Single [Hal60, Wet60]. Singular [Goo69e, Goo80o, GS86a, Goo88b, Goo88g]. Singularity [OG74l, OG74r, OG74q]. size [Goo80k, Goo83c, Goo84g]. sizes [Goo55d]. Skew [Goo86w, Goo86x, Goo86a]. Small [Goo56c, Goo84l, Goo85f, Goo86c]. smallest [Goo88b, Goo88g]. Smirnov [Goo78i, PG76, PG83]. Smith [Mar90]. Soc [Goo74b, Goo81-30, Goo89p, Goo89w]. Soc. [Goo82q]. social [GM72]. societal [GM72]. society [Goo82d]. soft [FOO01]. soft-decision [FOO01]. Solution [Goo57f, OK84, Wet60]. Solutions [AC84, And74, AF75, BKM^{+}67, BvL68, Ber73, BK75, CGI78, CB68, CGGR68, Ehr74, EB75, ES76, ECG^{+}77, EP78, Gil74, GBMW75, GK75, GRM76, GSA75, GT68, Goo73c, Goo75i, GR78, GB80, Goo81-31, GB82, GC68b, HK75, HSO^{+}81, HS82, HG68, KC75, KS68, LB68, MC75, MH70, MTD68, MR78, MW80, MS82, MM75, Oso78, PJR^{+}68, RF68, RC74, SB74, SS84, SH74, SAG68, SG68, TLM^{+}66, TQ68, VM68, WTR68, WF68, WV68]. Some [FOO01, Goo42, Goo56d, Goo69e, Goo70f, GG71c, Goo72f, Goo80u, Goo86y, Wet60, Goo60a, Goo61h, Goo62g, Goo79e, Goo79v, Goo83r, Goo84x, Goo85q, Goo87g, Goo90m, DE86, Goo81h, Goo86f, Goo86z, Goo87b, GS87, Goo87m, Goo89v, Hey86, Sin86, Sol86]. somewhat [Goo87i]. Sons [Goo56b]. source [Goo63e]. Space [Goo59c, GT76]. special [Goo86t, Goo94n, Goo97b, Goo97c, Goo98c, Goo99, Goo03b, Goo03c, McC96, Per00]. Species [Goo53a, GT56, Goo82o, Goo86g]. species-sam [Goo86g]. specified [Goo89m]. Speculates [GMM62, MGM62]. Speculations [Goo65e, Goo90u, Goo90a]. speculative [Goo89q]. Speed [Goo71c]. spherical [Goo79t, Goo92e]. Spies [Smi19]. Spirit [Goo78a]. Square [Goo85u, Goo93f, Goo93g]. Squared [Goo61l, Goo69a, Goo70b, Kha78, Goo78s, Goo81i, Goo81p, Goo81j, Goo81n, Goo82r, Goo82n, Goo83l, GS84, HG86, Mag81]. squareds [GG80a]. squares [Goo79g]. Stable [Goo61n]. standard [Goo82o]. standardized [Goo84g, Goo82p, Goo84m]. Statist [Goo74b, Goo78a, GM89]. Statistic [GGM70, GGM71, GC78, Goo85u, PG76, PG83, GG77, Goo78i, Goo85s]. Statistical [Bir62, CG96, DE86, Goo56b, Goo86y, Goo86z, Goo90v, Goo92g, GMT^{+}12, Goo19, Hey86, RDF12, SBC^{+}62, Sav62, Sin86, Sol86, Goo70f, Goo77b, Goo78a, Goo79e, Goo79-29, Goo80w, Goo81l, Goo81-30, Goo82h, Goo82d, Goo87m, GHDB89, Goo90j, Goo90d, Goo86b]. statistician [Goo84v]. statisticians [KG89]. Statistics [Goo58a, Goo70h, Goo72g, Goo77c, Goo88h, Goo88j, Goo90a, McS70, Sav54, Sav62, Sav72, Zel80, Goo79-29, Goo80v, Goo84z, Goo89z, Goo90m, Goo90u, Goo55a, Goo66a]. Stirling [GT77, Goo85f]. stochastic [Goo60c]. Stone [Goo95h]. stooge [Goo84v]. stoogian [Goo89h]. stopping [Goo91b]. Story [HS93]. Straddling [Goo03a]. strategies [Goo68h]. strengths [CG82]. Structural [JG81]. Structure [BG70, Goo59c, GT77]. structures [Ked12]. Student [GS86b]. Studies [Goo79-29]. study [Goo89u]. subintervals [Goo85l]. Subjective [Goo62i, Goo69f, Goo86-27, Goo90n]. Subjectivism [Aga75a, Goo76b]. subjectivity [Goo85n]. Sufficient [Kha78, Goo93a]. sufficientivity [Goo94l]. Sufftctmcy [Goo85m]. Suggested [Goo70g]. sum [Goo71b, Goo72e, GBSS84]. Summary [OG74s]. Summation [Goo62a, Goo41b, Goo42, Goo44, Goo46b]. summing [Goo79o]. Sums [Goo94o, Goo61b, Goo62e, Goo79g, Goo93d]. superintelligence [Nor12]. Suppes [Hum94]. supplement [Goo81n]. Support [Goo87o, Goo87k]. surfaces [GM72, Ked12]. Surprise [Goo56e, Goo57b, Goo71d, Goo83w, Goo84y, Goo85i, Goo93c, Goo89s]. survey [Goo85w, Goo87f, OG74s]. Suspicious [Goo90w, Goo89e]. sweep [Goo76b]. symbols [OG74n]. Symmetric [CG80, CG81, Goo75a, Goo76d]. Symmetry [Goo94o, JG81]. Symposium [Ano53]. synthesis [Goo83-30]. System [Hal60]. Systèmes [Goo66b].

T [Goo61d, Goo73a, WG82]. Table [CG77a, CG77b, Goo85u, GG80a, Goo79m, Goo86s]. Tables [CG80, CG81, CG96, Goo56c, Goo63b, Goo76d, GM87, GC87, PG86, CG82, GC77, Goo78p, Goo79e, Goo79x, Goo79i, Goo80a, Goo81s, Goo81x, Goo81w, GL81, Goo83f, GL83, Goo83-34, Goo85q, GM89, GF89, Mag81]. Tail [PG76, Goo82p, Goo84g, Goo84h, Goo84m, Goo84n, GS86b, HG86, Goo83-32, Goo85c]. tail-area [Goo82p, Goo84g, Goo84m, Goo84n, GS86b, Goo85c]. Tail-Areas [PG76]. tails [Goo84h, Goo86c]. Tapia [Goo82a]. tasting [Goo90l, Goo92c]. tax [Goo81t]. tea [Goo92c]. Technical [BRCG55]. techniques [GL86]. technology [Goo94j]. Teleprinter [GMT^{+}12, RDF12]. Tendencies [Goo93a]. tendency [Goo94b, Goo94l]. Tentative [Goo93e]. terminology [Goo56d, Goo82i, Goo83h, Goo86e, Goo86t, Goo93b, Goo80l]. terms [OG74i]. Terrell [Goo90i]. Test [Goo55e, Goo57c, Goo67b, GG68a, Goo69f, GC87, Goo53b, Goo57g, GG67, GG68b, Goo74b, Goo78s, Goo78j, Goo81h, Goo82h, Goo84r, GS84, Goo85p, Goo86u, Goo89m, Goo94i, Goo96a, LS86]. testimony [Goo81y, Goo81-32]. Testing [Goo85u, Goo87q, Goo81n, Goo96a]. Tests [Goo55h, Goo58i, CG82, Goo53b, Goo84w, Goo86c, Goo87h, Goo90f, Goo94j]. Their [CG80, CG81, Gre22, Goo76d, GH82a, GS87]. Theorem [Cox60, FG53, Goo68d, Goo83x, Goo88e, Goo89g, Goo94h, Jos96a, Goo62g, Goo62h, Goo68g, Goo69c]. Theorems [Goo62b]. Théorie [Goo56a]. Theory [Ano53, FOO01, Goo59d, Goo64, Goo66a, Goo72b, Goo86w, Goo86x, Goo41a, Goo56d, Goo77b, GT77, Goo80w, Goo81-30, Goo82c, Goo86a, Goo87j, Goo88i, Goo89p, Goo89d, Goo89w, Goo89k, Goo62b, Goo56b, Goo62c]. therapy [Goo78r]. there [GL85]. Thinking [Goo83-35, Goo09]. Thirty [GBS08]. Thomas [Goo86b]. Thompson [Goo82a]. those [Goo87b]. Thoughts [Goo74e]. Threat [Goo78a]. Three [CMU^{+}53, CMU^{+}93, Oyv52]. thumb [Goo84x, Goo84h, Goo84i, Goo84w, Goo92b]. Ties [Goo75h, Goo80d, Goo85s]. Time [Goo61k, Goo59c]. Timms [RDF12]. Today [Goo72g]. tolerant [Goo84n]. tomographic [Goo82f]. too [Goo84l]. topic [Goo94f]. topic-independent [Goo94f]. topological [Goo41c]. Total [Goo67d, Goo81t]. Trans [FOO01]. Transform [CG77a, CG77b, Goo79-28, Goo81p, Goo82n, Goo87f, HJB85]. transformation [Goo00a, Goo01]. transforms [GR48, Goo50a, Goo71e]. Transition [Goo61k]. transitivity [Goo89j]. transmission [OG76]. trap [Goo90c]. Trapping [GLGH79]. Treatise [Goo52a]. treatments [Goo78k, Goo86p]. trees [Goo65b, Goo68c]. trials [Goo81g]. trick [GF89]. Trigonometrical [Goo93d]. trigonometry [Goo79t, Goo92e, Goo94c]. trinomial [Goo58f]. trouble [Goo83-32]. true [HG86]. truncated [GT78]. Truth [Goo58a]. Truthlikeness [Ros75]. trying [AG81]. Tunny [GMT^{+}12, RDF12]. tuples [Goo85s]. Turing [Goo92g, Bri92, Goo79-29, Goo84-28, Goo94h, Goo00b, vdV09]. Twentieth [MHR80]. twin [Per00]. Two [Goo71e, Goo85u, GM87, MG65, Goo78m, Goo81e, Goo84h, Goo86p, GM89, GHDB89, OG74r]. Two-by-Two [GM87, GM89]. two-tails [Goo84h]. Two-Way [Goo85u]. Type [BG76b, BG76a].

UK [Shi25]. ultraintelligent [Goo65e]. unausgegorener [MA66]. unblocked [Goo87p, Goo91d]. uncertain [Goo81y, Goo81-32]. underflows [Goo78e]. undermining [Goo89g]. uniform [OG76]. unimodal [LR77]. unit [HG85]. Unity [Goo61d]. universal [Goo83b]. universe [Goo72h]. University [Sav62]. unpaired [Goo84g, Goo84i]. unsolved [Goo79v]. untitled [Goo72c, Goo78b, Goo78c, Goo78d, Goo79a, Goo79b, Goo79c, Goo80b, Goo80c, Goo83a, Goo85a, Goo87a]. updated [Goo94l]. updating [GJH90]. upper [HG86]. upper-tail [HG86]. use [Goo69g, Goo79r, Goo87f, Goo90d]. used [Goo96a]. Using [PG86, Goo81p, Goo82n, Goo84v, Goo84l]. usual [Goo88c]. Utility [Goo60g, Goo68e, Goo71d, Goo83g].

V [Goo03b, Goo03c]. vague [Goo87l]. valid [Goo69b]. Valley [Ano09e]. value [Goo80a, Goo80k, Goo81s, Goo83c, Goo83o, Goo84z, Goo88b, Goo88g, Goo90c]. Values [Goo63a, Goo87q, Goo86q, GS87, Goo89l, Goo89s, Goo90k, Goo92c, Goo93c]. variability [Goo89q]. Variables [KG77, Goo60c, Goo79s]. Variance [KG77, Goo80s, Goo81u, GS85a, MG80]. Variates [Goo73a]. Varieties [Goo96d]. various [Goo80l]. vectors [Goo80m, GL84]. Verisimilitude [Aga75b]. versus [Goo82e, Goo84h, Goo84i, Goo93b, Goo94k]. verticles [GJH90]. Very [Goo86c]. via [GJ82]. Vinograde [Goo80o]. Virginia [Ano09e]. Visions [Oyv26, Oyv52]. visitations [Goo97a]. Viterbi [FOO01]. vocabulary [Goo78q, Goo82o]. Vol [Goo63a]. Volkswagens [Lou94]. Volume [Bri92, JL95]. Voss [Goo91c]. voting [GT77, Goo79o]. vs [Goo89y].

walk [FG53]. walks [Goo58f]. War [Goo66f, Goo79-29, Goo92g]. wartime [Ano09f]. Washburne [Goo63a]. Watson [GG90]. Way [Goo85u, Goo86l]. Weak [Goo89j]. Weibull [Goo96a]. Weighing [Goo50d]. Weight [Goo60g, Goo61o, Goo85w, OG74l, Goo80f, Goo81c, Goo81y, Goo81-32, Goo83t, Goo84k, Goo85m, Goo89b, OG74i, OG74f, OG74h, OG74k, OG74p, Goo68e, Goo89i]. Weighted [Goo55h, Goo63e]. Weighting [Dav51]. weights [Goo83p, Goo03d].

Wetherill [BLH^{+}60]. Wexelblat [SWT^{+}81]. Which [Goo95h, Sto93]. White [Goo67e, Goo68i]. Who [Goo85h]. Wilcoxon [Goo85p, Goo94i]. Wiley [Goo56b]. Wilfred [Goo71g]. Will [Goo71c, Goo81o]. Wishart [GJ82]. Wissenschaft [MA66]. withheld [Goo88a]. within [Goo81p, Goo82n]. without [AG81]. Wolfowitz [Goo62b]. Words [Gre22]. Work [DE86, Goo79y, Goo80t, Goo86y, Goo86z, Goo92g, Goo95g, Hey86, Sin86, Sol86, Goo79z, Goo79-29]. worked [vdV09]. Works [Bri92]. World [Goo92g, Goo79-29]. Wright [Goo52a, Goo80p]. Writing [Goo55b, Goo68h]. WWII [Smi19].

xi [Goo56b]. XXXVII [Goo79-29].

Yates [Goo84r, Goo84t, Goo67a]. year [Goo81o]. Years [GM96, GBS08, Goo90o]. York [Goo56b]. Yule [GL83].

Zellner [Goo76a, Goo75b]. Zero [Goo61c]. Zoroastrian [Sin08].

References

Anscombe:1969:DPJ

[ABH^{+}69] F. J. Anscombe, Irwin D. J.

Bross, H. O. Hartley, Rolf E. Bargmann, H. A. David, Marvin Zelen, R. L. Anderson, Miles Davis, and I. J. Good. Discussion of paper by I. J. Good, April 9, 1968. Journal of the American Statistical Association, 64(325):50–66, March 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2283719. See [Goo69f].

Adelberg:1984:PSS

[AC84] Arnold Adelberg and Douglas E. Cameron. Problems and solutions: Solutions of elementary problems: E2887. American Mathematical Monthly, 91(1):58, January 1984. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HSO^{+}81].

Anderson:1975:APS

[AF75] C. W. Anderson and Neal Felsinger. Advanced problems and solutions: Solutions: 5949. American Mathematical Monthly, 82(5):536–538, May 1975. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ehr74].

Arnold:1981:CHG

[AG81] J. C. Arnold and I. J. Good. C118. How to get a large R^{2} without really trying. Journal of Statistical Computation and Simulation, 14(1):69–71, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Agassi:1975:SID

[Aga75a] Joseph Agassi. Subjectivism: From infantile disease to chronic illness. Synthese, 30(1–2): 3–14, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/
10.1007/BF00485289.

Agassi:1975:VCD

[Aga75b] Joseph Agassi. Verisimilitude: Comment on David Miller. Synthese, 30(1–2): 199–204, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/
10.1007/BF00485306.  See

[Mil75a] and reply [Mil75b].

Anderson:1974:PSA

[And74] C. W. Anderson. Problems and solutions: Advanced problems: 5949. American Mathematical Monthly, 81(1):89–91, January 1974. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AF75, EB75, GK75, GSA75, KC75, MM75].

Anderson:1989:CSC

[And89] Oliver D. Anderson. C307. Serial correlation: the bending moment analogy. Journal of Statistical Computation and Simulation, 31(1):40–44, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90].

Anonymous:1953:RPS

[Ano53]  Anonymous, editor.  Report

of Proceedings, Symposium on Information Theory, London, September, 1950, number 1 in Transactions on Information Theory. Institute of Radio Engineers, London, UK, February 1953.

Anonymous:1981:CHP

[Ano81] Anonymous. Corrections: How Polish Mathematicians Deciphered the Enigma, 3(3) 232, Reviews: H. H. Goldstine: A History of Numerical Analysis, 3(3) 293. Annals of the History of Computing, 3(4): 407, October/December 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

                                        dlib.computer.org/an/books/
                                        an1981/pdf/a4400.pdf.  See

[Rej81, SWT^{+}81].

Anonymous:1990:CEC

[Ano90] Anonymous. C350. Errata to C202, 305, 307, 310, 311, and 334. Journal of Statistical Computation and Simulation, 34(2– 3):176, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo84p, Goo88f, And89, JK89, Goo89a, Goo89u].

Anonymous:2009:IJG

[Ano09a] Anonymous. Irving John Good. Virginia Tech Web site., April 10,

                                        2009. URL https://www.stat.
                                        vt.edu/about/History/Good.
                                        html.

Anonymous:2009:JGa

[Ano09b] Anonymous. Jack Good. Roanoke Times, ??(??):??, April 7, 2009.

Anonymous:2009:JGb

[Ano09c] Anonymous. Jack Good. Daily Telegraph, ??(??):??, April 10,

Anonymous:2009:OPI

[Ano09d] Anonymous. Obituaries/passings/ Irving John Good. Los Angeles Times, ??(??):??, April 13, 2009. URL https:/

/www.latimes.com/archives/
la-xpm-2009-apr-13-me-passings13.
s1-story.html.

Anonymous:2009:OIJ

[Ano09e] Anonymous. Obituary: Irving John Good, New River Valley, Virginia. Web site, April 2009. URL https:/

/www.legacy.com/obituaries/
name/irving-good-obituary?
pid=125834454;   https://www.
legacy.com/us/obituaries/roanoke/
name/irving-good-obituary?
id=28686552.

Anonymous:2009:OPJ

[Ano09f] Anonymous. Obituary: Professor Jack Good: Mathematician and wartime codebreaker. The Times [London, UK], ??(??):??, April 16, 2009.

Anonymous:2009:PJG

[Ano09g] Anonymous. Professor Jack Good. The Telegraph, ??(??): ??, April ??, 2009. URL https:

//www.telegraph.co.uk/news/
obituaries/military-obituaries/
special-forces-obituaries/
5132599/Professor-Jack-Good.
html.

Anonymous:2013:IJG

[Ano13] Anonymous. Irving J. Good papers, 1925–2004. Special Collections and University Archives, University Libraries (0434), 560 Drillfield Drive, Newman Library, Virginia Tech, Blacksburg, Virginia 24061, USA,

2013. URL https://aspace.
lib.vt.edu/repositories/2/
resources/1378.

Anonymous:2021:LJG

[Ano21] Anonymous. Legends: I. J. Good. Virginia Tech Science Magazine, Fal 2021. URL

https://www.science.vt.edu/
magazine/stories/fall-2021/
legends.html.

Anonymous:20xx:CPI

[Anoxx]  Anonymous.  Collected Pub-

lications of Irving John Good (1916–2009), computer scientist. ????, ????, 20xx. ???? pp.

URL https://solo.bodleian.
ox.ac.uk/discovery/fulldisplay?
docid=alma990204606930107026&
context=L&vid=44OXF_INST:SOLO&
lang=en&search_scope=MyInst_
and_CI&adaptor=Local%20Search%
20Engine&tab=Everything&query=
creator,contains,irving%20john%
20good,AND&query=any,contains,
thesis,AND&mode=simple&offset=
0.

Banks:1996:CJG

[Ban96] David L. Banks. A conversation with I. J. Good. Statistical Science, 11(1):1–19, 1996. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Barnard:1990:C

[Bar90]  George A. Barnard.  Com-

ment. Journal of Statistical Planning and Inference, 25 (3):437–439, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/037837589090086A.

See reply [Goo90t].

Blough:1989:LE

[BBC^{+}89] David K. Blough, Daniel G.

Brooks, Steven Carroll, William A. Verdini, John T. Kent, S. R. Searle, Ravindra Khattree, Andrew R. Solow, Aaron A. Aardvark, I. J. Good, Gordon H. Fick, Hans Riedwyl, George D. Swanson, and Ralph B. D’Agostino. Letters to the Editor. The American Statistician, 43(3): 186–188, August 1989. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2685072.

Berry:1972:LE

[BDR^{+}72] Donald A. Berry, W. Edwards Deming, A. C. Rosander, Robert A. Agnew, C. Philip Cox, L. R. LaMotte, Warren Dent, I. J. Good, Franklin H. Silverman, Hans K. Ury, Marian B. Ury, John Havas, S. R. Searle, D. S. Robson, Lawrence L. Kupper, Stanley W. Nash, A. L. Wilson, S. L. George, M. M. Desu, Ralph E. Beals, and Herbert A. David. Letters to the Editor. The American Statistician, 26 (1):47–54, February 1972. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2682851.

Bernhart:1973:PSAa

[Ber73] Frank Bernhart. Problems and solutions: Advanced problems: 5895. American Mathematical Monthly, 80(2):208–209, February 1973. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BK75, Goo75i, RC74, SB74, SH74].

Beran:1970:RSI

[BG70] R. J. Beran and I. J. Good. Reviews: The Structure of Inference, by D. A. S. Fraser. American Mathematical Monthly, 77 (5):538–541, May 1970. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bruckman:1976:GSM

[BG76a] P. S. Bruckman and I. J. Good. A generalization of a series of de Morgan, with applications of Fibonacci type. Fibonacci Quarterly, 14(3):193–196, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:/

                                        /www.fq.math.ca/Scanned/14-
                                        3/bruckman.pdf.

Bruckman:1976:GSD

[BG76b] P. S. Bruckman and I. J. Good. A generalization of a series of DeMorgan with applications of Fibonacci type. Fibonacci Quarterly, 14(3):193–195, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:/

                                        /www.fq.math.ca/Scanned/14-
                                        3/bruckman.pdf.

Bacchus:1994:RBB

[BGHK94] Fahiem Bacchus, Adam Grove, Joseph Y. Halpern, and Daphne Koller. A response to “Believing on the basis of the evidence”. Computational Intelligence, 10 (1):21–25, February 1994. CO-DEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic).

Ben-Haim:1996:CMR

[BH96] Yakov Ben-Haim. C444. Must reliability be probabilistic? Journal of Statistical Computation and Simulation, 55(3):263–265, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See reply [Goo96b].

Birnbaum:1962:FSI

[Bir62]  Allan Birnbaum.  On the

foundations of statistical inference. Journal of the American Statistical Association, 57 (298):269–306, June 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2281640. See discussion

[SBC^{+}62].

Bernhart:1975:PSS

[BK75] Frank Bernhart and D. J. Kleitman. Problems and solutions: Solutions of advanced problems: 5895. American Mathematical Monthly, 82(5):531–532, May 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ber73].

Behrmann:1967:PSA

[BKM^{+}67] Adriano Behrmann, M. S. Klamkin, W. A. McWorter, W. G. Spohn, Jr., C. M. Petty, W. E. Johnson, M. S. Klamkin, Benjamin Volk, M. V. Subbarao, S. W. Williams, and Charles Green. Problems and solutions: Advanced problems: 5460–5469. American Mathematical Monthly, 74(2):206–208, February 1967. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BvL68, GC68b, KS68, MTD68, PJR^{+}68, SAG68, SG68, VM68, WF68].

Barnard:1960:DPM

[BLH^{+}60] G. A. Barnard, D. V. Lindley, B. Hill, F. J. Anscombe, I. J. Good, and G. Horsnell. Discussion of the papers of Messrs. Hald, Wetherill and Cox. Technometrics, 2(3):361– 372, August 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1266250. See [Hal60,

Wet60, Cox60].

Brenner:1955:TNS

[BRCG55] J. L. Brenner, G. W. Reitwiesner, C. W. Clenshaw, and I. J. Good. Technical notes and short papers. Mathematical Tables and Other Aids to Computation, 9 (51):117–121, July 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Bross:1991:LE

[BRG^{+}91] Irwin D. Bross, Marion R.

Reynolds, Jr., Timothy G. Gregoire, Stanley S. Schor, Beth Dawson-Saunders, Paul K. Jones, Steven J. Verhulst, Burt Holland, Per Lea, Dave Saville, Eugene Garfield, I. J. Good, Daniel Gianola, Jean-Louis Foulley, Laurence A. Baxter, Sherman Chottiner, Paul A. Herzberg, Judith D. Singer, John B. Willett, Ronald A. Thisted, Chap T. Le, Jean D. Gibbons, B. M. Pőtscher, C. M. Hurvich, and C. L. Tsai. Letters to the editor. The American Statistician, 45(2):163–172, May 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2684396.

Britton:1992:CWT

[Bri92] J. L. Britton, editor. Collected Works of A. M. Turing: Volume 1: Pure Mathematics. North-Hol-land, Amsterdam, The Netherlands, 1992. ISBN 0-444-88059-3. xxii + 287 pp. LCCN QA3 .T87 1992. Edited and with an introduction and postscript by J. L. Britton and Irving John Good. With a preface by P. N. Furbank.

Behrmann:1968:PSS

[BvL68] Adriano Behrmann and J. H. van Lint. Problems and solutions: Solutions of advanced problems: 5460. American Mathematical Monthly, 75(2):202, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Coxeter:1968:PSSa

[CB68] H. S. M. Coxeter and Leon Bankoff. Problems and solutions: Solutions of elementary problems: E1920. American Mathematical Monthly, 75(2):192, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Cornell:1970:MPC

[CG70] John A. Cornell and I. J. Good. The mixture problem for categorized components. Journal of the American Statistical Association, 65(329):339–355, March 1970. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL https://www.
                                        jstor.org/stable/2283597.

Cornell:1971:CMP

[CG71] John A. Cornell and I. J. Good. Corrigenda: “The Mixture Problem for Categorized Components. Journal of the American Statistical Association, 66 (333):229, March 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2284890.

Conolly:1977:ETD

[CG77a] Brian Conolly and I. J. Good. Errata: “A Table of Discrete Fourier Transform Pairs” [SIAM J. Appl. Math. 32 (1977), no. 4, 810– 822]. SIAM Journal on Applied Mathematics, 33(3):534, November 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). See [CG77b].

Conolly:1977:TDF

[CG77b] Brian Conolly and I. J. Good. A table of discrete Fourier transform pairs. SIAM Journal on Applied Mathematics, 32(4): 810–822, June 1977. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic). See errata [CG77a].

Crook:1980:ASD

[CG80] J. F. Crook and I. J. Good.

On the application of symmetric Dirichlet distributions and their mixtures to contingency tables, Part II. Annals of Statistics, 8(6):1198–1218, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.
aos/1176345194. See [CG81].

Crook:1981:CAS

[CG81] J. F. Crook and I. J. Good. Corrections to “On the Application of Symmetric Dirichlet Distributions and Their Mixtures to Contingency Tables, Part II”. Annals of Statistics, 9(5):1133, September 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.
aos/1176345597. See correction

[CG80].

Crook:1982:PST

[CG82] James Flinn Crook and Irving John Good. The powers and strengths of tests for multinomials and contingency tables. Journal of the American Statistical Association, 77(380): 793–802, December 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2287310.

Crook:1996:SAA

[CG96] J. F. Crook and I. J. Good. Statistical algorithms: Algorithm AS 301: The computation of a Bayes factor against independence in contingency tables. Applied Statistics, 45(2):255–265, 1996. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.
                                        cmu.edu/apstat/301.

Coxeter:1968:PSSb

[CGGR68] H. S. M. Coxeter, M. G. Greening, Ron Graham, and Jurg Ratz. Problems and solutions: Solutions of elementary problems: E1921. American Mathematical Monthly, 75(2):192–193, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Cano:1978:PSS

[CGI78] J. Cano, A. Gruebler, and Robert B. Israel. Problems and solutions: Solutions of advanced problems: 6133. American Mathematical Monthly, 85 (9):771, November 1978. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77].

Cherry:1953:DPT

[CMU^{+}53] E. C. Cherry, S. H. Moss, ?. Uttley, I. J. Good, W. Lawrence, W. P. Anderson, and Claude E. Shannon. Discussion of preceding three papers. In Anonymous [Ano53], pages 169–174.

Cherry:1993:DPT

[CMU^{+}93] E. C. Cherry, S. H. Moss, ?. Uttley, I. J. Good, W. Lawrence, W. P. Anderson, and Claude E. Shannon. Discussion of preceding three papers. In Sloane and Wyner [SW93], pages 184– 189. ISBN 0-470-54424-4, 0-7803-0434-9. LCCN TK5101 .S448 1993. Published in IRE Transactions Information Theory, No. 1, Feb. 1950. Originally published in [CMU^{+}53].

Cox:1960:SSA

[Cox60] D. R. Cox. Serial sampling acceptance schemes derived from Bayes’s theorem. Technometrics, 2(3):353–360, August 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1266249. See discussion

[BLH^{+}60].

Canfield:1977:LE

[CTL^{+}77] Ronald V. Canfield, David L.

Turner, Eugene Lehman, I. J. Good, Gottfried E. Noether, Benjamin J. Tepping, Bernard Most, Oliver D. Anderson, David A. Swanson, Roy H. Wampler, David Rothman, Russell V. Lenth, and Stephen Portnoy. Letters to the Editor. The American Statistician, 31(1): 50–54, February 1977. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2682982.

David:1951:BRB

[Dav51] F. N. David. Book review: Probability and the Weighting of Evidence, by I. J. Good. Biometrika, 38(3/4):485, December 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2332601.

Diaconis:1986:SSA

[DE86] Persi Diaconis and Eduardo Engel. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):171–174, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013691. See [Goo86y,

Sol86, Hey86, Sin86] and rejoinder [Goo86z].

Dawson:1957:EMP

[DG57] Reed Dawson and I. J. Good. Exact Markov probabilities from oriented linear graphs. Annals of Mathematical Statistics, 28(4): 946–956, December 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.
org/euclid.aoms/1177706795.

Dickey:1975:LE

[DGK^{+}75] James Dickey, N. T. Gridgeman, M. C. S. Kingsley, I. J. Good, James E. Carlson, Daniel Gianola, Michael H. Kutner, and Steve Selvin. Letters to the Editor. The American Statistician, 29(3):131–134, August 1975. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2683443.

Deming:1970:LEb

[DMK^{+}70] W. Edwards Deming, Donald V. Mathusz, Scott A. Krane, Keith Kussmaul, Charles De-Witt Roberts, I. J. Good, Haynes Phillips, Charles J. Kowalski, Kenneth Mullen, Samuel H. Plotkin, Jason Millman, Paul Cieslak, S. Chatterjee, Nathan Mantel, Alan B. Forsythe, Joseph M. Moser, Ayala Cohen, Evelio Fabbroni, and Victor Chew. Letters to the Editor. The American Statistician, 24(4):46–53, October 1970. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2681316.

Doyle:1994:IAC

[Doy94] Jon Doyle. Inference and acceptance: Comments on Kyburg’s “Believing on the basis of the evidence”. Computational Intelligence, 10(1):46–48, February 1994. CODEN COMIE6.

ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Deming:1970:LEa

[DTB^{+}70] W. Edwards Deming, H. Tamura, F. W. Bennett, Jr., I. H. LaValle, Eugene Dayhoff, William Knight, I. J. Good, L. R. Shenton, C. E. Stephenson, John D. Williams, Sudhir N. Dalal, Min-Te Chao, and M. N. Das. Letters to the Editor. The American Statistician, 24(2):35–39, April 1970. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2681931.

Ehrenfeucht:1975:APS

[EB75] Andrzej Ehrenfeucht and R. G. Burns. Advanced problems and solutions: Solutions: 5946. American Mathematical Monthly, 82 (3):310, March 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ehr74].

Esanu:1977:PSA

[ECG^{+}77] Mihai Esanu, J. Cano, A. Gruebler, Barbara Osofsky, Paul Erdős, H. L. Montgomery, and I. J. Good. Problems and solutions: Advanced problems: 6132–6137. American Mathematical Monthly, 84(2):140–141, February 1977. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CGI78, EP78, GR78, GB80, Goo81-31, MR78, MW80, Oso78].

Ehrenfeucht:1974:PSA

[Ehr74] Andrzej Ehrenfeucht. Problems and solutions: Advanced problems: 5946. American Mathematical Monthly, 81(1):89–91, January 1974. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AF75, EB75, GK75, GSA75, KC75, MM75].

Esanu:1978:APS

[EP78] Mihai Esanu and Purdue– Calumet Coffee Club. Advanced problems and solutions: Solutions: 6132. American Mathematical Monthly, 85(8):690, October 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77].

Erdos:1976:PSS

[ES76] Paul Erdős and Saul Stahl. Problems and solutions: Solutions of advanced problems: 5953. American Mathematical Monthly, 83 (7):574, August/September 1976. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gil74].

Foster:1953:GPR

[FG53] F. G. Foster and I. J. Good. On a generalization of Pólya’s randomwalk theorem. The Quarterly Journal of Mathematics. Oxford. Second Series, 4:120–126, 1953. ISSN 0033-5606,1464-3847.

Fienberg:2008:JGA

[Fie08] S. E. Fienberg. I. J. Good— an appreciation. In Good et al.

[GBS08], page ?? ISBN 0-89263-006-X (hardcover), 0-89263-007-8 (perfect). LCCN QA99 .G66 2008.

Frenger:2001:CAR

[FOO01] Pål K. Frenger, Pål Orten, and Tony Ottosson. Comments and additions to recent papers on new convolutional codes: “Rational rate punctured convolutional codes for soft-decision Viterbi decoding” [IEEE Trans. Inform. Theory 43 (1997), no. 4, 1305– 1313] by I. E. Bocharova and B. D. Kudryashov and “Some extended results on the search for good convolutional codes” [IEEE Trans. Inform. Theory 43 (1997), no. 5, 1682–1697; MR1476810 (98e:94035)] by J.-J. Chang, D.-J. Hwang and M.-C. Lin. IEEE Transactions on Information Theory, 47(3):1199–1201, 2001. CODEN IETTAW. ISSN 0018-9448,1557-9654.

Good:1980:PSS

[GB80] I. J. Good and P. T. Bateman. Problems and solutions: Solutions of advanced problems: 6137. American Mathematical Monthly, 87(6):495–496, June/ July 1980. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77, GR78, Goo81-31].

Good:1982:PSS

[GB82] I. J. Good and Anders Bager. Problems and solutions: Solutions of elementary problems: E2889. American Mathematical Monthly, 89(9):701–702, November 1982. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HSO^{+}81].

Good:1969:LE

[GBC69] I. J. Good, Richard M. Brugger, and W. P. Cooke. Letters to the Editor. The American Statistician, 23(1):35, February 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL https://www.
jstor.org/stable/2681692.

Gilles:1975:APS

[GBMW75] J. Gilles, Bruce C. Berndt, H. L.

Montgomery, and P. J. Weinberger. Advanced problems and solutions: Solutions: 5952. American Mathematical Monthly, 82(6):679–681, June/July 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gil74, GRM76].

Good:2008:GBT

[GBS08] Irving John Good, David Banks, and Eric P. Smith, editors. The Good Book: Thirty Years of Comments, Conjectures, and Conclusions by I. J. Good. Rice University Press, Houston, TX, USA, 2008. ISBN 0-89263-006-X (hardcover), 0-89263-007-8 (perfect). ???? pp. LCCN QA99 .G66 2008.

Good:1984:CPS

[GBSS84] I. J. Good, K. O. Bowman, L. R. Shenton, and G. Szekeres. C212. A Poisson sum up to the mean and a Ramanujan problem.

Journal of Statistical Computation and Simulation, 20(2):167– 173, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1968:RHP

[GC68a] I. J. Good and R. F. Churchhouse. The Riemann Hypothesis and pseudorandom features of the Möbius sequence. Mathematics of Computation, 22(104): 857–861, October 1968. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

                                        URL https://www.jstor.org/
                                        stable/pdfplus/2004584.pdf.

Green:1968:PSS

[GC68b] Charles Green and L. D. Crowson. Problems and solutions: Solutions of advanced problems: 5469. American Mathematical Monthly, 75(3):306, March 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Good:1974:BNB

[GC74] Irving John Good and James Flinn Crook. The Bayes/non-Bayes compromise and the multinomial distribution. Journal of the American Statistical Association, 69(347):711–720, September 1974. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL https://www.

                                        jstor.org/stable/2286006.

Good:1977:EAG

[GC77] I. J. Good and J. F. Crook. The enumeration of arrays and a generalization related to contingency tables. Discrete Mathematics, 19(1):23–45, 1977. CODEN DSMHA4. ISSN 0012-365X,1872-681X.

Good:1978:CED

[GC78] I. J. Good and J. F. Crook.

Corrigenda: Exact Distributions for χ^{2} and for the Likelihood-Ratio Statistic for the Equiprobable Multinomial Distribution. Journal of the American Statistical Association, 73(364): 900, December 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2286319.

Good:1987:RSM

[GC87] I. J. Good and J. F. Crook. The robustness and sensitivity of the mixed-Dirichlet Bayesian test for “Independence” in contingency tables. Annals of Statistics, 15 (2):670–693, June 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.
org/euclid.aos/1176350368.

Good:1990:CGR

[GC90] I. J. Good and I. N. Chengalur.

C340. A generalization of Rissanen’s distribution. Journal of Statistical Computation and Simulation, 34(2–3):155–158, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1958:PCR

[GD58] I. J. Good and K. Caj Doog. A paradox concerning rate of information. Information and Control, 1:113–126, 1958. CODEN IFCNA4. ISSN 0019-9958,1878-2981.

Good:1959:PCR

[GD59] I. J. Good and K. Caj Doog. A paradox concerning rate of information: corrections and additions. Information and Control, 2:195–197, 1959. CODEN IFCNA4. ISSN 0019-9958,1878-2981.

Good:1973:IJG

[GE73] I. J. Good and Christopher Riche Evans. An interview with I. J. Good. Harper & Row, New York, NY, USA, 1973. ???? pp. Audio recording.

Good:1989:CFP

[GF89] I. J. Good and R. V. Foutz. C329. The “Fisher–Poisson trick” extended to contingency tables. Journal of Statistical Computation and Simulation, 32(1–2):76– 81, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Goldsmith:1965:LE

[GFG^{+}65] Charles H. Goldsmith, Joseph L. Fleiss, A. F. Goodman, Gene Laska, Henry Tucker, I. J. Good, Paul Eldridge, Michael F. Capobianco, N. F. Laubscher, and Ira Marshak. Letters to the Editor. The American Statistician, 19(3):42–44, June 1965. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2681423.

Good:1977:LE

[GFM^{+}77] I. J. Good, Stephen E. Fienberg, James M. Malone, II, Maurice C. Bryson, Hugo C. Hamaker, R. M. Kozelka, Herbert F. Spirer, D. V. Lindley, and L. D. Phillips. Letters to the Editor. The American Statistician, 31(4):180–183, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL https://www.
jstor.org/stable/2683545.

Good:1967:GST

[GG67] I. J. Good and T. N. Gover. The generalized serial test and the binary expansion of \sqrt{2}. Journal of the Royal Statistical Society. Series A (General), 130 (1):102–107, ???? 1967. CO-DEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2344040. See cor-

rigendum [GG68a] and remark [GG68b].

Good:1968:CGS

[GG68a] I. J. Good and T. N. Gover. Corrigendum: The generalized serial test and the binary expansion of \sqrt{2}. Journal of the Royal Statistical Society. Series A (General), 131(3):434, ???? 1968. CO-DEN JSSAEF. ISSN 0035-9238.

URL https://www.jstor.org/
stable/2343530. See [GG67].

Good:1968:GST

[GG68b] I. J. Good and T. N. Gover. The generalized serial test and the binary expansion of \sqrt{2}. Journal of the Royal Statistical Society. Series A (General), 131(??):434, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. See [GG67].

Good:1970:CMN

[GG70] I. J. Good and R. A. Gaskins. The centroid method of numerical integration. Numerische Mathematik, 16:343–359, 1970. CO-DEN NUMMA7. ISSN 0029-599X,0945-3245.

Good:1971:CMN

[GG71a] I. J. Good and R. A. Gaskins. The centroid method of numerical integration. Numerische Mathematik, 16(4):343–359, January 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Good:1971:NRP

[GG71b] I. J. Good and R. A. Gaskins. Nonparametric roughness penalties for probability densities. Biometrika, 58(2):255– 277, August 1971. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2334515.

Good:1971:SRS

[GG71c] I. J. Good and R. A. Gaskins. Some relationships satisfied by additive and multiplicative congruential sequences, with implications for pseudorandom number generation. In A. O. L. (Arthur O. L.) Atkin and B. J. (Bryan John) Birch, editors, Computers in number theory: proceedings of the Science Research Council Atlas Symposium no. 2 held at Oxford, from 18–23 August, 1969, pages 125– 136. Academic Press, New York, NY, USA, 1971. ISBN 0-12-065750-3. LCCN QA241 .S35 1969.

Good:1972:GNE

[GG72] I. J. Good and R. A. Gaskins.

Global nonparametric estimation of probability densities. Virginia Journal of Science, 23:171–193, 1972. ISSN 0042-658X.

Gaskins:1977:CCD

[GG77] R. A. Gaskins and I. J. Good.

C4. Compenis on the distribution of the likelihood ratio statistic for the equiprobable multinomial. Journal of Statistical Computation and Simulation, 6(2): 150–152, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CBO

[GG78] I. J. Good and R. A. Gaskins. C19. Bayesian oedemology. Journal of Statistical Computation and Simulation, 7(3–4):299– 301, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gaskins:1980:CCB

[GG80a] R. A. Gaskins and I. J. Good.

C62. The covariance between the chi-squareds for the margins of a contingency table. Journal of Statistical Computation and Simulation, 10(2):161–162, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:DEBa

[GG80b] I. J. Good and R. A. Gaskins. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data. Journal of the American Statistical Association, 75 (369):42–56, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2287377.

Good:1980:DEBb

[GG80c] I. J. Good and R. A. Gaskins. Density estimation and bumphunting by the penalized likelihood method exemplified by scattering and meteorite data: Rejoinder. Journal of the American Statistical Association, 75 (369):70–73, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2287385.

Good:1990:CSC

[GG90] I. J. Good and N. L. Gilinsky. C346. A simple comment concerning the Galton–Watson– Bienayme process. Journal of Statistical Computation and Simulation, 34(2–3):165–166, 1990.

CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gartside:2001:LE

[GGF^{+}01] Peter S. Gartside, Joseph L.

Gastwirth, Boris Freidlin, W. Fred Webber, Walter A. Rosenkrantz, Sander Greenland, Alan Agresti, Herman Callaert, Reinhard Bergmann, and I. J. Good. Letters to the Editor. The American Statistician, 55(2):171–174, May 2001. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL https://www.

jstor.org/stable/2686010.

Good:1970:EDL

[GGM70] I. J. Good, T. N. Gover, and G. J. Mitchell. Exact distributions for χ^{2} and for the likelihood-ratio statistic for the equiprobable multinomial distribution. Journal of the American Statistical Association, 65 (329):267–283, March 1970. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2283591. See corrigenda

[GGM71].

Good:1971:CED

[GGM71] I. J. Good, T. N. Gover, and G. J. Mitchell. Corrigenda: Exact Distributions for χ^{2} and for the Likelihood-Ratio Statistic for the Equiprobable Multinomial Distribution. Journal of the American Statistical Association, 66 (333):229, March 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2284891. See [GGM70].

Good:1982:CFR

[GH82a] I. J. Good and Golde I. Holtzman. C141. On s-fold repeats and their additive and multiplicative properties. Journal of Statistical Computation and Simulation, 16(1): 66–69, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CMN

[GH82b] I. J. Good and Golde I. Hotzman. C142. The moments of the number of s-fold repeats. Journal of Statistical Computation and Simulation, 16(1):69–75, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CED

[GH83a] I. J. Good and C. I. Holtzman. C179. The exact distribution of χ^{2} for a non-equiprobable multinomial. Journal of Statistical Computation and Simulation, 18(2–3):238–239, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCCb

[GH83b] I. J. Good and C. I. Holtzman. C180. Conjectures concerning the cumulants of the multivariate lognormal distribution. Journal of Statistical Computation and Simulation, 18(2–3):239–241, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CDH

[GHDB89] I. J. Good, G. I. Holtzman, M. L.

Deaton, and L. H. Bernstein. C328. Diagnosis of heart attack from two enzyme measurements by means of bivariate probability density estimation: statistical details. Journal of Statistical Computation and Simulation, 32(1–2): 68–76, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1969:EPE

[GHJT69] I. J. Good, I. Hacking, R. C. Jeffrey, and H. Törnebohm. The estimation of probabilities: an essay on modern Bayesian methods. Synthese, 16(2):234–244, ???? 1969. ISSN 0039-7857 (print), 1573-0964 (electronic).

Gilles:1974:PSAa

[Gil74] J. Gilles. Problems and solutions:

Advanced problems: 5952. American Mathematical Monthly, 81 (2):175–176, February 1974. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ES76, GBMW75, GRM76, HK75, MC75].

Gillies:1986:DPM

[Gil86]  Donald Gillies.  In defense

of the Popper–Miller argument. Philosophy of Science, 53(1): 110–113, March 1986. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.jstor.org/
stable/187924. See response

[Goo87o].

Good:1982:CDN

[GJ82] I. J. Good and D. R. Jensen. C131. Determining the noncentral Wishart characteristic function via quadratic expressions. Journal of Statistical Computation and Simulation, 15(1):85– 87, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CNM

[GJH90] I. J. Good, D. R. Jensen, and Y. V. Hui. C345. Numerology for the masses of elementary verticles [sic]: an updating. Journal of Statistical Computation and Simulation, 34(2–3):164–165, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Gilles:1975:PSS

[GK75] J. Gilles and Torleiv Klove. Problems and solutions: Solutions of advanced problems: 5951. American Mathematical Monthly, 82(6): 678–679, June/July 1975. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ehr74].

Good:1981:PEC

[GL81] I. J. Good and B. C. Lewis. Probability estimation for 2×s contingency tables and predictive criteria [and discussion]. Bulletin de l’Institut International de Statistique, 49(1):97–113, 1981. Proceedings of the 43rd session of the International Statistical Institute, Vol. 1 (Buenos Aires, 1981).

Good:1983:CRB

[GL83] I. J. Good and Byron C. Lewis.

C147. The relationship between an informational measure of association and a generalization of one of Yule’s measures for 2 × s tables, and the relevance of generalized determinants and permanents. Journal of Statistical Computation and Simulation, 17(1): 63–66, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CGP

[GL84] I. J. Good and B. C. Lewis. C185.

Generation of pairs of related probability vectors. Journal of Statistical Computation and Simulation, 19(1):82–85, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:MLM

[GL85] I. J. Good and B. C. Lewis. A maximum-likelihood/minimum-discrepancy method for estimation of multinomial probabilities when there are related multinomials. Journal of Statistical Computation and Simulation, 21(3–4): 227–240, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CBT

[GL86] I. J. Good and B. C. Lewis.

The choice between techniques for choosing between methods of estimation. Journal of Statistical Computation and Simulation, 25(1–2):1–8, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CFN

[GL88] I. J. Good and Byron C. Lewis. C304. A further note on probability estimation for pairs of related multinomials. Journal of Statistical Computation and Simulation, 29(4):353–357, 1988. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:PER

[GLGH79] I. J. Good, Byron C. Lewis, Raymond A. Gaskins, and L. W. Howell. Population estimation by the removal method assuming proportional trapping. Biometrika, 66(3):485– 494, December 1979. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2335167.

Good:1972:SCS

[GM72] I. J. Good and Lawrence S. Mayer. On surfaces of constant societal loss in a model of social choice. J. Mathematical Sociology, 2:209–220, 1972. Collection of articles on social choice.

Good:1984:PME

[GM84] I. J. Good and Alan F. McMichael. A pragmatic modification of explicativity for the acceptance of hypotheses. Philosophy of Science, 51(1):120–127, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://

www.jstor.org/stable/187734.

Good:1987:AGT

[GM87] I. J. Good and Y. Mittal.

The amalgamation and geometry of two-by-two contingency tables. Annals of Statistics, 15 (2):694–711, June 1987. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.
org/euclid.aos/1176350369.

Good:1989:AAG

[GM89] I. J. Good and Y. Mittal.

Addendum: “The amalgamation and geometry of two-by-two contingency tables” [Ann. Statist. 15 (1987), no. 2, 694– 711; MR0888434 (88f:62081)]. Annals of Statistics, 17(2):947, June 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.
aos/1176347155.

Good:1996:CQD

[GM96] I. J. Good and Donald Michie.

Comments, queries, and debate: 50 years after breaking the codes: a disclaimer. IEEE Annals of the History of Computing, 18(4):65–66, October/December 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL https:/

                                        /ieeexplore.ieee.org/iel4/
                                        85/11673/00539920.pdf.

Good:1962:SSA

[GMM62] Irving John Good, Alan James Mayne, and John Maynard Smith. The Scientist Speculates. An Anthology of Partly-Baked Ideas. Heinemann and Basic Books, London, UK and New York, NY, USA, 1962. xvii + 413 pp. LCCN ????

Good:2012:BTC

[GMT^{+}12] Irving John Good, Donald Michie, G. (Geoffrey) Timms, James A. Reeds, Whitfield Diffie, and Judith Veronica Field, editors. Breaking Teleprinter Ciphers at Bletchley Park: General Report on Tunny with Emphasis on Statistical Methods (1945). Wiley, New York, NY, USA, 2012. ISBN 0-470-46589-1 (hardcover). cxi + 673 pp. LCCN D810.C88 G66 2015.

Good:1985:LE

[GN85] I. J. Good and Gottfried E. Noether. Letters to the Editor. The American Statistician, 39(1):81, February 1985. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2683919.

Goldszmidt:1994:HBK

[Gol94]  Moisés Goldszmidt.  Hedges,

background knowledge, and evidence: a reply to Kyburg’s “Believing on the basis of the evidence”. Computational Intelligence, 10(1):53–56, February 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Good:1940:ALM

[Goo40] I. J. Good. The approximate local monotony of measurable functions. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 36:9–13, 1940. CODEN PCPSA4. ISSN 0008-1981.

Good:1941:FDT

[Goo41a] I. J. Good. The fractional dimensional theory of continued fractions. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 37:199–228, 1941. CODEN PCPSA4. ISSN 0008-1981.

Good:1941:NSC

[Goo41b] I. J. Good. Note on the summation of a classical divergent series. Journal of the London Mathematical Society, 16:180–182, 1941. CODEN JLMSAK. ISSN 0024-6107,1469-7750.

Good:1941:TCP

[Goo41c] Irving John Good. The topological concept of partial dimension based on the ideas of Henri Lebesgue. Ph.D. thesis, Faculty of Mathematics, University of Cambridge, Cambridge, UK, 1941.

URL https://www.proquest.
com/pqdtglobal/docview/301212994/

.

Good:1942:SRB

[Goo42] I. J. Good. Some relations between certain methods of summation of infinite series. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 38:144–165, 1942. CODEN PCPSA4. ISSN 0008-1981.

Good:1944:RMM

[Goo44] I. J. Good. On the regularity of moment methods of summation. Journal of the London Mathematical Society, 19:141–143, 1944. CODEN JLMSAK. ISSN 0024-6107,1469-7750.

Good:1946:NRD

[Goo46a] I. J. Good. Normal recurring decimals. Journal of the London Mathematical Society, 21: 167–169, 1946. CODEN JLM-SAK. ISSN 0024-6107,1469-7750.

Good:1946:RGM

[Goo46b] I. J. Good. On the regularity of a general method of summation. Journal of the London Mathematical Society, 21:110–118, 1946. CODEN JLMSAK. ISSN 0024-6107,1469-7750.

Good:1947:NPD

[Goo47] I. J. Good. A note on positive determinants. Journal of the London Mathematical Society, 22: 92–95, 1947. CODEN JLMSAK. ISSN 0024-6107,1469-7750.

Good:1949:NIC

[Goo49] I. J. Good. The number of individuals in a cascade process. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 45:360– 363, 1949. CODEN PCPSA4. ISSN 0008-1981.

Good:1950:BIT

[Goo50a] I. J. Good. Bounded integral transforms. II. Quart. J. Math. Oxford Ser. (2), 1:185–190, 1950. ISSN 0033-5606,1464-3847.

Good:1950:ICM

[Goo50b] I. J. Good. On the inversion of circulant matrices. Biometrika, 37(1/2):185–186, June 1950. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2332165.

Good:1950:PLI

[Goo50c] I. J. Good. A proof of Liapounoff’s inequality. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 46:353, 1950. CO-DEN PCPSA4. ISSN 0008-1981.

Good:1950:PWE

[Goo50d] Irving John Good. Probability and the Weighing of Evidence. Charles Griffin & Co., Ltd. and Hafner Publishing Co., London, UK and New York, NY, USA, 1950. viii + 119 pp. LCCN QA273 .G65.

Good:1951:BRB

[Goo51a] I. J. Good. Book review: Calculating Instruments and Machines, by D. R. Hartree. Journal of the Royal Statistical Society. Series A (General), 114(1):106–107, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL https://www.

jstor.org/stable/2980914.

Good:1951:BRC

[Goo51b] I. J. Good. Book review: Calculating Instruments and Machines, by D. R. Hartree. Journal of the Royal Statistical Society. Series A (General), 114(1):106–107, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL https://www.

jstor.org/stable/2980914.

Good:1951:RMF

[Goo51c] I. J. Good. Random motion on a finite Abelian group. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 47:756–762, 1951. CODEN PCPSA4. ISSN 0008-1981.

Good:1952:BRB

[Goo52a] I. J. Good. Book review: A Treatise on Induction and Probability, by G. H. von Wright. Journal of the Royal Statistical Society. Series A (General), 115 (2):283–285, ???? 1952. CO-DEN JSSAEF. ISSN 0035-9238.

URL https://www.jstor.org/
stable/2981161.

Good:1952:CRM

[Goo52b] I. J. Good. Corrigenda: Random motion on a finite Abelian group. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 48:368, 1952. CODEN PCPSA4. ISSN 0008-1981.

Good:1952:GDM

[Goo52c] I. J. Good. A generalisation of Dirichlet’s multiple integral. Edinburgh Math. Notes, 38:7–8, 1952. ISSN 0950-1843,2051-2031.

Good:1952:RD

[Goo52d] I. J. Good. Rational decisions.

Journal of the Royal Statistical Society. Series B (Methodological), 14(1):107–114, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL https://www.jstor.
org/stable/2984087.

Good:1953:PFS

[Goo53a] I. J. Good. The population frequencies of species and the estimation of population parameters. Biometrika, 40(3/4):237– 264, December 1953. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2333344.

Good:1953:STS

[Goo53b] I. J. Good. The serial test for sampling numbers and other tests for randomness. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 49(??):276–284, 1953. CODEN PCPSA4. ISSN 0008-1981.

Good:1955:BRB

[Goo55a] I. J. Good. Book review: The Foundations of Statistics, by L. J. Savage. Journal of the Royal Statistical Society. Series A (General), 118(2):245–246, ???? 1955.

CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2343133.

Good:1955:CRW

[Goo55b] I. J. Good. A classification of rules for writing informative English. Methodos, 7:193–200, ???? 1955. ISSN 0369-2361.

Good:1955:CCM

[Goo55c] I. J. Good. Conjectures concerning the Mersenne numbers. Mathematical Tables and Other Aids to Computation, 9(51):120– 121, September 1955. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Good:1955:JDS

[Goo55d] I. J. Good. The joint distribution for the sizes of the generations in a cascade process. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 51:240–242, 1955. CO-DEN PCPSA4. ISSN 0008-1981.

Good:1955:LRT

[Goo55e] I. J. Good. The likelihood ratio test for Markoff chains. Biometrika, 42(3/4): 531–533, December 1955. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2333404. See corrigenda

[Goo57c].

Good:1955:NFS

[Goo55f] I. J. Good. A new finite series for Legendre polynomials. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 51:385– 388, 1955. CODEN PCPSA4. ISSN 0008-1981.

Good:1955:MCP

[Goo55g] I. J. Good. On the marking of chess-players. Mathematical Gazette, 39:292–296, 1955. CO-DEN MAGAAS. ISSN 0025-5572,2056-6328.

Good:1955:WCS

[Goo55h] I. J. Good. On the weighted combination of significance tests. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):264–265, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL https://www.jstor.
org/stable/2983961.

Good:1956:BRBb

[Goo56a] I. J. Good. Book review: Fondements d’une Théorie Positive des Choix Comportant un Risque et Critique des Postulats et Axiomes de l’École Americaine, by M. Allais. Journal of the Royal Statistical Society. Series A (General), 119(2):213–214, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2342890.

Good:1956:BRBa

[Goo56b] I. J. Good. Book review: Theory of Games and Statistical Decisions. D. Blackwell and M. A. Girschick, New York: John Wiley and Sons; London: Chapman and Hall, 1954. Pp. xi, 355. Journal of the American Statistical Association, 51(274):388–390, June 1956.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

                                        tronic).  URL https://www.
                                        jstor.org/stable/2281364.

Good:1956:ESF

[Goo56c] I. J. Good. On the estimation of small frequencies in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 18(1): 113–124, ???? 1956. CO-DEN JSTBAJ. ISSN 0035-9246.

                                        URL https://www.jstor.org/
                                        stable/2983740.

Good:1956:STN

[Goo56d] I. J. Good. Some terminology and notation in information theory. Proc. Inst. Elec. Engrs. C., 103:200–204, 1956. ISSN 0369-8904,2054-0426.

Good:1956:SIM

[Goo56e] I. J. Good. The surprise index for the multivariate normal distribution. Annals of Mathematical Statistics, 27(4):1130–1135, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        aoms/1177728079. See correction

[Goo57b].

Good:1957:BRB

[Goo57a] I. J. Good. Book review: Language as Choice and Chance, by G. Herdan. Journal of the Royal Statistical Society. Series A (General), 120(1):89–91, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL https://www.

jstor.org/stable/2342558.

Good:1957:CSI

[Goo57b] I. J. Good. Corrections to “The Surprise Index for the Multivariate Normal Distribution”. Annals of Mathematical Statistics, 28(4):1055, December 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.
org/euclid.aoms/1177706815.

See [Goo56e].

Good:1957:CLR

[Goo57c] I. J. Good. Corrigenda: “The Likelihood Ratio Test for Markoff Chains”. Biometrika, 44(1/2): 301, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2333289. See [Goo55e].

Good:1957:NL

[Goo57d] I. J. Good. A needle for the lecturer. New Scientist, 2(46):38, October 3, 1957. CODEN NWS-CAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Good:1957:MPN

[Goo57e] I. J. Good. On the masses of the proton, neutron and hyperons. J. Roy. Naval Sci. Service, 12:144, 1957.

Good:1957:NSI

[Goo57f] I. J. Good. On the numerical solution of integral equations. Mathematical Tables and Other Aids to Computation, 11 (58):82–83, April 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

URL https://www.jstor.org/
stable/2002156.

Good:1957:STR

[Goo57g] I. J. Good. On the serial test for random sequences. Annals of Mathematical Statistics, 28(1):262–264, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.
aoms/1177707053;   https://
www.jstor.org/stable/2237039.

Good:1957:SPM

[Goo57h] I. J. Good. Saddle-point methods for the multinomial distribution. Annals of Mathematical Statistics, 28(4):861–881, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.
aoms/1177706790. See correc-

tions [Goo61i].

Good:1958:BRB

[Goo58a] I. J. Good. Book review: Probability, Statistics and Truth, by Richard von Mises; Hilda Geiringer. Journal of the Royal Statistical Society. Series A (General), 121(2):238–240, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL https://www.

jstor.org/stable/2343367.

Good:1958:CMMa

[Goo58b] I. J. Good. Could a machine make probability judgments? I. Computers and Automation, 8:14–16, 1958. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549.

Good:1958:CMMb

[Goo58c] I. J. Good. Could a machine make probability judgments? II. Computers and Automation, 8:24–26, 1958. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549.

Good:1958:HMS

[Goo58d] I. J. Good. How much science can you have at your fingertips? IBM Journal of Research and Development, 2(4):282–288, ???? 1958. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL https:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5392640.

Good:1958:IAP

[Goo58e] I. J. Good. The interaction algorithm and practical Fourier analysis. Journal of the Royal Statistical Society. Series B (Methodological), 20 (2):361–372, July 1958. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2983896.  Addendum

in same journal, 22(2) 429–430, July 1960 [Goo60d].

Good:1958:LPT

[Goo58f] I. J. Good. Legendre polynomials and trinomial random walks. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 54:39– 42, 1958. CODEN PCPSA4. ISSN 0008-1981.

Good:1958:RMA

[Goo58g] I. J. Good. Random motion and analytic continued fractions. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 54:43– 47, 1958. CODEN PCPSA4. ISSN 0008-1981.

Good:1958:R

[Goo58h] I. J. Good. Reviews. British Journal for the Philosophy of Science, IX(35):251–255, November 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

                                        //bjps.oxfordjournals.org/
                                        content/IX/35/251.full.pdf+
                                        html;   https://www.jstor.
                                        org/stable/685660.

Good:1958:STP

[Goo58i] I. J. Good. Significance tests in parallel and in series. Journal of the American Statistical Association, 53(284):799–813, December 1958. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL https://www.

                                        jstor.org/stable/2281953.

Good:1959:CFA

[Goo59a] I. J. Good. A classification of fallacious arguments and interpretations. Methodos, 11:147–159, 1959. ISSN 0369-2361.

Good:1959:KP

[Goo59b] I. J. Good. Kinds of probability. Science, 129:443–447, 1959.

CODEN SCIEAS. ISSN 0036-8075,1095-9203.

Good:1959:LSS

[Goo59c] I. J. Good. Lattice structure of space–time. British Journal for the Philosophy of Science, IX(36):317–319, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/IX/36/317.full.pdf+
html;   https://www.jstor.
org/stable/685192.

Good:1959:TC

[Goo59d] I. J. Good. A theory of causality.

British Journal for the Philosophy of Science, IX(36):307–310, February 1959. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/IX/36/307.full.pdf+
html;   https://www.jstor.
org/stable/685190.

Good:1960:CSM

[Goo60a] I. J. Good. A comparison of some methods of calculating covariance functions on an electronic computer. Computer Journal, 3: 262–265, 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Good:1960:ESR

[Goo60b] I. J. Good. Effective sampling rates for signal detection; or can the Gaussian model be salvaged? Information and Control, 3:116– 140, 1960. CODEN IFCNA4. ISSN 0019-9958,1878-2981.

Good:1960:GSV

[Goo60c] I. J. Good. Generalizations to several variables of Lagrange’s expansion, with applications to stochastic processes. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 56:367–380, 1960. CODEN PCPSA4. ISSN 0008-1981.

Good:1960:IAP

[Goo60d] I. J. Good. The interaction algorithm and practical Fourier analysis: an addendum. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):372–375, ???? 1960. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984108. See [Goo58e].

Good:1960:MCM

[Goo60e] I. J. Good. Monte Carlo method. In McGraw-Hill Encyclopedia of Science and Technology, volume 8, pages 586–587. McGraw-Hill, New York, NY, USA, 1960. LCCN ????

Good:1960:PC

[Goo60f] I. J. Good. The paradox of confirmation. British Journal for the Philosophy of Science, XI(42):145b–145, August 1960. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://bjps.
oxfordjournals.org/content/
XI/42/145-b.full.pdf+html;

https://www.jstor.org/stable/
685588.

Good:1960:WEC

[Goo60g] I. J. Good. Weight of evidence, corroboration, explanatory power, information and the utility of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):319–331, ???? 1960. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984102. See corrigen-

dum [Goo68e].

Good:1961:ADD

[Goo61a] I. J. Good. Amount of deciding and decisionary effort. Information and Control, 4:271–281, 1961. CODEN IFCNA4. ISSN 0019-9958,1878-2981.

Good:1961:ACS

[Goo61b] I. J. Good. Analysis of cumulative sums by multiple contour integration. Quart. J. Math. Oxford Ser. (2), 12:115–122, 1961. ISSN 0033-5606,1464-3847.

Good:1961:AFD

[Goo61c] I. J. Good. An asymptotic formula for the differences of the powers at zero. Annals of Mathematical Statistics, 32(1):249–256, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.
aoms/1177705156.

Good:1961:BRB

[Goo61d] I. J. Good. Book review: Entropy and the Unity of Knowledge, by P. T. Landsberg. Journal of the Royal Statistical Society. Series A (General), 124(4):578, ???? 1961. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2342943.

Good:1961:CC

[Goo61e] I. J. Good. A causal calculus (I). British Journal for the Philosophy of Science, 11(44):305–318, February 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/XI/44/305.full.pdf+
html;   https://www.jstor.
org/stable/685131.

Good:1961:CCI

[Goo61f] I. J. Good. A causal calculus (II). British Journal for the Philosophy of Science, 12(45):43–51, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/XII/45/43.full.pdf+
html;   https://www.jstor.
org/stable/685537.

Good:1961:CMC

[Goo61g] I. J. Good. The colleague matrix, a Chebyshev analogue of the companion matrix. The Quarterly Journal of Mathematics. Oxford. Second Series, 12:61–68, 1961. ISSN 0033-5606,1464-3847.

Good:1961:CSMa

[Goo61h] I. J. Good. A comparison of some methods of calculating covariance functions on an electronic computer. Computer Journal, 3(4):262–265, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
3/4/262.full.pdf+html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/030262.sgm.abs.html;   http:
//www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_
04/tiff/262.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/
263.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/264.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/265.tif.

Good:1961:CNC

[Goo61i] I. J. Good. Corrections notes:

Corrections to “Saddle-point Methods for the Multinomial Distribution”. Annals of Mathematical Statistics, 32(2):619, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.
aoms/1177705072. See [Goo57h].

Good:1961:CSMb

[Goo61j] I. J. Good. Corrections to “Saddlepoint methods for the multinomial distributions”. Annals of Mathematical Statistics, 32: 619, 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

Good:1961:FCM

[Goo61k] I. J. Good. The frequency count of a Markov chain and the transition to continuous time. Annals of Mathematical Statistics, 32(1):41–48, March 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.aoms/1177705137.

Good:1961:MSM

[Goo61l] I. J. Good. The multivariate saddlepoint method and chisquared for the multinomial distribution. Annals of Mathematical Statistics, 32(2):535–548, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        aoms/1177705059.

Good:1961:PCI

[Goo61m] I. J. Good. The paradox of confirmation (II). British Journal for the Philosophy of Science, XII(45):63–64, May 1961. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

                                        //bjps.oxfordjournals.org/
                                        content/XII/45/63.full.pdf+
                                        html;   https://www.jstor.
                                        org/stable/685539.

Good:1961:RSC

[Goo61n] I. J. Good. The real stable characteristic functions and chaotic acceleration. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):

180–183, ???? 1961. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2983855.

Good:1961:WEC

[Goo61o] I. J. Good. Weight of evidence, causality and false-alarm probabilities. In Information theory (Symposium, London, 1960), pages 125–136. Butterworths, Washington, DC, USA, 1961.

Good:1962:APS

[Goo62a] I. J. Good. Analogues of Poisson’s summation formula. American Mathematical Monthly, 69 (4):259–266, April 1962. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Good:1962:BRBb

[Goo62b] I. J. Good. Book review: Coding Theorems of Information Theory, by J. Wolfowitz. Journal of the Royal Statistical Society. Series A (General), 125(4):643–645, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL https://www.

jstor.org/stable/2982637.

Good:1962:BRBa

[Goo62c] I. J. Good. Book review: Theory of Probability, by Harold Jeffreys. Journal of the Royal Statistical Society. Series A (General), 125(3):487–489, ???? 1962. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2982417.

Good:1962:CFA

[Goo62d] I. J. Good. A classification of fallacious arguments and interpretations. Technometrics, 4(1): 125–132, February 1962. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1266177.

Good:1962:CAC

[Goo62e] I. J. Good. Corrigenda: “Analysis of cumulative sums by multiple contour integration”. Quart. J. Math. Oxford Ser. (2), 13:80, 1962. ISSN 0033-5606,1464-3847.

Good:1962:EC

[Goo62f] I. J. Good. Errata and corrigenda. British Journal for the Philosophy of Science, XIII (49):88, May 1962. CO-DEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://bjps.
oxfordjournals.org/content/
XIII/49/88.full.pdf+html;
https://www.jstor.org/stable/
685237.

Good:1962:PSB

[Goo62g] I. J. Good. Proofs of some ‘binomial’ identities by means of MacMahon’s ‘Master Theorem’. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 58:161–162, 1962. CODEN PCPSA4. ISSN 0008-1981.

Good:1962:SPMa

[Goo62h] I. J. Good. A short proof of MacMahon’s ‘Master Theorem’.

Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 58:160, 1962. CODEN PCPSA4. ISSN 0008-1981.

Good:1962:SPMb

[Goo62i] I. J. Good. Subjective probability as the measure of a non-measureable set. In Logic, Methodology and Philosophy of Science (Proceedings of the 1960 International Congress), pages 319–329. Stanford University Press, Stanford, CA, USA, 1962.

Good:1963:BRD

[Goo63a] I. J. Good. Book review: Decisions, Values and Groups, Vol. 2, by Norman F. Washburne. Biometrics, 19(3):502, September 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2527943.

Good:1963:MEH

[Goo63b] I. J. Good. Maximum entropy for hypothesis formulation, especially for multidimensional contingency tables. Annals of Mathematical Statistics, 34(3): 911–934, September 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.
org/euclid.aoms/1177704014.

Good:1963:IQE

[Goo63c] I. J. Good. On the independence of quadratic expressions. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):377–382, ???? 1963.

CODEN JSTBAJ. ISSN 0035-

                                        9246. URL https://www.jstor.
                                        org/stable/2984305. See corri-

genda [Goo66e].

Good:1963:QMC

[Goo63d] I. J. Good. Quadratics in Markov-chain frequencies, and the binary chain of order 2. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):383–391, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL https://www.jstor.
                                        org/stable/2984306.

Good:1963:WCE

[Goo63e] I. J. Good. Weighted covariance for estimating the direction of a Gaussian source. In Proc. Sympos. Time Series Analysis (Brown Univ., 1962), pages 447–470. Wiley, New York, NY, USA, 1963.

Good:1964:BRB

[Goo64] I. J. Good. Book review: Information Theory and Coding, by Norman Abramson. Journal of the Royal Statistical Society. Series A (General), 127 (3):454–455, ???? 1964. CO-DEN JSSAEF. ISSN 0035-9238.

                                        URL https://www.jstor.org/
                                        stable/2343598.

Good:1965:BMM

[Goo65a] I. J. Good. Brains, machines and mathematics. Computer Journal, 8(1):88, April 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://comjnl.

oxfordjournals.org/content/
8/1/88.full.pdf+html.

Good:1965:GLE

[Goo65b] I. J. Good. The generalization of Lagrange’s expansion and the enumeration of trees. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 61:499–517, 1965. CODEN PCPSA4. ISSN 0008-1981.

Good:1965:LE

[Goo65c] I. J. Good. Letter to the Editor. The American Statistician, 19(3):43, June 1965. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2681423.

Good:1965:EPE

[Goo65d] Irving John Good. The Estimation of Probabilities; an Essay on Modern Bayesian Methods, volume 30 of MIT Press Research Monographs. MIT Press, Cambridge, MA, USA, 1965. xii + 109 pp. LCCN QA273 .G64.

Good:1965:SCF

[Goo65e] Irving John Good. Speculations concerning the first ultraintelligent machine. In Franz L. Alt and Morris Rubinoff, editors, Advances in Computers, volume 6, pages 31–88. Academic Press, New York, NY, USA, 1965. See [Nor12, Yud13].

Good:1966:BRBa

[Goo66a] I. J. Good. Book review: Mathematical Theory of Probability and Statistics, by Richard von Mises; Hilda Geiringer. Journal of the Royal Statistical Society. Series A (General), 129(2):289–291, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL https://www.

jstor.org/stable/2343628.

Good:1966:BRBb

[Goo66b] I. J. Good. Book review: Nombres et Systèmes aléatoires, by Octav Onicescu. Journal of the Royal Statistical Society. Series A (General), 129(2):291, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2343629.

Good:1966:CIQ

[Goo66c] I. J. Good. Corrigenda: “On the independence of quadratic expressions”. Journal of the Royal Statistical Society. Series B (Methodological), 28: 584, 1966. CODEN JSTBAJ. ISSN 0035-9246. URL http:/

/links.jstor.org/sici?sici=
0035-9246(1966)28:3<584:OTIOQE>
2.0.CO;2-0&origin=MSN.

Good:1966:DPE

[Goo66d] I. J. Good. A derivation of the probabilistic explication of information. Journal of the Royal Statistical Society. Series B (Methodological), 28(3): 578–581, ???? 1966. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984450.

Good:1966:IQE

[Goo66e] I. J. Good. On the independence of quadratic expressions: Corrigenda. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):584, ???? 1966. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984452. See [Goo63c].

Good:1966:PW

[Goo66f] I. J. Good. The probability of war. Journal of the Royal Statistical Society. Series A (General), 129(2):268–269, ???? 1966. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.
org/stable/2343620.

Good:1967:NCY

[Goo67a] I. J. Good. 234. Note: Checks on Yates’s algorithm. Biometrics, 23(3):573, September 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic).  URL https://www.
jstor.org/stable/2528018.

Good:1967:BST

[Goo67b] I. J. Good. A Bayesian significance test for multinomial distributions. (With discussion). Journal of the Royal Statistical Society. Series B (Methodological), 29(3):399–431, ???? 1967. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984384. See [Goo74b].

Good:1967:HML

[Goo67c] I. J. Good. Human and machine logic. British Journal for the Philosophy of Science, 18(2):144–147, August 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/18/2/144.full.pdf+
html;   https://www.jstor.
org/stable/686582.

Good:1967:PTE

[Goo67d] I. J. Good. On the principle of total evidence. British Journal for the Philosophy of Science, 17(4):319–321, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/17/4/319.full.pdf+
html;   https://www.jstor.
org/stable/686773.

Good:1967:WSR

[Goo67e] I. J. Good. The white shoe is a red herring. British Journal for the Philosophy of Science, 17(4):322, February 1967. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/17/4/322.full.pdf+
html;   https://www.jstor.
org/stable/686774.

Good:1968:CFFa

[Goo68a] I. J. Good. The characteristic functions of functions. Nature, 218:603–604, 1968. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Good:1968:CFFb

[Goo68b] I. J. Good. Characteristic functions of functions. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 307:317–334, 1968. CODEN PRLAAZ. ISSN 0962-8444,2053-9169.

Good:1968:CGL

[Goo68c] I. J. Good. Correction: “The generalization of Lagrange’s expansion and the enumeration of trees”. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 64:489, 1968. CODEN PCPSA4. ISSN 0008-1981.

Good:1968:CGT

[Goo68d] I. J. Good. Correspondence:

Gödel’s theorem. Computer Journal, 10(4):417, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/417.
tif.

Good:1968:CWE

[Goo68e] I. J. Good. Corrigendum: Weight of evidence, corroboration, explanatory power, information and the utility of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):203, ???? 1968. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984472. See [Goo60g].

Good:1968:CEE

[Goo68f] I. J. Good. Corroboration, explanation, evolving probability, simplicity and a sharpened razor. British Journal for the Philosophy of Science, 19(2):123–143, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/19/2/123.full.pdf+
html;   https://www.jstor.
org/stable/686791.

Good:1968:GT

[Goo68g] I. J. Good. Gödel’s Theorem. Computer Journal, 10 (4):417, February 1968. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.
oxfordjournals.org/content/
10/4/417.1.full.pdf+html.

Good:1968:LEN

[Goo68h] I. J. Good. Letter to the Editor: number of possible strategies when writing compilers. Communications of the ACM, 11(7):474, July 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Good:1968:WSQ

[Goo68i] I. J. Good. The white shoe qua herring is pink. British Journal for the Philosophy of Science, 19(2):156–157, August 1968. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/

content/19/2/156.full.pdf+
html;   https://www.jstor.
org/stable/686795.

Good:1969:CQF

[Goo69a] I. J. Good. Conditions for a quadratic form to have a chisquared distribution. Biometrika, 56(1):215–216, March 1969. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2334707. See correction

[Goo70b] and remark [Kha78].

Good:1969:DBF

[Goo69b] I. J. Good. Discussion of Bruno de Finetti’s paper ‘Initial probabilities: A prerequisite for any valid induction’. Synthese, 20(1):17–24, June 1969. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/article/
10.1007/BF00567233.

Good:1969:GTR

[Goo69c] I. J. Good. Gödel’s Theorem is a red herring. British Journal for the Philosophy of Science, 19(4):357–358, February 1969. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/19/4/357.full.pdf+
html;   https://www.jstor.
org/stable/686295.

Good:1969:HRR

[Goo69d] I. J. Good. How random are random numbers? The American Statistician, 23(4): 42–45, October 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2681742.

Good:1969:SAS

[Goo69e] I. J. Good. Some applications of the singular decomposition of a matrix. Technometrics, 11(4): 823–831, November 1969. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL https://www.jstor.org/
                                        stable/1266902.

Good:1969:SEB

[Goo69f] I. J. Good. A subjective evaluation of Bode’s law and an ’objective’ test for approximate numerical rationality. Journal of the American Statistical Association, 64(325):23–49, March 1969. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2283718. See discussion

[ABH^{+}69].

Good:1969:WUD

[Goo69g] I. J. Good. What is the use of a distribution? In Multivariate Analysis, II (Proc. Second Internat. Sympos., Dayton, Ohio, 1968), pages 183–203. Academic Press, New York, NY, USA, 1969.

Good:1970:BR

[Goo70a] I. J. Good. Book review. Computer Journal, 13(3):236, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/
content/13/3/236.full.pdf+
html.

Good:1970:CCQ

[Goo70b] I. J. Good. Corrections: “Conditions for a Quadratic Form to Have a Chi-Squared Distribution”. Biometrika, 57(1): 225, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2334962. See [Goo69a].

Good:1970:GBE

[Goo70c] I. J. Good. A generalization of the Bernoulli–Euler partition formula. Scripta Mathematica, 28 (??):319–320, ???? 1970. ISSN 0036-9713.

Good:1970:ICM

[Goo70d] I. J. Good. The inverse of a centrosymmetric matrix. Technometrics, 12(4):925–928, November 1970. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1267339.

Good:1970:SPC

[Goo70e] I. J. Good. Short proof of a conjecture by Dyson. Journal of Mathematical Physics, 11(6):1884, June 1970. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/
v11/i6/p1884_s1.

Good:1970:SSM

[Goo70f] I. J. Good. Some statistical methods in machine intelligence research. Mathematical Biosciences, 6:185–208, 1970. CO-DEN MABIAR. ISSN 0025-5564,1879-3134.

Good:1970:SRM

[Goo70g] I. J. Good. A suggested resolution of Miller’s Paradox. British Journal for the Philosophy of Science, 21(3):288–289, August 1970. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

                                        //bjps.oxfordjournals.org/
                                        content/21/3/288.full.pdf+
                                        html;   https://www.jstor.
                                        org/stable/686203.

Good:1970:RPM

[Goo70h] Irving John Good. Review of Philip McShane, Randomness, Statistics, and Emergence, Gill and MacMillan, 1970. Times Literary Supplement, ??(??):1043, September 18, 1970. See book [McS70].

Good:1971:DOS

[Goo71a] I. J. Good. Does an observed sequence of numbers follow a simple rule? (another look at Bode’s law): Comment. Journal of the American Statistical Association, 66(335):559– 562, September 1971. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2283523.

Good:1971:FSM

[Goo71b] I. J. Good. The factorization of a sum of matrices and the multivariate cumulants of a set of quadratic expressions. Journal of Combinatorial Theory (Series A), 11:27–37, 1971. CO-DEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Good:1971:FWS

[Goo71c] I. J. Good. Free will and speed of computation. British Journal for the Philosophy of Science, 22(1):48–50, February 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (elec-

tronic).  URL http://bjps.
oxfordjournals.org/content/
22/1/48.full.pdf+html.

Good:1971:PEI

[Goo71d] I. J. Good. The probabilistic explication of information, evidence, surprise, causality, explanation, and utility. In Foundations of statistical inference (Proc. Sympos., Univ. Waterloo, Waterloo, Ont., 1970), pages 108–141. Holt, Rinehart and Winston of Canada, Ltd., Toronto, ON, Canada, 1971. With comments by G. A. Barnard, I. D. J. Bross, V. P. Godambe and H. E. Kyburg and a reply by the author.

Good:1971:RBT

[Goo71e] I. J. Good. The relationship between two Fast Fourier Transforms. IEEE Transactions on Computers, C-20(3):310–317, March 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=1671829.

Good:1971:R

[Goo71f] I. J. Good. Reviews. British Journal for the Philosophy of Science, 22(4):382–387, November 1971. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

                                        //bjps.oxfordjournals.org/
                                        content/22/4/382.full.pdf+
                                        html;   https://www.jstor.
                                        org/stable/686322.

Good:1971:WP

[Goo71g] I. J. Good. Wilfred Perks. Journal of the Royal Statistical Society. Series A (General), 134(1):124, ???? 1971. CO-DEN JSSAEF. ISSN 0035-9238.

                                        URL https://www.jstor.org/
                                        stable/2343998.

Good:1972:NCP

[Goo72a] I. J. Good. 342. Note: Correlation for power functions. Biometrics, 28(4):1127– 1129, December 1972. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

                                        URL https://www.jstor.org/

stable/2528645. See correction, ibid. 29 (1973), 829.

Good:1972:BRM

[Goo72b] I. J. Good. Book review: Multitype Branching Processes. Theory and Applications (Charles J. Mode). SIAM Review, 14(3):

512–513, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1972:BRB

[Goo72c] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 67 (337):245–246, March 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2284742.

Good:1972:CPF

[Goo72d] I. J. Good. Correlation for power functions. Biometrics, 28:1127– 1129; correction, ibid. 29 (1973), 829, 1972. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Good:1972:EFS

[Goo72e] I. J. Good. Errata: “The factorization of a sum of matrices and the multivariate cumulants of a set of quadratic expressions”. Journal of Combinatorial Theory (Series A), 12:309, 1972. CO-DEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Good:1972:SPP

[Goo72f] I. J. Good. Some probability paradoxes in choice from among random alternatives: Comment. Journal of the American Statistical Association, 67 (338):374–375, June 1972. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2284385.

Good:1972:STP

[Goo72g] I. J. Good. Statistics and Today’s problems. The American Statistician, 26(3):11–19, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic).  URL https://www.
                                        jstor.org/stable/2682859.

Good:1972:WMA

[Goo72h] I. J. Good. What is the most amazing approximate integer in the universe? Pi Mu Epsilon Journal, 5:314–315, 1972. CO-DEN PMEJBR. ISSN 0031-952X.

Good:1973:BRB

[Goo73a] I. J. Good. Book review: The Generation of Random Variates by T. G. Newman; P. L. Odell. International Statistical Review = Revue Internationale de Statistique, 41(1):139–140, April 1973. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

                                        tronic).  URL https://www.
                                        jstor.org/stable/1402798.

Good:1973:JPG

[Goo73b] I. J. Good. The joint probability generating function for run-lengths in regenerative binary Markov chains, with applications. Annals of Statistics, 1(5): 933–939, September 1973. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.aos/1176342513.

Good:1973:PSA

[Goo73c] I. J. Good. Problems and solutions: Advanced problems:

5897. American Mathematical Monthly, 80(2):208–209, February 1973. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BK75, Goo75i, RC74, SB74, SH74].

Good:1973:WDF

[Goo73d] I. J. Good. What are degrees of freedom? The American Statistician, 27(5):227–228, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL https://www.

jstor.org/stable/3087407.

Good:1974:CCC

[Goo74a] I. J. Good. A correction concerning complexity. British Journal for the Philosophy of Science, 25(3):289, September 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/25/3/289.full.pdf+
html;   https://www.jstor.
org/stable/686911.

Good:1974:CBS

[Goo74b] I. J. Good. Corrigendum:

“A Bayesian significance test for multinomial distributions” (J. Roy. Statist. Soc. Ser. B 29 (1967), 399–431). Journal of the Royal Statistical Society. Series B (Methodological), 36(1):109, ???? 1974. CO-DEN JSTBAJ. ISSN 0035-9246.

URL https://www.jstor.org/
stable/2984778. See [Goo67b].

Good:1974:LLC

[Goo74c] I. J. Good. A little learning can be dangerous. British Journal for the Philosophy of Science, 25(4):340–342, December 1974. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

                                        //bjps.oxfordjournals.org/
                                        content/25/4/340.full.pdf+
                                        html;   https://www.jstor.
                                        org/stable/686463.

Good:1974:RSF

[Goo74d] I. J. Good. A reciprocal series of Fibonacci numbers. Fibonacci Quarterly, 12(4):346–??, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:/

                                        /www.fq.math.ca/Scanned/12-
                                        4/good.pdf.

Good:1974:RTA

[Goo74e] Irving John Good. Random thoughts about randomness. In Schaffner and Cohen [SC74], pages 117–135. ISBN 90-277-0408-2, 90-277-0409-0 (paperback). ISSN 0068-0346. LCCN Q175 .B7312. URL https:/

                                        /link.springer.com/chapter/
                                        10.1007/978-94-010-2140-1_9.

Invited lecture in the symposium on the Concept of Randomness, dedicated to the memory of L. J. Savage, in the Biennial Meeting of the Philosophy of Science Association, Olds Plaza Hotel, Lansing, Michigan, October 27– 29, 1972.

Good:1975:BFA

[Goo75a] I. J. Good. The Bayes factor against equiprobability of a multinomial population assuming a symmetric Dirichlet prior. Annals of Statistics, 3(1):246–250, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.
aos/1176343015.

Good:1975:BRB

[Goo75b] I. J. Good. Book review: An Introduction to Bayesian Inference in Econometrics by Arnold Zellner. Technometrics, 17(1): 137–138, February 1975. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1268014. See amend-

ment [Goo76a].

Good:1975:CDM

[Goo75c] I. J. Good. Comments on David Miller. Synthese, 30 (1–2):205–206, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/
10.1007/BF00485307.  See

[Mil75a] and reply [Mil75b].

Good:1975:CJA

[Goo75d] I. J. Good. Comments on Joseph Agassi. Synthese, 30 (1–2):31, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF00485292.  See

[Aga75a], comments [Har75, Wet75], and reply [Goo75j].

Good:1975:CRG

[Goo75e] I. J. Good. Comments on ronald giere. Synthese, 30(1– 2):133, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF00485300.

Good:1975:LDB

[Goo75f] I. J. Good. The Lagrange distributions and branching processes. SIAM Journal on Applied Mathematics, 28(2):270–275, March 1975. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic).

Good:1975:NFC

[Goo75g] I. J. Good. A new formula for cumulants. Mathematical Proceedings of the Cambridge Philosophical Society, 78(2):333–337, 1975. CODEN MPCPCO. ISSN 0305-0041,1469-8064.

Good:1975:NOC

[Goo75h] I. J. Good. The number of orderings of n candidates when ties are permitted. Fibonacci Quarterly, 13(1):11–18, February 1975. CO-DEN FIBQAU. ISSN 0015-0517.

                                        URL http://www.fq.math.ca/
                                        Scanned/13-1/good.pdf.

Good:1975:PSS

[Goo75i] I. J. Good. Problems and solutions: Solutions of advanced problems: 5897. American Mathematical Monthly, 82(5):532–533, May 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ber73].

Good:1975:R

[Goo75j] I. J. Good. Replies. Synthese, 30(1–2):83–93, February/ March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/article/
10.1007/BF00485297.  See

[Goo75d, Aga75a, Har75, Wet75].

Good:1975:ECR

[Goo75k] Irving John Good. Explicativity, corroboration, and the relative odds of hypotheses. Synthese, 30(1–2):39–73, February/ March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/article/
10.1007/BF00485294.

Good:1976:ARA

[Goo76a] I. J. Good. Amendment to review of Arnold Zellner’s An Introduction to Bayesian Inference in Econometrics. Technometrics, 18(1):123, February 1976. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1267931. See [Goo75b].

Good:1976:BIH

[Goo76b] I. J. Good. The Bayesian influence, or how to sweep subjectivism under the carpet. In Foundations of probability theory, statistical inference, and statistical theories of science (Proc. Internat. Res. Colloq., Univ. Western Ontario, London, Ont., 1973), Vol. II, volume Vol. 6 of University of Western Ontario Ser. Philos. Sci., pages 125–174. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0618-2.

Good:1976:Ma

[Goo76c] I. J. Good. Minicommunications. Communications in Statistics: Simulation and Computation, 5(1):81–84, 1976. CODEN CSSCDB. ISSN 0361-0918.

Good:1976:ASD

[Goo76d] I. J. Good. On the application of symmetric Dirichlet distributions and their mixtures to contingency tables. Annals of Statistics, 4(6): 1159–1189, November 1976. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.
org/euclid.aos/1176343649.

Good:1976:RFC

[Goo76e] I. J. Good. The relationship of a formula of Carlitz to the generalized Lagrange expansion. SIAM Journal on Applied Mathematics, 30(1):103, January 1976. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic).

Good:1976:R

[Goo76f] I. J. Good. Reviews. British Journal for the Philosophy of Science, 27(3):289–292, September 1976. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/27/3/289.full.pdf+
html;   https://www.jstor.
org/stable/686127.

Good:1977:CDL

[Goo77a] I. J. Good. C2. Does −2 log(likelihoodratio)

exceed X^{2}? Journal of Statistical Computation and Simulation, 6 (1):70, 1977. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1977:EMT

[Goo77b] I. J. Good. Explicativity: a mathematical theory of explanation with statistical applications. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 354 (1678):303–330, 1977. CODEN PRLAAZ. ISSN 0962-8444,2053-9169.

Good:1977:NFS

[Goo77c] I. J. Good. A new formula for kstatistics. Annals of Statistics, 5 (1):224–228, January 1977. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.
org/euclid.aos/1176343757.

Good:1978:AOT

[Goo78a] I. J. Good. Alleged objectivity; a threat to the human spirit? A reply to “Bayesianism; a threat to the statistical profession?” (Internat. Statist. Rev. 45 (1977), 111–115) by H. C. Hamaker. International Statistical Review = Revue Internationale de Statistique, 46(1):65–66, April 1978. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic).  URL https://www.
jstor.org/stable/1402508.

Good:1978:BRBb

[Goo78b] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 73 (362):441, June 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2286689.

Good:1978:BRBc

[Goo78c] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 73 (362):441–442, June 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2286690.

Good:1978:BRBd

[Goo78d] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 73 (362):442, June 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2286691.

Good:1978:CMA

[Goo78e] I. J. Good. C 24. A method for avoiding overflows and underflows. Journal of Statistical Computation and Simulation, 8(2): 162–163, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CML

[Goo78f] I. J. Good. C10. Are maximumlikelihood estimates invariant? Journal of Statistical Computation and Simulation, 7(1):80– 81, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CCN

[Goo78g] I. J. Good. C11. The complexity of nutrition advice. Journal of Statistical Computation and Simulation, 7(1):81, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CSO

[Goo78h] I. J. Good. C12. Should outliers be dragged in? Journal of Statistical Computation and Simulation, 7(2):163–164, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMK

[Goo78i] I. J. Good. C14. Moments of the Kolmogorov–Smirnov onesample statistic. Journal of Statistical Computation and Simulation, 7(3–4):289–290, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CST

[Goo78j] I. J. Good. C15. A signs test when estimating probability densities. Journal of Statistical Computation and Simulation, 7(3–4): 290–292, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CET

[Goo78k] I. J. Good. C16. Ethical treatments. Journal of Statistical Computation and Simulation, 7(3–4):292–295, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CHC

[Goo78l] I. J. Good. C17. Histograms with 2^{n} classes. Journal of Statistical Computation and Simulation, 7(3–4):295, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CFB

[Goo78m] I. J. Good. C18. A fuzzy Bayesian method for estimating probabilities given two related multinomial distributions. Journal of Statistical Computation and Simulation, 7(3–4):296–299, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CPA

[Goo78n] I. J. Good. C22. Path analysis and correlation for power functions. Journal of Statistical Computation and Simulation, 8 (1):80, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMC

[Goo78o] I. J. Good. C23. Monozygotic criminals. Journal of Statistical Computation and Simulation, 8(2):161–162, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CIT

[Goo78p] I. J. Good. C5. Interpolation in tables of the normal distribution. Journal of Statistical Computation and Simulation, 6(3–4): 313, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CMV

[Goo78q] I. J. Good. C6. A minimalvocabulary algorithm? Journal of Statistical Computation and Simulation, 6(3–4):314, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CAA

[Goo78r] I. J. Good. C7. Adenine arabinoside therapy. Journal of Statistical Computation and Simulation, 6(3–4):314–315, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:CIC

[Goo78s] I. J. Good. C9. An improvement in the chi-squared goodness-of-fit test. Journal of Statistical Computation and Simulation, 7(1): 79–81, 1978. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1978:RBF

[Goo78t] I. J. Good. Review: Fractals: Form, Chance, and Dimension, by Benoît B. Mandelbrot. Journal of the American Statistical Association, 73 (362):438, June 1978. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2286682.

Good:1979:BRBa

[Goo79a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 74 (366):501–502, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2286363.

Good:1979:BRBb

[Goo79b] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 74 (366):502–503, June 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2286364.

Good:1979:BRBc

[Goo79c] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 74 (367):732, September 1979. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2287011.

Good:1979:CCRa

[Goo79d] I. J. Good. C25. A comment on runs of signs. Journal of Statistical Computation and Simulation, 8(3–4):311–312, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCS

[Goo79e] I. J. Good. C26. A comparison of some statistical estimates for the numbers of contingency tables. Journal of Statistical Computation and Simulation, 8(3–4): 312–314, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CIM

[Goo79f] I. J. Good. C28. The importance of mathematical consistency, at least in a discriminiation problem. Journal of Statistical Computation and Simulation, 8(3–4): 315–316, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCCa

[Goo79g] I. J. Good. C29. A comment on Cuthbert Daniel’s adjustment of sums of squares. Journal of Statistical Computation and Simulation, 8(3–4):316, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CPF

[Goo79h] I. J. Good. C38. “Proper fees” in multiple-choice examinations. Journal of Statistical Computation and Simulation, 9(1):73– 74, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CPX

[Goo79i] I. J. Good. C40. Partitioning of X^{2} and of the likelihood ratio for contingency tables. Journal of Statistical Computation and Simulation, 9(1):74–77, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCJ

[Goo79j] I. J. Good. C41. On the combination of judgements concerning quantiles of a distribution with potential application to the estimation of mineral resources. Journal of Statistical Computation and Simulation, 9(1):77– 79, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo79k].

Good:1979:CCCb

[Goo79k] I. J. Good. C42. Comment on C41. Journal of Statistical Computation and Simulation, 9(2): 159, 1979. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo79j].

Good:1979:CBM

[Goo79l] I. J. Good. C43. The Bayesian meaning of invariant estimation. Journal of Statistical Computation and Simulation, 9(2):160– 161, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CBB

[Goo79m] I. J. Good. C44. Bayes’s billiard-table argument extended to multinomials. Journal of Statistical Computation and Simulation, 9(2):161–163, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CII

[Goo79n] I. J. Good. C45. An index of influentiality of a scientist. Journal of Statistical Computation and Simulation, 9(2):163, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CVP

[Goo79o] I. J. Good. C46. A voting proposal: power-summing. Journal of Statistical Computation and Simulation, 9(2):164, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCP

[Goo79p] I. J. Good. C47. A comment on “proper fees” (C38). Journal of Statistical Computation and Simulation, 9(2):164–165, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CBE

[Goo79q] I. J. Good. C49. A Boolean error? Journal of Statistical Computation and Simulation, 9(3): 236–238, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comments [Goo80h, Goo81q].

Good:1979:CPS

[Goo79r] I. J. Good. C50. Predictive sample re-use and the estimation of probabilities. Journal of Statistical Computation and Simulation, 9(3):238–239, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CCRb

[Goo79s] I. J. Good. C52. The clustering of random variables. Journal of Statistical Computation and Simulation, 9(3):241–243, ???? 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CPC

[Goo79t] I. J. Good. C53. Partial correlation and spherical trigonometry. Journal of Statistical Computation and Simulation, 9(3):243– 245, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CMG

[Goo79u] I. J. Good. C54. Moments of a generalized gamma density. Journal of Statistical Computation and Simulation, 9(4):311, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CFM

[Goo79v] I. J. Good. C56. Fractional moments and cumulants: some unsolved problems. Journal of Statistical Computation and Simulation, 9(4):314–315, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CFG

[Goo79w] I. J. Good. C57. Further generallzations of X^{2}. Journal of Statistical Computation and Simulation, 9(4):315, 1979. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:CGX

[Goo79x] I. J. Good. C58. Generalizations of X^{2} for contingency tables. Journal of Statistical Computation and Simulation, 9(4): 315–316, 1979. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1979:EWCb

[Goo79y] I. J. Good. Early work on computers at Bletchley. Annals of the History of Computing, 1(1):38–48, July/September 1979. CODEN AHCOE5.

ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1979/pdf/a1038.pdf;   http:
//www.computer.org/annals/
an1979/a1038abs.htm.

Good:1979:EWCa

[Goo79z] I. J. Good. Early work on computers at Bletchley Park. Cryptologia, 3(2):65–77, April 1979. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http://

www.informaworld.com/smpp/
content~content=a741902783~
db=all~order=page.

Good:1979:ICM

[Goo79-27] I. J. Good. The impossibility of complete mutual observation. Physics Letters A, 70(2):81–82, 1979. CODEN PYLAAG. ISSN 0375-9601,1873-2429.

Good:1979:IDG

[Goo79-28] I. J. Good. The inversion of the discrete Gauss transform. Applicable Analysis, 9(3):205–218, 1979. CODEN APANCC. ISSN 0003-6811.

Good:1979:SHP

[Goo79-29] I. J. Good. Studies in the history of probability and statistics. XXXVII. A. M. Turing’s statistical work in World War II. Biometrika, 66(2):393– 396, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2335677.

Good:1980:AVB

[Goo80a] I. J. Good. An approximation of value in the Bayesian analyses of contingency tables. Journal of Statistical Computation and Simulation, 12(2):145–147, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:BRBa

[Goo80b] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 75 (369):239, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2287418.

Good:1980:BRBb

[Goo80c] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 75 (370):480, June 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2287498.

Good:1980:CNO

[Goo80d] I. J. Good. C59. The number of orderings of n candidates when ties and omissions are both allowed. Journal of Statistical Computation and Simulation, 10 (2):159, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo80j].

Good:1980:CPD

[Goo80e] I. J. Good. C64. The probability distribution of the gaps between the Mersenne primes. Journal of Statistical Computation and Simulation, 10(2):165, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CAR

[Goo80f] I. J. Good. C67. Another relationship between weight of evidence and errors of the first and second kinds. Journal of Statistical Computation and Simulation, 10(3–4):315–316, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CGC

[Goo80g] I. J. Good. C68. The grading of competitors. Journal of Statistical Computation and Simulation, 10(3–4):316–318, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CCCa

[Goo80h] I. J. Good. C70. Comment concerning C49. Journal of Statistical Computation and Simulation, 11(1):74, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo79q].

Good:1980:CDC

[Goo80i] I. J. Good. C71. Degrees of causation in regression analysis.

Journal of Statistical Computation and Simulation, 11(2):153– 155, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CCCb

[Goo80j] I. J. Good. C72. Comment concerning C59. Journal of Statistical Computation and Simulation, 11(2):155–156, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo80d].

Good:1980:CDS

[Goo80k] I. J. Good. C73. The diminishing significance of a p-value as the sample size increases. Journal of Statistical Computation and Simulation, 11(3–4):307–313, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CTV

[Goo80l] I. J. Good. C74. Terminology for various kinds of Bayesian methods. Journal of Statistical Computation and Simulation, 11(3–4): 309–313, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo83-30].

Good:1980:CFD

[Goo80m] I. J. Good. C77. Functions of distinct random vectors having identical distributions. Journal of Statistical Computation and Simulation, 12(1):68–70, ???? 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CGP

[Goo80n] I. J. Good. C78. On Godambe’s paradox. Journal of Statistical Computation and Simulation, 12(1):70–72, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CVL

[Goo80o] I. J. Good. C79. Vinograde’s lemma, singular decompositions and k-frames. Journal of Statistical Computation and Simulation, 12(1):72–74, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CFP

[Goo80p] I. J. Good. C80. Feynman’s path integrals and Sewall Wright’s path analysis. Journal of Statistical Computation and Simulation, 12(1):74–77, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CSC

[Goo80q] I. J. Good. C81. A simplification in “the causal calculus”. Journal of Statistical Computation and Simulation, 12(1):77– 78, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:CAF

[Goo80r] I. J. Good. C90. The autocorrelation functions ρ^{|r|^{c}}. Journal of Statistical Computation and Simulation, 12(2):148–151, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:GDG

[Goo80s] I. J. Good. Generalized determinants and generalized generalized variance. Journal of Statistical Computation and Simulation, 12(3–4):311–315, 1980. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1980:PWC

[Goo80t] I. J. Good. Pioneering work on computers at Bletchley. In Metropolis et al. [MHR80], pages 31–46. ISBN 0-12-491650-3, 1-4832-9668-7 (ebook). LCCN QA75.5 .I63

1976.  URL http://www.
sciencedirect.com/science/
article/pii/B9780124916500500125.

Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Good:1980:SHH

[Goo80u] I. J. Good. Some history of the hierarchical Bayesian methodology. In Bayesian statistics (Valencia, 1979), pages 489–519. University Press, Valencia, Spain, 1980. ISBN 84-370-0172-2. With discussion.

Good:1980:CJB

[Goo80v] Irving J. Good. The contributions of Jeffreys to Bayesian statistics. In Zellner [Zel80], pages 21– 34. ISBN 0-444-85270-0. LCCN HB139 .B39.

Good:1980:EMT

[Goo80w] Irving J. Good. Explicativity: a mathematical theory of explanation with statistical applications. In Zellner [Zel80], chapter 26, pages 397–426. ISBN 0-444-85270-0. LCCN HB139 .B39.

Good:1981:CCS

[Goo81a] I. J. Good. C100. Can scientific induction be meaningfully questioned? Journal of Statistical Computation and Simulation, 13 (2):154, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CDP

[Goo81b] I. J. Good. C101. Dynamic probability does not contradict reductio ad absurdum. Journal of Statistical Computation and Simulation, 13(2):154–155, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CEPa

[Goo81c] I. J. Good. C102. An error by Peirce concerning “weight of evidence”. Journal of Statistical Computation and Simulation, 13(2):155–157, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CEPb

[Goo81d] I. J. Good. C103. The effect of permutations of rows and columns on measures of association. Journal of Statistical Computation and Simulation, 13(3–4): 309–312, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CICa

[Goo81e] I. J. Good. C104. Informational correlations between two distributions. Journal of Statistical Computation and Simulation, 13(3–4):312–314, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CWS

[Goo81f] I. J. Good. C105. When should marginal cases be dropped in the interests of simplicity? Journal of Statistical Computation and Simulation, 13(3–4):314–315, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CICb

[Goo81g] I. J. Good. C106. International clinical trials. Journal of Statistical Computation and Simulation, 13(3–4):315, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CSCa

[Goo81h] I. J. Good. C107. Some comments about the Sherman–Foutz test. Journal of Statistical Computation and Simulation, 13(3–4): 316–319, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAC

[Goo81i] I. J. Good. C108. Analogue of chisquared that is powerful against bumpy alternatives. Journal of Statistical Computation and Simulation, 13(3–4):319–323, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CNM

[Goo81j] I. J. Good. C109. A new Monte Carlo computation for the characteristic function of Pearson’s chi-squared. Journal of Statistical Computation and Simulation, 13(3–4):323–324, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CRC

[Goo81k] I. J. Good. C110. Randomly connected genetic nets. Journal of Statistical Computation and Simulation, 13(3–4):324–327, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CSD

[Goo81l] I. J. Good. C114. On the statistical determination of authorship. Journal of Statistical Computation and Simulation, 14(1): 55–60, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAE

[Goo81m] I. J. Good. C115. Assessment of eye-ball descriptions. Journal of Statistical Computation and Simulation, 14(1):60–63, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CSCb

[Goo81n] I. J. Good. C116. A supplement to chi-squared when testing homogeneity. Journal of Statistical Computation and Simulation, 14(1):63–67, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CMCb

[Goo81o] I. J. Good. C117. A Monte Carlo method that will be impracticable before the year 2001. Journal of Statistical Computation and Simulation, 14(1):67–69, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CFC

[Goo81p] I. J. Good. C119. The fast calculation of the exact distribution of Pearson’s chi-squared and of the number of repeats within the cells of a multinomial by using a Fast Fourier Transform. Journal of Statistical Computation and Simulation, 14(1):71–78, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CCC

[Goo81q] I. J. Good. C84. A comment on C49, “A Boolean error”. Journal of Statistical Computation and Simulation, 12(2):137–138, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo79q].

Good:1981:CRP

[Goo81r] I. J. Good. C87. Roughness penalties, invariant under rotation for multidimensional probability density estimation. Journal of Statistical Computation and Simulation, 12(2):142–144, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CAV

[Goo81s] I. J. Good. C88. An approximation of value in the Bayesian analyses of contingency tables. Journal of Statistical Computation and Simulation, 12(2):145– 147, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CTC

[Goo81t] I. J. Good. C89. The total cost of collecting income tax. Journal of Statistical Computation and Simulation, 12(2):147–148, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CGDa

[Goo81u] I. J. Good. C91. Generalized determinants and generalized variance. Journal of Statistical Computation and Simulation, 12(3–4): 311–313, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CN

[Goo81v] I. J. Good. C92. Neologisms.

Journal of Statistical Computation and Simulation, 12(3–4): 313, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CWP

[Goo81w] I. J. Good. C94. When is g positive in the mixed Dirichlet approach to contingency tables? Journal of Statistical Computation and Simulation, 13(1):49– 52, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CMCa

[Goo81x] I. J. Good. C95. The Monte Carlo computation of Bayes factors for contingency tables. Journal of Statistical Computation and Simulation, 13(1):52–56, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CWE

[Goo81y] I. J. Good. C96. The weight of evidence provided by uncertain testimony or from an uncertain event. Journal of Statistical Computation and Simulation, 13(1): 56–60, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CGDb

[Goo81z] I. J. Good. C97. Generalized determinants and generalized Jacobians. Journal of Statistical Computation and Simulation, 13(1): 60–62, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CFP

[Goo81-27] I. J. Good. C98. Flexible priors for estimating a normal distribution. Journal of Statistical Computation and Simulation, 13(2): 151–153, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:CPS

[Goo81-28] I. J. Good. C99. The probability of a scientific law. Journal of Statistical Computation and Simulation, 13(2):153, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:EA

[Goo81-29] I. J. Good. Errata and annenda. Journal of Statistical Computation and Simulation, 14 (1):79, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1981:EEM

[Goo81-30] I. J. Good. Erratum: “Explicativity: a mathematical theory of explanation with statistical applications” [Proc. Roy. Soc. London Ser. A 354 (1977), no. 1678, 303– 330; MR 58 #24636]. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 377(1771):504, 1981. CODEN PRLAAZ. ISSN 0962-8444,2053-9169.

Good:1981:PSS

[Goo81-31] I. J. Good. Problems and solutions: Solutions of advanced problems: 6137. American Mathematical Monthly, 88(3): 215, March 1981. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77, GR78, GB80].

Good:1981:WEP

[Goo81-32] I. J. Good. The weight of evidence provided by uncertain testimony or from an uncertain event. Journal of Statistical Computation and Simulation, 13(1): 56–60, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:BRN

[Goo82a] I. J. Good. Book review:

Nonparametric Probability Density Estimation (Richard A. Tapia and James R. Thompson). SIAM Review, 24(3):364–365, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1982:CPH

[Goo82b] I. J. Good. C121. The postulate of honest judgement. Journal of Statistical Computation and Simulation, 14(3–4):321, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPN

[Goo82c] I. J. Good. C122. A paradox in the Neyman–Pearson theory. Journal of Statistical Computation and Simulation, 14(3– 4):322, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CMS

[Goo82d] I. J. Good. C123. Are the methods of statistical mechanics applicable to society? Journal of Statistical Computation and Simulation, 14(3–4):322–323, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPD

[Goo82e] I. J. Good. C124. Probability densities versus discriminant functions. Journal of Statistical Computation and Simulation, 14(3–4):321–323, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPP

[Goo82f] I. J. Good. C126. PROTOM: a projection-tomographic method for density estimation from multivariate data. Journal of Statistical Computation and Simulation, 15(1):73–75, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CWM

[Goo82g] I. J. Good. C127. When is maximum likelihood estimation O.K. for a Bayesian? Journal of Statistical Computation and Simulation, 15(1):75–77, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CISa

[Goo82h] I. J. Good. C129. An index of separateness of clusters and a permutation test for its statistical significance. Journal of Statistical Computation and Simulation, 15(1):81–84, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CBL

[Goo82i] I. J. Good. C130. Bayesian likelihood: a point of terminology. Journal of Statistical Computation and Simulation, 15(1): 84–85, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CMR

[Goo82j] I. J. Good. C133. On the measurement of regret. Journal of Statistical Computation and Simulation, 15(2–3):237–238, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CPI

[Goo82k] I. J. Good. C134. A proposal for the improvement of the refereeing process. Journal of Statistical Computation and Simulation, 15(2–3):238, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CIB

[Goo82l] I. J. Good. C135. Interactions between indexes of separateness. Journal of Statistical Computation and Simulation, 15(2–3): 239–242, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CISb

[Goo82m] I. J. Good. C136. The irregular shapes of polypeptide chains. Journal of Statistical Computation and Simulation, 15(2–3): 243–247, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CFC

[Goo82n] I. J. Good. C138. The fast calculation of the exact distribution of Pearson’s chi-squared and of the number of repeats within the cells of a multinomial by using a Fast Fourier Transform: an addendum. Journal of Statistical Computation and Simulation, 15(4): 336–337, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CSE

[Goo82o] I. J. Good. C139. The standard error of the estimated “coverage” of a sample of species or vocabulary. Journal of Statistical Computation and Simulation, 15 (4):337, 1982. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CST

[Goo82p] I. J. Good. C140. Standardized tail-area probabilities. Journal of Statistical Computation and Simulation, 16(1):65–66, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:CNF

[Goo82q] I. J. Good. Corrigenda: “A new formula for cumulants” [Math. Proc. Cambridge Philos. Soc. 78 (1975), no. 2, 333–337; MR 52 #6957]. Mathematical Proceedings of the Cambridge Philosophical Society, 91(2):343, 1982. CO-DEN MPCPCO. ISSN 0305-0041,1469-8064.

Good:1982:CAP

[Goo82r] I. J. Good. The cumulants of an analogue of Pearson’s chisquared. Journal of Statistical Computation and Simulation, 15(2–3):171–181, 1982. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1982:DCM

[Goo82s] I. J. Good. Diversity as a concept and its measurement: Comment. Journal of the American Statistical Association, 77(379): 561–563, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2287710.

Good:1982:GEE

[Goo82t] I. J. Good. A good explanation of an event is not necessarily corroborated by the event. Philosophy of Science, 49(2):251–253, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://

                                        www.jstor.org/stable/186922.

Good:1982:LPC

[Goo82u] I. J. Good. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77 (378):342–344, June 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2287248.

Good:1983:BRB

[Goo83a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 78(384): 987–989, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

                                        URL https://www.jstor.org/
                                        stable/2288219.

Good:1983:CSIa

[Goo83b] I. J. Good. C143. Scientific induction: “universal and predictive”. Journal of Statistical Computation and Simulation, 16(3–4): 311–312, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CDS

[Goo83c] I. J. Good. C144. The diminishing significance of a fixed p-value as the sample size increases: a discrete model. Journal of Statistical Computation and Simulation, 16(3–4):312–3 14, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CMA

[Goo83d] I. J. Good. C145. A measure of adhockery. Journal of Statistical Computation and Simulation, 16(3–4):314, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCB

[Goo83e] I. J. Good. C146. Correlation between power functions. Journal of Statistical Computation and Simulation, 16(3–4):314–316, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPE

[Goo83f] I. J. Good. C148. Probability estimation by maximum penalized likelihood for large contingency tables and for other categorical data. Journal of Statistical Computation and Simulation, 17(1): 66–67, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CSC

[Goo83g] I. J. Good. C149. A simple consequence of Daniel Bernoulli’s logarithmic formula for utility. Journal of Statistical Computation and Simulation, 17(1):67– 68, 1983. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CDP

[Goo83h] I. J. Good. C150. “Density” a point of terminology. Journal of Statistical Computation and Simulation, 17(1):68, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CA

[Goo83i] I. J. Good. C151. Antisurprise. Journal of Statistical Computation and Simulation, 17(1):69– 71, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CND

[Goo83j] I. J. Good. C152. On the number of duplicated fingerprints. Journal of Statistical Computation and Simulation, 17(1):71– 74, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCA

[Goo83k] I. J. Good. C154. A conjecture about the history of errors of the first and second kinds. Journal of Statistical Computation and Simulation, 17(2):155, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CIA

[Goo83l] I. J. Good. C157. An improved algorithm for the fast calculation of the exact distribution of Pearson’s chi-squared. Journal of Statistical Computation and Simulation, 17(3):236–242, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CNE

[Goo83m] I. J. Good. C158. A note on the extraction of coefficients from power series. Journal of Statistical Computation and Simulation, 17(3):243–244, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CFCa

[Goo83n] I. J. Good. C160. A further comment on multiple-choice examinations. Journal of Statistical Computation and Simulation, 17(4): 311–313, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CWF

[Goo83o] I. J. Good. C161. When are free observations of positive expected value. Journal of Statistical Computation and Simulation, 17(4): 313–315, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CMC

[Goo83p] I. J. Good. C162. Moments and cumulants of weights of evidence. Journal of Statistical Computation and Simulation, 17(4):315– 319, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CME

[Goo83q] I. J. Good. C163. The moments of an exponential. Journal of Statistical Computation and Simulation, 17(4):320–321, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCL

[Goo83r] I. J. Good. C164. The cumulants of the lognormal distribution, including some conjectures. Journal of Statistical Computation and Simulation, 17(4):321– 328, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CCCa

[Goo83s] I. J. Good. C165. A correction concerning my interpretation of Peirce, and the Bayesian interpretation of Neyman–Pearson “hypothesis determination. Journal of Statistical Computation and Simulation, 18(1):71–84, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CMW

[Goo83t] I. J. Good. C166. The moments of weight of evidence gives intermediate hypotheses. Journal of Statistical Computation and Simulation, 18(1):74–79, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPO

[Goo83u] I. J. Good. C167. The probability of odd partial quotients in continued fractions. Journal of Statistical Computation and Simulation, 18(1):79–80, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CFD

[Goo83v] I. J. Good. C168. Fractional dimensions, continued fractions, and Markovity. Journal of Statistical Computation and Simulation, 18(1):80–84, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See erratum [Goo83-28].

Good:1983:CBNa

[Goo83w] I. J. Good. C169. Barndorff– Nielsen’s plausibility function and a principle of least surprise. Journal of Statistical Computation and Simulation, 18(2–3): 215–218, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPM

[Goo83x] I. J. Good. C170. A potential misstatement of Bayes’s theorem. Journal of Statistical Computation and Simulation, 18(2–3): 218–219, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CFCb

[Goo83y] I. J. Good. C171. Further combinatorial identities related to the cumulants of a lognormal distribution. Journal of Statistical Computation and Simulation, 18(2–3):219–226, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CPR

[Goo83z] I. J. Good. C172. Probabilities of residues for partial quotients of continued fractions, and a question about moments. Journal of Statistical Computation and Simulation, 18(2–3):226–227, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See erratum [Goo89o].

Good:1983:CGD

[Goo83-27] I. J. Good. C173. A geometry of distributions. Journal of Statistical Computation and Simulation, 18(2–3):227–229, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:CSIb

[Goo83-28] I. J. Good. C174. Scientific induction, Jeffreys’s and my approaches compared. Journal of Statistical Computation and Simulation, 18(2–3):229–231, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comments [Goo83-29, Goo83-31].

Good:1983:CHB

[Goo83-29] I. J. Good. C175. Hierarchical Bayes, mixed Dirichlet priors, penalized likelihood, etc. (see the first paragraph of C174.). Journal of Statistical Computation and Simulation, 18(2–3): 231–234, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83-28]. This note was my response to comments by Tom Leonard.

Good:1983:CBNb

[Goo83-30] I. J. Good. C176. The Bayes/ non-Bayes compromise or synthesis and Box’s current philosophy (see the first paragraph of C74.). Journal of Statistical Computation and Simulation, 18(2–3): 234–236, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo80l].

Good:1983:CDI

[Goo83-31] I. J. Good. C177. The device of imaginary results, “WHAT IF”, and preposterior analysis (see the first paragraph of C174). Journal of Statistical Computation and Simulation, 18(2–3): 236, 1983. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83-28].

Good:1983:CTT

[Goo83-32] I. J. Good. C178. Tail trouble when sampling. Journal of Statistical Computation and Simulation, 18(2–3):237, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1983:PED

[Goo83-33] I. J. Good. The philosophy of exploratory data analysis. Philosophy of Science, 50(2):283–295, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://

                                        www.jstor.org/stable/188015.

Good:1983:RHM

[Goo83-34] I. J. Good. The robustness of a hierarchical model for multinomials and contingency tables. In Scientific inference, data analysis, and robustness (Madison, Wisconsin, 1981), volume 48 of Publications of the Mathematical Research Center, University of Wisconsin, pages 191–211. Academic Press, New York, NY, USA, 1983. ISBN 0-12-121160-6.

Good:1983:GTF

[Goo83-35] Irving John Good. Good Thinking: the Foundations of Probability and Its Applications. University of Minnesota Press, Minneapolis, MN, USA, 1983. ISBN 0-8166-1141-6, 0-8166-1142-4 (paperback). xviii + 332 pp. LCCN QA273.4 .G66 1983; QA273.4 .G66 1983. URL http:

                                        //site.ebrary.com/lib/yale/
                                        Doc?id=10231244.

Good:1984:BAP

[Goo84a] I. J. Good. A Bayesian approach in the philosophy of inference. British Journal for the Philosophy of Science, 35(2):161–166, June 1984. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/35/2/161.full.pdf+
html;   https://www.jstor.
org/stable/686966.

Good:1984:BRB

[Goo84b] I. J. Good. Book review:

The Fractal Geometry of Nature, by Benoît B. Mandelbrot. SIAM Review, 26(1):131–132, January 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//link.aip.org/link/siread/
v26/i1/p131/s1;   https://
www.jstor.org/stable/2029696.

Good:1984:BRF

[Goo84c] I. J. Good. Book review:

The Fractal Geometry of Nature, by Benoît B. Mandelbrot. SIAM Review, 26(1):131–132, January 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//link.aip.org/link/siread/
v26/i1/p131/s1;   https://
www.jstor.org/stable/2029696.

Good:1984:CRR

[Goo84d] I. J. Good. C186. The repeat rate with respect to a given multinomial population, and the Dirichlet–multinomial, distribution. Journal of Statistical Computation and Simulation, 19(1): 85–89, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CLE

[Goo84e] I. J. Good. C187. Laplace’s estimate of the mass of Jupiter.

Journal of Statistical Computation and Simulation, 19(1):89– 90, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CCD

[Goo84f] I. J. Good. C189. A commonsense design for multiple comparisons. Journal of Statistical Computation and Simulation, 19 (1):94, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CHS

[Goo84g] I. J. Good. C191. How should tail-area probabilities be standardized for sample size in unpaired comparisons? Journal of Statistical Computation and Simulation, 19(2):174, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:COT

[Goo84h] I. J. Good. C192. One tail versus two-tails, and the harmonicmean rule of thumb. Journal of Statistical Computation and Simulation, 19(2):174–176, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPV

[Goo84i] I. J. Good. C193. Paired versus unpaired comparisons and the Hakvunic-mean rule of thumb. Journal of Statistical Computation and Simulation, 19(2):176– 177, 1984. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CEC

[Goo84j] I. J. Good. C195. Editorial comment on Kempthorne’s C194. Journal of Statistical Computation and Simulation, 19(2):182– 184, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CBE

[Goo84k] I. J. Good. C197. The best explicatum for weight of evidence. Journal of Statistical Computation and Simulation, 19(4):294– 299, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPT

[Goo84l] I. J. Good. C198. A probability too small to achieve using a computer. Journal of Statistical Computation and Simulation, 19(4): 299–300, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CST

[Goo84m] I. J. Good. C199. Standardized tail-area probabilities: a possible interpretation. Journal of Statistical Computation and Simulation, 19(4):299–300, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CTB

[Goo84n] I. J. Good. C200. The tolerant Bayesian’s interpretation of a tail-area probability. Journal of Statistical Computation and Simulation, 19(4):300–302, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CBI

[Goo84o] I. J. Good. C201. A Bayesian interpretation of ancillarity. Journal of Statistical Computation and Simulation, 19(4):302–308, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPA

[Goo84p] I. J. Good. C202. Probability applied to Diophantine equations. Journal of Statistical Computation and Simulation, 19(4):308– 313, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90].

Good:1984:CQD

[Goo84q] I. J. Good. C203. A quick and dirty method for nonparametric estimation of a density function. Journal of Statistical Computation and Simulation, 19(4):313– 314, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CEH

[Goo84r] I. J. Good. C204. The early history of the Fisher–Yates–Irwin formula and Fisher’s ’exact test’. Journal of Statistical Computation and Simulation, 19(4):315– 319, 1984. CODEN JSCSAJ.

ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CMP

[Goo84s] I. J. Good. C205. Monotonic properties of the moments of a Bayes factor and the relationship to measures of divergence. Journal of Statistical Computation and Simulation, 19(4):320– 325, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CFN

[Goo84t] I. J. Good. C208. A further note on the early history of the Fisher– Yates–Irwin formula. Journal of Statistical Computation and Simulation, 20(2):155–159, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CEN

[Goo84u] I. J. Good. C209. An error by Neyman noticed by Dickey. Journal of Statistical Computation and Simulation, 20(2):159– 160, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:COJ

[Goo84v] I. J. Good. C210. The onlyjust effect in nonparametric procedures, and the achievement of objectivity by using a statistician’s stooge. Journal of Statistical Computation and Simulation, 20(2):160–164, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CSH

[Goo84w] I. J. Good. C213. A sharpening of the harmonic-mean rule of thumb for combining tests “in parallel”. Journal of Statistical Computation and Simulation, 20(2):173– 176, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CHM

[Goo84x] I. J. Good. C214. The harmonicmean rule of thumb: some classes of applications. Journal of Statistical Computation and Simulation, 20(2):176–179, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPL

[Goo84y] I. J. Good. C215. Penalized least surprise. Journal of Statistical Computation and Simulation, 20(2):179–180, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CCC

[Goo84z] I. J. Good. Comment concerning the conjectured value of a determinant occurring in statistics. Journal of Statistical Computation and Simulation, 19(1): 95, 1984. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:RBF

[Goo84-27] I. J. Good. Review: The Fractal Geometry of Nature, by Benoît B. Mandelbrot. SIAM Review, 26 (1):131–132, January 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL https://www.jstor.org/
stable/2029696.

Good:1984:TC

[Goo84-28] I. J. Good. Turing and the computer. Nature, 307(5952):663– 664, February 16, 1984. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/
nature/journal/v307/n5952/
pdf/307663a0.pdf.

Good:1985:BRB

[Goo85a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 80(392): 1067–1068, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2288583.

Good:1985:CPI

[Goo85b] I. J. Good. C216. Probabilistic induction is inevitable. Journal of Statistical Computation and Simulation, 20(4):323–324, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CTA

[Goo85c] I. J. Good. C217. Tail-area probabilities and Bayes factors as “distances” from the null hypothesis. Journal of Statistical Computation and Simulation, 20(4): 325, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CCC

[Goo85d] I. J. Good. C218. A correction concerning degrees of causation. Journal of Statistical Computation and Simulation, 20(4): 326, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CNN

[Goo85e] I. J. Good. C220. The number of numerical multiplications when multiplying polynomials by recursive bisection. Journal of Statistical Computation and Simulation, 20(4):329–331, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CAS

[Goo85f] I. J. Good. C223. The accuracy of Stirling’s formula for x! when x is small. Journal of Statistical Computation and Simulation, 21(1): 84–86, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CHN

[Goo85g] I. J. Good. C224. A historical note regarding the mass of Jupiter. Journal of Statistical Computation and Simulation, 21(1):86–87, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CWD

[Goo85h] I. J. Good. C225. Who discovered the God–Koheleth law?

Journal of Statistical Computation and Simulation, 21(1):87– 88, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CNM

[Goo85i] I. J. Good. C226. A new measure of surprise. Journal of Statistical Computation and Simulation, 21(1):88–89, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CHC

[Goo85j] I. J. Good. C227. A historical comment concerning the index of separateness. Journal of Statistical Computation and Simulation, 21(1):89–90, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CNF

[Goo85k] I. J. Good. C228. The number of functionals: a combinatorial curiosity. Journal of Statistical Computation and Simulation, 21 (1):90, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CPDa

[Goo85l] I. J. Good. C229. Probability density estimation over subintervals of the range. Journal of Statistical Computation and Simulation, 21(1):90–91, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CSW

[Goo85m] I. J. Good. C232. Sufftctmcy and weight of evidence. Journal of Statistical Computation and Simulation, 21(3–4):334–336, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CFS

[Goo85n] I. J. Good. C233. Fisher and subjectivity. Journal of Statistical Computation and Simulation, 21(3–4):336–337, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CDL

[Goo85o] I. J. Good. C234. In defense of the law of the excluded middle, or all logic is probabilistic. Journal of Statistical Computation and Simulation, 21(3–4): 337–339, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CAA

[Goo85p] I. J. Good. C235. Another application of the Wilcoxon signedrank test. Journal of Statistical Computation and Simulation, 21(3–4):339, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CMN

[Goo85q] I. J. Good. C236. Mathematically natural generalizations of some measures of association and dependence for contingency tables. Journal of Statistical Computation and Simulation, 22(1): 83–91, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CEP

[Goo85r] I. J. Good. C238. Explanatory power depends on more than probabilities. Journal of Statistical Computation and Simulation, 22(2):184–186, 1985. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CED

[Goo85s] I. J. Good. C241. The exact distribution of a statistic for matched n-tuples allowing for ties. Journal of Statistical Computation and Simulation, 23(1–2): 142–146, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CPDb

[Goo85t] I. J. Good. C244. The probability distribution of the gaps between the Mersenne primes, II. Journal of Statistical Computation and Simulation, 23(1–2): 154–156, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo86d].

Good:1985:DTI

[Goo85u] I. J. Good. Discussion: Testing for independence in a twoway table: New interpretations of the chi-square statistic. Annals of Statistics, 13(3):

885–886, September 1985. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.aos/1176349638.

Good:1985:HCC

[Goo85v] I. J. Good. A historical comment concerning novel confirmation. British Journal for the Philosophy of Science, 36(2):184–185, June 1985. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

                                        //bjps.oxfordjournals.org/
                                        content/36/2/184.full.pdf+
                                        html;   https://www.jstor.
                                        org/stable/687034.

Good:1985:WEB

[Goo85w] I. J. Good. Weight of evidence: a brief survey. In Bayesian statistics, 2 (Valencia, 1983), pages 249–269. North-Holland, Amsterdam, The Netherlands, 1985. ISBN 0-444-87746-0. With discussion and a reply by the author.

Good:1986:ASC

[Goo86a] I. J. Good. An addendum: “Skew circulants and the theory of numbers”. Fibonacci Quarterly, 24(2):176–177, 1986. CO-DEN FIBQAU. ISSN 0015-0517.

Good:1986:BRS

[Goo86b] I. J. Good. Book review: Statistical Inference Based on Ranks (Thomas P. Hettmansperger). SIAM Review, 28(4):587–588, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1986:CVS

[Goo86c] I. J. Good. C246. Very small tails of the t distribution, and significance tests for clustering. Journal of Statistical Computation and Simulation, 23(3):243– 248, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CCCa

[Goo86d] I. J. Good. C248. Comment on C244. Journal of Statistical Computation and Simulation, 23(3): 250, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo85t].

Good:1986:CSM

[Goo86e] I. J. Good. C250. “Saturated model” or “quasimodel”, a point of terminology. Journal of Statistical Computation and Simulation, 24(2):168–169, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CSB

[Goo86f] I. J. Good. C260. Some Bayes/ non-Bayes compromises based on covert Bayesianism. Journal of Statistical Computation and Simulation, 25(3–4):298–300, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CSS

[Goo86g] I. J. Good. C261. The speciessam pling problem again. Journal of Statistical Computation and Simulation, 25(3–4):301– 304, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CGA

[Goo86h] I. J. Good. C264. A geometrical approach to the amalgamation paradox. Journal of Statistical Computation and Simulation, 26(1–2):129–132, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CCI

[Goo86i] I. J. Good. C265. The computerintensive form of a Bayes/non-Bayes compromise. Journal of Statistical Computation and Simulation, 26(1–2):132–133, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CRP

[Goo86j] I. J. Good. C266. Recognizability, permutability, and homogeneity. Journal of Statistical Computation and Simulation, 26(1–2): 133–134, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CME

[Goo86k] I. J. Good. C267. The 3d monocular effect. Journal of Statistical Computation and Simulation, 26(1–2):134–139, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CLR

[Goo86l] I. J. Good. C268. The likelihoodratio approach to the generalized Behrens–Fisher problem: oneway Anova under heteroscedasticity. Journal of Statistical Computation and Simulation, 26(1–2): 139–145, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CND

[Goo86m] I. J. Good. C269. The newly discovered poem, possibly by Shakespeare. Journal of Statistical Computation and Simulation, 26(1–2):145–146, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CNP

[Goo86n] I. J. Good. C270. Newcomb’s problem and the foundations of probability. Journal of Statistical Computation and Simulation, 26(1–2):127–148, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CAE

[Goo86o] I. J. Good. C271. Age of earliest memory: an informal sampling experiment. Journal of Statistical Computation and Simulation, 26(3–4):299–301, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CFB

[Goo86p] I. J. Good. C272. A flexible Bayesian model for comparing two treatments. Journal of Statistical Computation and Simulation, 26(3–4):301–305, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CHC

[Goo86q] I. J. Good. C273. A historical comment concerning the Monte Carlo estimation of pvalues. Journal of Statistical Computation and Simulation, 26(3–4):306, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CMP

[Goo86r] I. J. Good. C274. Multinomial prediction and improved approximations to the shrinking parameter that maximizes a Dirichlet–multinomial likelihood. Journal of Statistical Computation and Simulation, 26(3–4): 306–311, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CMB

[Goo86s] I. J. Good. C275. The maximum of a Bayes factor against “independence” in a contingency table, and generalizations to higher dimensions. Journal of Statistical Computation and Simulation, 26(3–4):312–316, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CTM

[Goo86t] I. J. Good. C276. On terminology for matrices of special forms. Journal of Statistical Computation and Simulation, 26(3–4): 316–317, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1986:CCCb

[Goo86u] I. J. Good. Comments, conjectures, and conclusions C258. Editorial note on C257 regarding the t-test. Journal of Statistical Computation and Simulation, 25(3–4): 296–297, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [LS86].

Good:1986:ND

[Goo86v] I. J. Good. (Neo)^{n}-Darwinism.

Physica. D, Nonlinear Phenomena, 22(1–3):13–30, 1986. CO-DEN PDNPDT. ISSN 0167-2789,1872-8022. Evolution, games and learning (Los Alamos, N.M., 1985).

Good:1986:SCTa

[Goo86w] I. J. Good. Skew circulants and the theory of numbers. Fibonacci Quarterly, 24(1):47–60, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/24-
1/good.pdf.

Good:1986:SCTb

[Goo86x] I. J. Good. Skew circulants and the theory of numbers: An addendum. Fibonacci Quarterly, 24(2):176–177, May 1986. CO-DEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/
Scanned/24-2/good.pdf.

Good:1986:SSAa

[Goo86y] I. J. Good. Some statistical applications of Poisson’s work. Statistical Science, 1(2):157–170, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013690.  See [DE86,

Sol86, Hey86, Sin86] and rejoinder [Goo86z].

Good:1986:SSAb

[Goo86z] I. J. Good. [Some statistical applications of Poisson’s work]: Rejoinder. Statistical Science, 1(2):179–180, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013695.  See [DE86,

Sol86, Hey86, Sin86, Goo86y].

Good:1986:ASP

[Goo86-27] I. J. Good. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):346–347, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013612.

Good:1987:BRB

[Goo87a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 82 (398):697–698, June 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2289507.

Good:1987:CSC

[Goo87b] I. J. Good. C277. Some checks on the consistency of probability judgments and a conjectured method for improving those judgments. Journal of Statistical Computation and Simulation, 28(1):63–71, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CIE

[Goo87c] I. J. Good. C278. Indices of economic concentration. Journal of Statistical Computation and Simulation, 28(1):71–72, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CCN

[Goo87d] I. J. Good. C283. The calculation of the normal or error function with normally acceptable proportional error. Journal of Statistical Computation and Simulation, 28(1):83–84, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CC

[Goo87e] I. J. Good. C284. Confidence. Journal of Statistical Computation and Simulation, 28(1):85– 87, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CSU

[Goo87f] I. J. Good. C285. A survey of the use of the Fast Fourier Transform for computing distributions. Journal of Statistical Computation and Simulation, 28(1):87– 93, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CCL

[Goo87g] I. J. Good. C286. Correlation and linearity: some lapses in exposition. Journal of Statistical Computation and Simulation, 28(1): 93–94, 1987. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CCC

[Goo87h] I. J. Good. C288. A correction concerning ’ancillarity’ for significance tests. Journal of Statistical Computation and Simulation, 28(2):182–183, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CMP

[Goo87i] I. J. Good. C289. Maximum penalized likelihood for somewhat parametric estimation. Journal of Statistical Computation and Simulation, 28(3):261–264, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CED

[Goo87j] I. J. Good. C294. An example of dynamic probability in number theory. Journal of Statistical Computation and Simulation, 28(4):349–350, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:DRR

[Goo87k] I. J. Good. Discussion: a reinstatement, in response to Gillies, of Redhead’s argument in support of induction. Philosophy of Science, 54(3):470–472, 1987. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Good:1987:FGV

[Goo87l] I. J. Good. F6. Graig the vague. Journal of Statistical Planning and Inference, 16(??): 262–263, ???? 1987. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0378375887900772.

Good:1987:FSI

[Goo87m] I. J. Good. F7. Some improvements in statistical and other ‘literature’. Journal of Statistical Planning and Inference, 16 (??):263–265, ???? 1987. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0378375887900784.

Good:1987:FAS

[Goo87n] I. J. Good. F8. Is the arc-sine law that counter-intuitive? Journal of Statistical Planning and Inference, 16(??):429–430, ???? 1987. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0378375887900930.

Good:1987:RRG

[Goo87o] I. J. Good. A reinstatement, in response to Gillies, of Redhead’s argument in support of induction. Philosophy of Science, 54(3):470–472, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

                                        tronic).  URL https://www.
                                        jstor.org/stable/187586. See

[Gil86].

Good:1987:SUD

[Goo87p] I. J. Good. Serial unblocked designs balanced for correlation between neighbours. Communications in Statistics: Theory and Methods, 16(4):1153–1159, 1987. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1987:TPN

[Goo87q] I. J. Good. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82 (397):125–128, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2289133.

Good:1988:CBR

[Goo88a] I. J. Good. C297. Bayes’s red billiard ball is also a herring, and why Bayes withheld publication. Journal of Statistical Computation and Simulation, 29(4):335– 340, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CIS

[Goo88b] I. J. Good. C299. An inequality for the smallest singular value of a matrix products. Journal of Statistical Computation and Simulation, 29(4):343–346, 1988. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CWS

[Goo88c] I. J. Good. C300. What should be the usual shape of a loss function? Journal of Statistical Computation and Simulation, 29(4): 346–347, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CER

[Goo88d] I. J. Good. C302. An early risk assessment for an atomic reactor. Journal of Statistical Computation and Simulation, 29(4): 351, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:CPI

[Goo88e] I. J. Good. C303. The pragmatic irrelevance of De Finetti’s theorem. Journal of Statistical Computation and Simulation, 29(4): 351–353, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See apology [Goo89r].

Good:1988:CNC

[Goo88f] I. J. Good. C305. The number of correct significant digits in a piece of physical numerology. Journal of Statistical Computation and Simulation, 29(4): 358–359, 1988. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90].

Good:1988:ISS

[Goo88g] I. J. Good. An inequality for the smallest singular value of a matrix product. Journal of Statistical Computation and Simulation, 29(4):343–346, 1988. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1988:IBSa

[Goo88h] I. J. Good. The interface between statistics and philosophy of science. Statistical Science, 3(4):386–397, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        ss/1177012754. See [Goo88j].

Good:1988:SGA

[Goo88i] I. J. Good. A simple generalization of analytic function theory. Exposition. Math., 6(4):289–311, 1988. ISSN 0723-0869 (print), 1878-0792 (electronic).

Good:1988:IBSb

[Goo88j] I. J. Good. [The interface between statistics and philosophy of science]: Rejoinder. Statistical Science, 3(4):406–412, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177012759. See [Goo88h].

Good:1989:CCP

[Goo89a] I. J. Good. C311. On the combination of pieces of evidence. Journal of Statistical Computation and Simulation, 31(1):54– 58, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90].

Good:1989:CYA

[Goo89b] I. J. Good. C312. Yet another argument for the explicatum of weight of evidence. Journal of Statistical Computation and Simulation, 31(1):58–59, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CRD

[Goo89c] I. J. Good. C313. The rate of diagnosing a fault, for example, in a communication channel. Journal of Statistical Computation and Simulation, 31(1): 60–61, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CCN

[Goo89d] I. J. Good. C314. Chaotics and number theory. Journal of Statistical Computation and Simulation, 31(1):61–62, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CSF

[Goo89e] I. J. Good. C315. A suspicious feature of the Popper/Miller argument. Journal of Statistical Computation and Simulation, 31(1):62–63, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CIL

[Goo89f] I. J. Good. C316. Interpretation of a large coefficient of determination. Journal of Statistical Computation and Simulation, 31(1): 63–64, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CTC

[Goo89g] I. J. Good. C317. The theorem of corroboration and undermining, and Popper’s demarcation rule. Journal of Statistical Computation and Simulation, 31(2): 119–120, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CPC

[Goo89h] I. J. Good. C318. The paradox of confirmation and stoogian observations. Journal of Statistical Computation and Simulation, 31 (2):121, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CWE

[Goo89i] I. J. Good. C319. Weight of evidence and a compelling metaprinciple. Journal of Statistical Computation and Simulation, 31(2): 121–123, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CWA

[Goo89j] I. J. Good. C320. Weak attacks on transitivity of preference. Journal of Statistical Computation and Simulation, 31(2): 123–124, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CNN

[Goo89k] I. J. Good. C321. The need for a normative theory of rationality. Journal of Statistical Computation and Simulation, 31(2): 124–126, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CMV

[Goo89l] I. J. Good. C322. Medialized p-values. Journal of Statistical Computation and Simulation, 31(2):126–127, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CTS

[Goo89m] I. J. Good. C323. A test for a specified Markov model or branching process. Journal of Statistical Computation and Simulation, 31(2):127–128, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CEC

[Goo89n] I. J. Good. C324. Erratum to C168 on fractional dimensions, continued fractions and Markovity. Journal of Statistical Computation and Simulation, 31 (2):128, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83v].

Good:1989:CEA

[Goo89o] I. J. Good. C325. Erratum and addendum to C172. Journal of Statistical Computation and Simulation, 31(2):128–129, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo83z].

Good:1989:CAFa

[Goo89p] I. J. Good. C327. Addenda for “The fractional dimensional theory of continued fractions”, Proc. Cambridge Philos. Soc. 37 (1941), 199–228. Journal of Statistical Computation and Simulation, 32(1–2):64–68, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CVE

[Goo89q] I. J. Good. C330. The variability of evaluations of speculative ideas. Journal of Statistical Computation and Simulation, 32(1–2): 81–88, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CAFb

[Goo89r] I. J. Good. C331. An apology to De Finetti. Journal of Statistical Computation and Simulation, 32(1–2):88–90, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Goo88e].

Good:1989:CSI

[Goo89s] I. J. Good. C332. Surprise indexes and p-values. Journal of Statistical Computation and Simulation, 32(1–2):90–92, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CJN

[Goo89t] I. J. Good. C333. On the judgement of “numerological” assertions. Journal of Statistical Computation and Simulation, 32(1–2): 92–103, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CNP

[Goo89u] I. J. Good. C334. On the neutron and proton masses: a numerological case study. Journal of Statistical Computation and Simulation, 32(1–2):103–111, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90].

Good:1989:CSN

[Goo89v] I. J. Good. C335. Some numerology for the mass ratio of proton to electron, and a partially Bayesian evaluation. Journal of Statistical Computation and Simulation, 32(1–2):111–115, 1989. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1989:CFD

[Goo89w] I. J. Good. Corrigenda: “The fractional dimensional theory of continued fractions” [Proc. Cambridge Philos. Soc. 37 (1941), 199–228; MR 3, 75]. Mathematical Proceedings of the Cambridge Philosophical Society, 105(3):607, 1989. CODEN MPCPCO. ISSN 0305-0041,1469-8064.

Good:1989:FER

[Goo89x] I. J. Good. F20. An enveloping regression procedure. Journal of Statistical Planning and Inference, 21(1):134, January 1989. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/037837588990030X.

Good:1989:FCV

[Goo89y] I. J. Good. F21. Chaotics vs. chaology. Journal of Statistical Planning and Inference, 21 (1):134–135, January 1989. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0378375889900311.

Good:1989:IBS

[Goo89z] I. J. Good. The interface between statistics and the philosophy of science. In Logic, methodology and philosophy of science, VIII (Moscow, 1987), volume 126 of Stud. Logic Found. Math., pages 393–411. North-Holland, Amsterdam, The Netherlands, 1989. ISBN 0-444-70520-1.

Good:1990:ASC

[Goo90a] I. J. Good. Abstract of “Speculations concerning the future of statistics”. The American Statistician, 44(2):132–133, May 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL https://www.
jstor.org/stable/2684151.

Good:1990:BRF

[Goo90b] I. J. Good. Book review: Fractals Everywhere (Michael Barnsley). SIAM Review, 32(2): 317–318, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1990:CVT

[Goo90c] I. J. Good. C342.5. The p-value trap, an editorial comment on C342. Journal of Statistical Computation and Simulation, 34(2– 3):162, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [Joh90].

Good:1990:CUD

[Goo90d] I. J. Good. C344. The use of divergent series in statistical problems. Journal of Statistical Computation and Simulation, 34(2– 3):164, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CRR

[Goo90e] I. J. Good. C347. Randomresponse questionnaires and liedetectors. Journal of Statistical Computation and Simulation, 34(2–3):166–168, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CMT

[Goo90f] I. J. Good. C356. Multivariate tests after dimensionality reduction. Journal of Statistical Computation and Simulation, 36 (1):45, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CI

[Goo90g] I. J. Good. C357. Isolikelihoods. Journal of Statistical Computation and Simulation, 36(1):45– 47, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CFD

[Goo90h] I. J. Good. C359. Factorial designs and the programming of games. Journal of Statistical Computation and Simulation, 36(1):49–50, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CEC

[Goo90i] I. J. Good. C361. Editorial concerning Terrell’s conjecture. Journal of Statistical Computation and Simulation, 36(1):53– 54, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CBF

[Goo90j] I. J. Good. C363. A Besselfunction inequality derived by a statistical argument. Journal of Statistical Computation and Simulation, 36(1):56–57, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CMV

[Goo90k] I. J. Good. C367. Medialized p-values, an earlier reference. Journal of Statistical Computation and Simulation, 36(2– 3):177, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CEDa

[Goo90l] I. J. Good. C368. On the exact distribution of Pearson’s χ^{2} for the lady tasting beer. Journal of Statistical Computation and Simulation, 36(2–3):177–179, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CID

[Goo90m] I. J. Good. C369. Invariant distance in statistics: some miscellaneous comments. Journal of Statistical Computation and Simulation, 36(2–3):179–186, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CCB

[Goo90n] I. J. Good. C370. A compromise between credibility and subjective probability. Journal of Statistical Computation and Simulation, 36(2–3):186–193, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CFY

[Goo90o] I. J. Good. C372. Fifty years of errors. Journal of Statistical Computation and Simulation, 36(2–3):193–194, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CEDb

[Goo90p] I. J. Good. C377. Exploratory data analysis. Journal of Statistical Computation and Simulation, 37(3–4):243–245, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:CCC

[Goo90q] I. J. Good. C378. Comments concerning the hadron quantal hypothesis. Journal of Statistical Computation and Simulation, 37(3–4):245–247, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1990:FEM

[Goo90r] I. J. Good. Fractals everywhere (Michael Barnsley). SIAM Review, 32(2):317–??, June 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1990:QHH

[Goo90s] I. J. Good. A quantal hypothesis for hadrons and the judging of physical numerology. In Disorder in physical systems, Oxford Sci. Publ., pages 129–165. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853215-6.

Good:1990:RCG

[Goo90t] I. J. Good. A reply to the comment by G. A. Barnard. Journal of Statistical Planning and Inference, 25(3):439–440, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/037837589090087B.

See [Bar90].

Good:1990:SCF

[Goo90u] I. J. Good. Speculations concerning the future of statistics. Journal of Statistical Planning and Inference, 25(3):441–466, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/037837589090088C.

Good:1990:SDF

[Goo90v] I. J. Good. Statistical discussion forum. Journal of Statistical Planning and Inference, 26(1):109–125, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0378375890901009.

See comment [Jos96a].

Good:1990:SFP

[Goo90w] I. J. Good. A suspicious feature of the Popper/Miller argument. Philosophy of Science, 57(3):535–536, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

                                        tronic).  URL https://www.
                                        jstor.org/stable/187644.

Good:1991:BRS

[Goo91a] I. J. Good. Book review: The Science of Fractal Images (Heinz-Otto Peitgen and Dietmar Saupe, eds.). SIAM Review, 33(3): 496–499, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1991:CCC

[Goo91b] I. J. Good. C383. A comment concerning optional stopping. Journal of Statistical Computation and Simulation, 39(3): 191–192, 1991. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1991:RBS

[Goo91c] I. J. Good. Review: The Science of Fractal Images. by Heinz-Otto Peitgen, Dietmar Saupe, M. F. Barnsley, R. L. Devaney, B. B. Mandelbrot, R. F. Voss, Y. Fisher, M. McGuire. SIAM Review, 33(3):496–499, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.

jstor.org/stable/2031463.

Good:1991:SUD

[Goo91d] I. J. Good. Serial unblocked designs balanced for correlation between neighbours, II. Communications in Statistics: Theory and Methods, 20(4):1225–1231, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1992:BNB

[Goo92a] I. J. Good. The Bayes/non-Bayes compromise: a brief review. Journal of the American Statistical Association, 87(419): 597–606, September 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2290192.

Good:1992:CRR

[Goo92b] I. J. Good. C386. Ridge regression and the harmonic-mean rule of thumb. Journal of Statistical Computation and Simulation, 40(3–4):262, 1992. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CFC

[Goo92c] I. J. Good. C387. Further comments concerning the lady tasting tea or beer: p-values and restricted randomization. Journal of Statistical Computation and Simulation, 40(3–4):263– 267, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CFN

[Goo92d] I. J. Good. C388. A further note on physical numerology. Journal of Statistical Computation and Simulation, 40(3–4): 267–268, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CPCa

[Goo92e] I. J. Good. C389. Partial correlation and spherical trigonometry, II. Journal of Statistical Computation and Simulation, 43(3– 4):253, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:CPCb

[Goo92f] I. J. Good. C390. Partial correlation coefficients are analogous to partial differential coefficients. Journal of Statistical Computation and Simulation, 43(3–4): 254–256, 1992. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1992:IRA

[Goo92g] I. J. Good. Introductory remarks for the article in Biometrika 66 (1979), “A. M. Turing’s Statistical Work in World War II”. In Britton [Bri92], pages 211–223. ISBN 0-444-88059-3. LCCN QA3 .T87 1992. Edited and with an introduction and postscript by J. L. Britton and Irving John Good. With a preface by P. N. Furbank.

Good:1993:CTS

[Goo93a] I. J. Good. C394. Tendencies to be sufficient or necessary causes. Journal of Statistical Computation and Simulation, 44(3–4): 261–262, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CRR

[Goo93b] I. J. Good. C397. Refutation and rejection versus inexactification, and other comments concerning terminology. Journal of Statistical Computation and Simulation, 47(1–2):91–92, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CVL

[Goo93c] I. J. Good. C398. P -values, luck, and surprise. Journal of Statistical Computation and Simulation, 47(1–2):92–94, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CTF

[Goo93d] I. J. Good. C399. Trigonometrical factorization of the Fibonacci and Lucas numbers, and sums of powers of cotangents. Journal of Statistical Computation and Simulation, 47(1–2):94–99, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CTM

[Goo93e] I. J. Good. C400. A tentative measure of probabilistic causation relevant to the philosophy of the law. Journal of Statistical Computation and Simulation, 47(1–2):99–105, 1993. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CFL

[Goo93f] I. J. Good. Complex Fibonacci and Lucas numbers, continued fractions, and the square root of the golden ratio. Fibonacci Quarterly, 31(1):7–20, February 1993. CODEN FIBQAU. ISSN 0015-0517,2641-340X. URL http:/

                                        /www.fq.math.ca/Scanned/31-

1/good.pdf. See erratum [Goo93g].

Good:1993:ECF

[Goo93g] I. J. Good. Erratum: “Complex Fibonacci and Lucas numbers, continued fractions, and the square root of the golden ratio” [Fibonacci Quart. 31 (1993), no. 1, 7–20; MR1202337 (93k:11010)]. Fibonacci Quarterly, 31(3):274–??, August 1993. CODEN FIBQAU. ISSN 0015-0517,2641-340X. URL http:/

                                        /www.fq.math.ca/Scanned/31-

3/good.pdf. See [Goo93f].

Gook:1993:EF

[Goo93h] I. J. Gook. Enigma and Fish. In Hinsley and Stripp [HS93], pages 149–166. ISBN 0-19-820327-6, 0-19-285304-X (paperback). LCCN D810.C88 M46 1993.

URL https://archive.org/
details/codebreakersinsi0000unse_
r2h6.

Good:1994:CCE

[Goo94a] I. J. Good. C407. A comment on the evaluation of numerology. Journal of Statistical Computation and Simulation, 49(3–4): 216–217, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFA

[Goo94b] I. J. Good. C415. A fourth argument for the explication of causal tendency. Journal of Statistical Computation and Simulation, 49(3–4):235–237, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFL

[Goo94c] I. J. Good. C416. Fibonacci and Lucas numbers and trigonometry. Journal of Statistical Computation and Simulation, 49(3– 4):237, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFE

[Goo94d] I. J. Good. C417. The flybutter effect: the obverse of the butterfly effect. Journal of Statistical Computation and Simulation, 49(3–4): 237–238, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CFP

[Goo94e] I. J. Good. C418. Fuzzy probability and hierarchical Bayes.

Journal of Statistical Computation and Simulation, 49(3–4): 238–239, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CKT

[Goo94f] I. J. Good. C419. A key to the topic-independent philosophy of science. Journal of Statistical Computation and Simulation, 49(3–4):239–240, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CES

[Goo94g] I. J. Good. C420. The existence of sharp null hypotheses. Journal of Statistical Computation and Simulation, 49(3–4): 241–242, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CTL

[Goo94h] I. J. Good. C421. Turing’s little theorem is not really paradoxical. Journal of Statistical Computation and Simulation, 49(3–4): 242–244, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CCW

[Goo94i] I. J. Good. C423. A comment on the Wilcoxon sign-rank test and a conjecture concerning hypothyroidism. Journal of Statistical Computation and Simulation, 49(3–4):249–250, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CST

[Goo94j] I. J. Good. C424. Significance tests: are they more important in science than in technology? Journal of Statistical Computation and Simulation, 49(3–4): 250–251, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CRV

[Goo94k] I. J. Good. C426. Rationality versus minimax. Journal of Statistical Computation and Simulation, 49(3–4):253, 1994. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:CTN

[Goo94l] I. J. Good. Causal tendency, necessitivity and sufficientivity: an updated review. In Humphreys [Hum94], pages 293–315. ISBN 0-7923-2552-4 (vol. 1), 0-7923-2553-2 (vol. 2), 0-7923-2862-0 (vol. 3), 0-7923-2554-0 (set). LCCN Q175 .P3425 1994. URL https:/

/link.springer.com/book/10.
1007/978-94-011-0774-7.

Good:1994:CDH

[Goo94m] I. J. Good. Contribution to the discussion of Henry E. Kyburg, Jr.’s “Believing on the basis of the evidence”. Computational Intelligence, 10(1):57–58, February 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Good:1994:SCK

[Goo94n] I. J. Good. The self-consistency of the kinematics of special relativity. Physics Essays, 7(4):436–441, 1994. CODEN PHESEM. ISSN 0836-1398.

Good:1994:SPA

[Goo94o] I. J. Good. A symmetry property of alternating sums of products of reciprocals. Fibonacci Quarterly, 32(3):284–287, June 1994. CODEN FIBQAU. ISSN 0015-0517,2641-340X. URL http:/

                                        /www.fq.math.ca/Scanned/32-
                                        3/good.pdf.

Good:1995:CCC

[Goo95a] I. J. Good. C435. A comment concerning physical numerology. Journal of Statistical Computation and Simulation, 52(2):190– 191, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CAR

[Goo95b] I. J. Good. C436. An apology related to blurred hypotheses. Journal of Statistical Computation and Simulation, 52(2): 191–192, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CHR

[Goo95c] I. J. Good. C437. Haphazardness is not randomness, and the defeat of Hitler. Journal of Statistical Computation and Simulation, 52 (2):192, 1995. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CRA

[Goo95d] I. J. Good. C438. Reliability always depends on probability of course. Journal of Statistical Computation and Simulation, 52(2):192–193, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:CDD

[Goo95e] I. J. Good. C439. Dual density functions. Journal of Statistical Computation and Simulation, 52(2):193–194, 1995. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1995:LMR

[Goo95f] I. J. Good. Lorentz matrices:

a review. International Journal of Theoretical Physics, 34(5):779– 799, 1995. CODEN IJTPBM. ISSN 0020-7748,1572-9575.

Good:1995:MPB

[Goo95g] I. J. Good. The mathematics of philosophy: A brief review of my work. In Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I [JL95], chapter 13, pages 211–238. ISBN 94-010-4211-X, 94-011-0471-9 (e-book). ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN ????

URL https://link.springer.
com/chapter/10.1007/978-94-
011-0471-5_13.

Good:1995:RRS

[Goo95h] I. J. Good. Review of Richard Stone, The Assumptions on Which Causal Inferences Rest. Mathematical Reviews, 95F(??): 3546–3547, June 1995. CODEN MAREAR. ISSN 0025-5629. See [Sto93].

Good:1996:CCI

[Goo96a] I. J. Good. C443. Can the idea of the QH test for normality be used for testing the Weibull distribution? Journal of Statistical Computation and Simulation, 55(3): 257–263, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1996:CRC

[Goo96b] I. J. Good. C445. A reply to C444. Journal of Statistical Computation and Simulation, 55(3): 265, 1996. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [BH96].

Good:1996:FQA

[Goo96c] I. J. Good. F44A. A question about F44. Journal of Statistical Planning and Inference, 55 (2):258, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0378375896900671.

See [Jos96a] and reply [Jos96b].

Good:1996:FVR

[Goo96d] I. J. Good. F46. Varieties of the regression fallacy. Journal of Statistical Planning and Inference, 55(2):260, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0378375896888495.

Good:1996:FCF

[Goo96e] I. J. Good. F49A. A comment on F49. Journal of Statistical Planning and Inference, 55(2): 264, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0378375896900695.

Good:1997:RVM

[Goo97a] I. J. Good. The roughness of visitations in a Markov chain, a review with extensions. In Advances in the theory and practice of statistics, Wiley Ser. Probab. Statist. Appl. Probab. Statist., pages 67–87. Wiley, New York, NY, USA, 1997. ISBN 0-471-15574-8.

Good:1997:SCKa

[Goo97b] I. J. Good. The self-consistency of the kinematics of special relativity. II. Physics Essays, 10(2):327– 333, 1997. CODEN PHESEM. ISSN 0836-1398.

Good:1997:SCKb

[Goo97c] I. J. Good. The self-consistency of the kinematics of special relativity. III. Physics Essays, 10 (3):454–465, 1997. CODEN PH-ESEM. ISSN 0836-1398.

Good:1998:CEE

[Goo98a] I. J. Good. C446. The excess of even numbers of carbon atoms in organic compounds, and Eddington’s fisherman’s fallacy: dogs, hot dogs and astronomers. Journal of Statistical Computation and Simulation, 60(2):157– 160, 1998. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1998:CCC

[Goo98b] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 61(4):287–312, 1998. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1998:SCK

[Goo98c] I. J. Good. The self-consistency of the kinematics of special relativity. IV. Physics Essays, 11 (2):248–263, 1998. CODEN PH-ESEM. ISSN 0836-1398.

Good:1999:ESC

[Goo99] I. J. Good. Erratum: “The self-consistency of the kinematics of special relativity. IV” [Phys. Essays 11 (1998), no. 2, 248– 263; MR1646418 (99g:83008)]. Physics Essays, 12(1):190, 1999. CODEN PHESEM. ISSN 0836-1398.

Good:2000:COE

[Goo00a] I. J. Good. A correction to one of Einstein’s derivations of the Lorentz transformation. Physics Essays, 13(4):580–586, 2000. CO-DEN PHESEM. ISSN 0836-1398. See erratum [Goo00a].

Good:2000:TAE

[Goo00b] I. J. Good. Turing’s anticipation of empirical Bayes in connection with the cryptanalysis of the Naval Enigma. Journal of Statistical Computation and Simulation, 66(2):101–111, 2000. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. 50th Anniversary of the Department of Statistics, Virginia Tech, Part II (Blacksburg, VA, 1999).

Good:2001:ECO

[Goo01] I. J. Good. Erratum: “A correction to one of Einstein’s derivations of the Lorentz transformation” [Phys. Essays 13, (2000), no. 4, 580–586; MR1867077 (2003a:83004)]. Physics Essays, 14(2):180, 2001. CODEN PH-ESEM. ISSN 0836-1398. See [Goo00a].

Good:2002:CCC

[Goo02] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 72(1):1–31, 2002. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2003:CSI

[Goo03a] I. J. Good. C473. Straddling the interquantile range. Journal of Statistical Computation and Simulation, 73(8):612–616, 2003. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2003:SCKa

[Goo03b] I. J. Good. The self-consistency of the kinematics of special relativity. V(A). Physics Essays, 16(1): 29–42, 2003. CODEN PHESEM. ISSN 0836-1398.

Good:2003:SCKb

[Goo03c] I. J. Good. The self-consistency of the kinematics of special relativity. V(B). Physics Essays, 16 (2):174–179, 2003. CODEN PH-ESEM. ISSN 0836-1398.

Gook:2003:AIW

[Goo03d] I. J. Gook. The accumulation of imprecise weights of evidence. In ????, editor, Proceedings of the 2003 International Symposium on Imprecise Probabilities and Their Applications (ISIPTA-03), Lugano, Switzerland, July 14–17, 2003, page ?? ????, ????, 2003.

Good:2004:CCC

[Goo04] I. J. Good. Comments, conjectures, and conclusions. Journal of Statistical Computation and Simulation, 74(12):913–922, 2004. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2006:CCCa

[Goo06a] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 76(8):737–740, 2006. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2006:CCCb

[Goo06b] I. J. Good. Comments, conjectures and conclusions. Journal of Statistical Computation and Simulation, 76(11):1027–1030, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:2009:GTF

[Goo09] Irving John Good. Good thinking: the foundations of probability and its applications. Dover, New York, NY, USA, 2009. ISBN 0-486-47438-0. xviii + 332 pp. LCCN QA273.4 .G66 2009. URL

http://www.loc.gov/catdir/
enhancements/fy1004/2009024055-
d.html;   http://www.loc.gov/
catdir/enhancements/fy1318/
2009024055-t.html.

Good:2019:SDF

[Goo19] I. J. Good. Statistical discussion forum. Journal of Statistical Planning and Inference, 15(??):129, ???? 2019. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0378375886900911.

Good:1948:BIT

[GR48] I. J. Good and G. E. H. (Gerd Edzard Harry) Reuter. Bounded integral transforms. Quarterly Journal of Mathematics. Oxford series., 19:224–234, December 1948. CODEN QJMAAT. ISSN 0033-5606,1464-3847.

Good:1978:PSS

[GR78] I. J. Good and Reinhard Razen. Problems and solutions: Solutions of advanced problems: 6137. American Mathematical Monthly, 85(10):830–831, December 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77, GB80, Goo81-31].

Greenberg:2022:BPC

[Gre22] Joel Greenberg. The Bletchley Park Codebreakers in Their Own Words. Greenhill Books, Barnsley, South Yorkshire, UK, 2022. ISBN 1-78438-811-4 (hardcover), 1-78438-812-2 (e-book), 1-78438-814-9. xvii + 334 + 16 pp. LCCN D810.S7 G74 2022. URL

                                        http://digital.casalini.it/
                                        9781784388140. With an intro-

duction by Tony Comer.

Gilles:1976:PSS

[GRM76] J. Gilles, J. Barkley Rosser, and Hugh L. Montgomery. Problems and solutions: Solutions of advanced problems: 5952. American Mathematical Monthly, 83 (10):819–820, December 1976. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gil74, GBMW75].

Good:1983:CEH

[GS83] I. J. Good and Eric P. Smith. C155. The early history of the quadratic index of diversity or repeat rate. Journal of Statistical Computation and Simulation, 17(3):231–233, 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1984:CPB

[GS84] I. J. Good and Eric P. Smith.

C184. The possible bias of the Pearson chi-squared test in nonequiprobable cases. Journal of Statistical Computation and Simulation, 19(1):79–82, 1984. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:VCG

[GS85a] I. J. Good and E. P. Smith. The variance and covariance of a generalized index of similarity especially for a generalization on an index of Hellinger and Bhattacharyya. Communications in Statistics: Theory and Methods, 14(12):3053–3061, 1985. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1985:CCP

[GS85b] I. J. Good and Eric P. Smith.

C240. A continuum of principal component methods. Journal of Statistical Computation and Simulation, 23(1–2):136–142, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See correction [Mar90].

Good:1986:AAA

[GS86a] I. J. Good and Eric P. Smith.

An additive algorithm analogous to the singular decomposition; or a comparison of polarization and multiplicative models: an example of qualitative robustness. Comm. Statist. B— Simulation Comput., 15(2):545– 569, 1986. ISSN 0361-0918.

Good:1986:CPS

[GS86b] I. J. Good and Eric P. Smith. C247. A power series for the tailarea probability of Student’s t distribution. Journal of Statistical Computation and Simulation, 23(3):248–250, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1987:CSG

[GS87] I. J. Good and Eric P. Smith. C280. Some general measures of similarity and their values under multinormality. Journal of Statistical Computation and Simulation, 28(1):75–79, 1987. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Glasser:1975:PSS

[GSA75] M. G. Glasser, B. L. R. Shawyer, and G. E. Andrews. Problems and solutions: Solutions of advanced problems: 5950. American Mathematical Monthly, 82 (4):414–415, April 1975. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ehr74].

Good:1976:Mb

[GSJ76] I. J. Good, Eugene Seneta, and Donald R. Jensen. Minicommunications. Communications in Statistics: Simulation and Computation, 5(2–3):141–147, 1976. CODEN CSSCDB. ISSN 0361-0918.

Good:1976:Mc

[GSK76] I. J. Good, S. R. Searle, and A. I. Khuri. Minicommunications. Communications in Statistics: Simulation and Computation, 5(4):207–211, 1976. CO-DEN CSSCDB. ISSN 0361-0918.

Good:1956:NNS

[GT56] I. J. Good and G. H. Toulmin. The number of new species, and the increase in population coverage, when a sample is increased. Biometrika, 43 (1/2):45–63, June 1956. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2333577.

Good:1968:PSS

[GT68] I. J. Good and T. Teichmann.

Problems and solutions: Solutions of elementary problems: E1923. American Mathematical Monthly, 75(2):194–195, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Good:1976:ICO

[GT76] I. J. Good and T. N. Tideman.

From individual to collective ordering through multidimensional attribute space. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 347(1650):371–385, 1976. CODEN PRLAAZ. ISSN 0962-8444,2053-9169.

Good:1977:SNG

[GT77] I. J. Good and T. N. Tideman. Stirling numbers and a geometric structure from voting theory. Journal of Combinatorial Theory (Series A), 23(1):34–45, 1977. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Good:1978:IST

[GT78] I. J. Good and T. N. Tideman. Integration over a simplex, truncated cubes, and Eulerian numbers. Numerische Mathematik, 30 (4):355–367, December 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Good:1981:CRI

[GT81] I. J. Good and T. N. Tideman. C93. The relevance of imaginary alternatives. Journal of Statistical Computation and Simulation, 12(3–4):313–315, 1981. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1985:CJE

[GT85] I. J. Good and Gordon Tullock. C221. On judicial errors. Journal of Statistical Computation and Simulation, 21(1):75– 76, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hald:1960:CHD

[Hal60] A. Hald. The compound hypergeometric distribution and a system of single sampling inspection plans based on prior distributions and costs. Technometrics, 2 (3):275–340, August 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL https://www.jstor.org/
stable/1266247. See discussion

[BLH^{+}60].

Harper:1975:CJG

[Har75]  William L. Harper.  Com-

ments on I. J. Good. Synthese, 30(1–2):75–78, February/ March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http:/

/link.springer.com/article/
10.1007/BF00485295.  See

[Goo75d].

Heyde:1986:SSA

[Hey86] C. C. Heyde. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):176–177, May 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.
org/euclid.ss/1177013693.

Henderson:1968:PSS

[HG68] Mason Henderson and Michael Goldberg. Problems and solutions: Solutions of elementary problems: E1924. American Mathematical Monthly, 75(2): 195–196, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Holtzman:1985:CHU

[HG85] Golde I. Holtzman and I. J. Good. C242. Heaviside’s unit function applied to midrank procedures. Journal of Statistical Computation and Simulation, 23(1–2): 146–148, 1985. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Holtzman:1986:PCS

[HG86] Golde I. Holtzman and I. J. Good. The Poisson and chi-squared approximations as compared with the true upper-tail probability of Pearson’s X^{2} for equiprobable multinomials. Journal of Statistical Planning and Inference, 13(3):283–295, ???? 1986. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0378375886901400.

Heideman:1985:GHF

[HJB85] Michael T. Heideman, Don H. Johnson, and C. Sidney Burrus. Gauss and the history of the Fast Fourier Transform. Archive for History of Exact Sciences, 34(3): 265–277, September 1985. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=0003-9519&
volume=34&issue=3&spage=265.

Hammer:1975:APS

[HK75] F. D. Hammer and William Knight. Advanced problems and solutions: Solutions: 5955. American Mathematical Monthly, 82 (4):415–416, April 1975. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gil74].

Harrell:1987:LE

[HLT^{+}87] Frank E. Harrell, Jr., Kerry L.

Lee, James R. Thompson, Rick L. Edgeman, Lawrence M. Leemis, Nathan Mantel, Paul D. James, R. W. Farebrother, Jerome P. Keating, Robert L. Mason, Michael E. Welch, George T. Diderrich, I. J. Good, and George Casella. Letters to the Editor. The American Statistician, 41 (1):87–92, February 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2684337.

Hamdan:1971:LE

[HPG^{+}71] M. A. Hamdan, John W. Pratt, Peter Gottlieb, I. J. Good, M. A. Hamdan, S. G. Carmer, W. M. Walker, T. J. Valentine, Ralph B. D’Agostino, A. M. Kshirsagar, I. Gwyn Evans, D. G. Kabe, Edward E. Cureton, J. R. Rutherford, J. K. Sharma, and James R. Harvey. Letters to the Editor. The American Statistician, 25(5): 56–63, December 1971. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2686089.

Harris:1982:PSS

[HS82] Lawrence Harris and I. J. Schoenberg. Problems and solutions: Solutions of elementary problems: E2884. American Mathematical Monthly, 89(9):700–701, November 1982. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HSO^{+}81].

Hinsley:1993:CIS

[HS93] F. H. (Francis Harry) Hinsley and Alan Stripp, editors. Codebreakers: the Inside Story of Bletchley Park. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-820327-6, 0-19-285304-X (paperback). xxi + 321 + 8 pp. LCCN D810.C88 M46 1993.

                                        URL https://archive.org/
                                        details/codebreakersinsi0000unse_
                                        r2h6.

Harris:1981:PSE

[HSO^{+}81] Lawrence Harris, T. Sekiguchi, C. O. Oakley, Arnold Adelberg, T. F. Mori, G. J. Szekely, and I. J. Good. Problems and solutions: Elementary problems: E2884– E2889. American Mathematical Monthly, 88(5):349, May 1981. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AC84, GB82, HS82, MS82, OK84, SS84].

Humphreys:1994:PSS

[Hum94] Paul Humphreys, editor. Patrick Suppes: Scientific Philosopher, volume 233–235 of Synthese library. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-2552-4 (vol. 1), 0-7923-2553-2 (vol. 2), 0-7923-2862-0 (vol. 3), 0-7923-2554-0 (set). xix + 418 (vol. 1), viii + 314 (vol. 2), viii + 336 (vol. 3) pp. LCCN Q175 .P3425 1994. URL https:/

/link.springer.com/book/10.
1007/978-94-011-0774-7.

Israel:1994:CK

[Isr94] David Israel. Commentary on Kyburg. Computational Intelligence, 10(1):63–69, February 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Jabine:1983:LE

[JEE^{+}83] Thomas B. Jabine, Robert M.

Eissner, Marie D. Eldridge, Benjamin H. Eichhorn, Murray Aitkin, R. Kirk Steinhorst, I. J. Good, C. Philip Cox, Lynn Roy LaMotte, J. E. Mann, Stephen M. Stigler, V. M. Joshi, Ronaldo Iachan, and C. G. Khatri. Letters to the Editor. The American Statistician, 37 (1):92–97, February 1983. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2685853.

Jensen:1981:IDA

[JG81] D. R. Jensen and I. J. Good.

Invariant distributions associated with matrix laws under structural symmetry. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):

327–332, ???? 1981. CO-DEN JSTBAJ. ISSN 0035-9246.

                                        URL https://www.jstor.org/
                                        stable/2984944.

Jensen:1983:RED

[JG83] D. R. Jensen and I. J. Good. A representation for ellipsoidal distributions. SIAM Journal on Applied Mathematics, 43(5): 1194–1200, October 1983. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic).

Johnson:1989:CAA

[JK89] Norman L. Johnson and Samuel Kotz. C310. Axiomatic approaches to formulas for combining likelihoods or evidence. Journal of Statistical Computation and Simulation, 31(1):49– 54, 1989. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See errata [Ano90].

Jarvie:1995:CRM

[JL95] I. C. Jarvie and Nathaniel Laor. Critical Rationalism, Metaphysics and Science: Essays for Joseph Agassi. Volume I, volume 161 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. ISBN 94-010-4211-X, 94-011-0471-9 (ebook). ISSN 0068-0346 (print), 2214-7942 (electronic). xxii + 266 pp. LCCN ???? URL https:/

                                        /link.springer.com/book/10.
                                        1007/978-94-011-0471-5.

Johnstone:1990:CCR

[Joh90] David Johnstone. C342. Consistency and relative plausibility. Journal of Statistical Computation and Simulation, 34(2–3): 160–162, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo90c].

Joshi:1996:FFP

[Jos96a] V. M. Joshi. F44. Fallacy in the proof of Birnbaum’s theorem: the reply by Berger (F28). Journal of Statistical Planning and Inference, 55(2):257–258, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/037837589688846X.

See [Goo90v, Goo96c].

Joshi:1996:FRF

[Jos96b] V. M. Joshi. F44B. A reply to F44A. Journal of Statistical Planning and Inference, 55 (2):258–259, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0378375896888471.

See [Goo96c].

Kendall:1973:LE

[KBE^{+}73] Barbie Kendall, Irwin D. J.

Bross, Laszlo Engelman, Boyd Z. Palmer, E. Whyte, John Wiorkowski, Nathan Mantel, Marvin S. Margolis, J. Phillip Cooper, Y. S. Sathe, I. J. Good, George Zyskind, and Paul A. Johnson. Letters to the Editor. The American Statistician, 27 (1):39–44, February 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2682908.

Kimberling:1975:PSS

[KC75] C. H. Kimberling and L. Carlitz. Problems and solutions: Solutions of advanced problems: 5947. American Mathematical Monthly, 82(4):411–413, April 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ehr74].

Kedlaya:2012:EGF

[Ked12] Kiran S. Kedlaya. Errata to “Good formal structures for flat meromorphic connections, I: Surfaces,” Duke Math. J. 154 (2010), 243–418 [MR2682186]. Duke Mathematical Journal, 161(4): 733–734, 2012. CODEN DUM-JAO. ISSN 0012-7094,1547-7398.

Khuri:1977:DQF

[KG77] A. I. Khuri and I. J. Good. The distribution of quadratic forms in non-normal variables and an application to the variance ratio. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):217–221, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL https://www.jstor.
                                        org/stable/2984798.

Khuri:1989:POM

[KG89] André I. Khuri and I. J. Good.

The parameterization of orthogonal matrices: a review mainly for statisticians. South African Statist. J., 23(2):231–250, 1989. ISSN 0038-271X,1996-8450.

Khatri:1978:RNS

[Kha78]  C. G. Khatri.  A remark

on the necessary and sufficient conditions for a quadratic form to be distributed as chisquared. Biometrika, 65(1): 239–240, April 1978. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL https://www.jstor.org/
stable/2335305. See [Goo69a].

Krotki:1988:LE

[KLC^{+}88] Karol J. Krotki, Roderick J. A.

Little, Ronald Christensen, W. A. Thompson, Jr., I. J. Good, Martin Rutsch, H. N. Nagaraja, Irwin D. Bross, Fred L. Ramsey, Hanspeter Thoni, Ludo M. J. Gypen, and Joseph Lee Rodgers. Letters to the Editor. The American Statistician, 42(4):288– 291, November 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2685150.

Klamkin:1968:PSSa

[KS68] M. S. Klamkin and Seymour Schuster. Problems and solutions: Solutions of advanced problems: 5461. American Mathematical Monthly, 75(2):202–203, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Koop:1974:LE

[KsLG^{+}74] J. C. Koop, Jack Chao sheng Lee, Fernando Garagorry, Jack Mardekian, Farhad Mehran, James C. Impara, Irwin D. J. Bross, John P. Hoyt, Richard L. Haden, Don Self, Marvin A. Johnston, A. R. Sampson, David Hinkley, and I. J. Good. Letters to the Editor. The American Statistician, 28(2):73–77, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL https://www.

                                        jstor.org/stable/2683609.

Kyburg:1994:BBE

[Kyb94] Henry E. Kyburg, Jr. Believing on the basis of the evidence. Computational Intelligence, 10 (1):3–20, February 1994. CO-DEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See comments [Wel94, Sme94, Lou94, Isr94, She94, Goo94m, Doy94, BGHK94, Gol94].

Lewis:1968:PSS

[LB68] J. E. Lewis and M. G. Beumer. Problems and solutions: Solutions of elementary problems: E1916. American Mathematical Monthly, 75(2):190–191, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Lee:1995:CPM

[Lee95] J. A. N. Lee. Computer pioneers: Irving John (Jack) Good. IEEE Web site.,

1995. URL https://history.
computer.org/pioneers/hestenes.
html.

Leaverton:1972:LE

[LFB^{+}72] Paul E. Leaverton, R. J. Freund, D. E. Barton, Bruce R. Boals, Lee R. Abramson, James L. Bicksler, T. Cacoullos, I. J. Good, Masoodul Haq, James R. Harvey, Joel Huber, D. M. Olsson, Tony Lancaster, S. R. Searle, R. Choudary Hanumara, Edward F. Gocka, James R. Maar, and Richard M. Brugger. Letters to the Editor. The American Statistician, 26 (2):42–52, April 1972. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2683463.

Lindley:2010:IJG

[Lin10] Dennis V. Lindley. Irving John Good, 1916–2009. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (1):259–261, January 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/
stable/20622588.

Lindley:1969:LE

[LOG^{+}69] D. V. Lindley, Edward T.

O’Donnell, I. J. Good, M. Subramanian, R. Van Wyk, James Pachares, Max L. Eaton, James Stapleton, Bale S. Gurunanjappa, and Richard M. Brugger. Letters to the Editor. The American Statistician, 23(3):48–50, June 1969. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL https://www.

                                        jstor.org/stable/2682583.

Loui:1994:KV

[Lou94] Ronald Loui. Kyburg and Volkswagens. Computational Intelligence, 10(1):77–78, February 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Levin:1977:CML

[LR77] Bruce Levin and James Reeds. Compound multinomial likelihood functions are unimodal: proof of a conjecture of I. J. Good. Annals of Statistics, 5(1):79–87, 1977. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

                                        URL http://links.jstor.org/
                                        sici?sici=0090-5364(197701)
                                        5:1<79:CMLFAU>2.0.CO;2-W&origin=
                                        MSN.

Lefante:1986:CNO

[LS86] John J. Lefante, Jr. and Arvind K. Shah. C257. A note on the onesample t-test. Journal of Statistical Computation and Simulation, 25(3–4):295–296, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See comment [Goo86u].

Lu:1989:MMS

[LSG89] Ruey-Pyng Lu, Eric P. Smith, and I. J. Good. Multivariate measures of similarity and niche overlap. Theoret. Population Biol., 35 (1):1–21, 1989. ISSN 0040-5809 (print), 1096-0325 (electronic).

Meyer-Abich:1966:BRPb

[MA66] Adolf Meyer-Abich. Book review:

Phantasie in der Wissenschaft. Eine Anthologie unausgegorener Ideen by Irving John Good. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 50(1): 106, ???? 1966. CODEN SUARAH. ISSN 0039-4564.

URL http://www.jstor.org/
stable/20775481.

Maghsoodloo:1981:CCC

[Mag81] Saeed Maghsoodloo. C111. The continuity correction for chisquared for 2 × 2 tables. Journal of Statistical Computation and Simulation, 14(1):47–51, 1981. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marx:1990:CCC

[Mar90] Brian D. Marx. C371. A correction for C240, Good and Smith, JSCS 22, pp. 136–142. Journal of Statistical Computation and Simulation, 36(2–3):193, 1990. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163. See [GS85b].

Marsal:1975:APS

[MC75] Dietrich Marsal and Paul Chauveheid. Advanced problems and solutions: Solutions: 5957. American Mathematical Monthly, 82 (7):770, August/September 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gil74].

McCausland:1996:RSC

[McC96] I. McCausland. A reply to: “The self-consistency of the kinematics of special relativity” [Phys. Essays 7 (1994), no. 4, 436–441; MR1324249 (96c:83004)] by I. J. Good. Physics Essays, 9(3):484– 486, 1996. CODEN PHESEM. ISSN 0836-1398.

McShane:1970:RSE

[McS70] P. McShane. Randomness, Statistics and Emergence. University of Notre Dame Press, South Bend, IN, USA, 1970. x + 268 pp. LCCN ????

Meltzer:1965:TFP

[MG65] B. Meltzer and I. J. Good. Two forms of the prediction paradox. British Journal for the Philosophy of Science, XVI(61):50–51, May 1965. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http:

//bjps.oxfordjournals.org/
content/XVI/61/50.full.pdf+
html;   https://www.jstor.
org/stable/686139.

Marsaglia:1980:CGN

[MG80] George Marsaglia and I. J. Good. C69. Generating a normal sample with given sample mean and variance. Journal of Statistical Computation and Simulation, 11(1):

71–74, 1980. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Mayne:1962:SSA

[MGM62] Alan J. Mayne, Irving John Good, and John Maynard Smith, editors. The Scientist Speculates: an Anthology of Partlybaked Ideas. Heinemann, London, UK, 1962. xvii + 413 pp. LCCN ????

May:1970:PSS

[MH70] K. O. May and D. W. Hadwin. Problems and solutions: Solutions of elementary problems: E1917. American Mathematical Monthly, 77(5):527–528, May 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (ebook). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Miller:1975:AP

[Mil75a] David Miller. The accuracy of predictions. Synthese, 30 (1–2):159–191, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF00485304. See com-

ments [Ros75, Aga75b, Goo75c] and reply [Mil75b].

Miller:1975:APR

[Mil75b] David Miller. The accuracy of predictions: A reply. Synthese, 30 (1–2):207–219, February/March 1975. CODEN SYNTAE. ISSN 0039-7857(print) 1573-0964 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF00485308. See com-

ments [Ros75, Aga75b, Goo75c].

Mullin:1975:PSSb

[MM75] A. A. Mullin and William Margolis. Problems and solutions: Solutions of advanced problems: 5948. American Mathematical Monthly, 82(4):413, April 1975. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ehr74].

Montgomery:1978:PSSb

[MR78] H. L. Montgomery and Adam Riese. Problems and solutions: Solutions of advanced problems: 6136. American Mathematical Monthly, 85(9):772, November 1978. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77, MW80].

Mori:1982:PSS

[MS82] T. F. Mori and G. J. Szekely.

Problems and solutions: Solutions of elementary problems: E2888. American Mathematical Monthly, 89(9):701, November 1982. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HSO^{+}81].

McWorter:1968:PSSa

[MTD68] W. A. McWorter, E. J. Taft, and J. E. Delany. Problems and solutions: Solutions of advanced problems: 5462. American Mathematical Monthly, 75(2):203–204, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Montgomery:1980:PSSa

[MW80] H. L. Montgomery and David Wigner. Problems and solutions: Solutions of advanced problems: 6136. American Mathematical Monthly, 87(3):225–226, March 1980. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77, MR78].

Norvig:2012:AIC

[Nor12] Peter Norvig. Artificial intelligence: The coming superintelligence. New Scientist, 216(2889): viii, November 3, 2012. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). See [Goo65e].

Osteyee:1974:CBG

[OG74a] David Bridston Osteyee and Irving John Good. Comparison between I(R_{T}: S_{T}), W_{T} (H_{N}/H_{S+N}), and J_{T} (N, S + N) for Gaussian signals and noise. Lecture Notes in Mathematics, 376:96–97, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064139.

Osteyee:1974:CAA

[OG74b] David Bridston Osteyee and Irving John Good. Conclusions and areas for additional research. Lecture Notes in Mathematics, 376:130, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064144.

Osteyee:1974:D

[OG74c] David Bridston Osteyee and Irving John Good. Divergence. Lecture Notes in Mathematics, 376:44–46, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064134/.

Osteyee:1974:E

[OG74d] David Bridston Osteyee and Irving John Good. Entropy. Lecture Notes in Mathematics, 376: 13–17, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064130/.

Osteyee:1974:EMIa

[OG74e] David Bridston Osteyee and Irving John Good. Expected mutual information. Lecture Notes in Mathematics, 376: 26–38, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064132/.

Osteyee:1974:EMIb

[OG74f] David Bridston Osteyee and Irving John Good. Expected mutual information, expected weight of evidence, and divergence for random processes. Lecture Notes in Mathematics, 376: 47–53, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064135/.

Osteyee:1974:EMIc

[OG74g] David Bridston Osteyee and Irving John Good. Expected mutual information rate. Lecture Notes in Mathematics, 376: 98–111, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0064140/.

Osteyee:1974:EWE

[OG74h] David Bridston Osteyee and Irving John Good. Expected weight of evidence. Lecture Notes in Mathematics, 376: 39–43, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0064133/.

Osteyee:1974:EEW

[OG74i] David Bridston Osteyee and Irving John Good. Expressing expected weight of evidence for Gaussian generalized processes in terms of integral operators. Lecture Notes in Mathematics, 376:92–95, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0064138/.

Osteyee:1974:GPE

[OG74j] David Bridston Osteyee and Irving John Good. Gaussian processes with equal covariance functions including nonrandom signals in Gaussian noise. Lecture Notes in Mathematics, 376: 118–122, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064142/.

Osteyee:1974:IEW

[OG74k] David Bridston Osteyee and Irving John Good. Information in events and weight of evidence. Lecture Notes in Mathematics, 376:8–12, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064129/.

Osteyee:1974:IWE

[OG74l] David Bridston Osteyee and Irving John Good. Information, Weight of Evidence. The Singularity Between Probability Measures and Signal Detection, volume 376 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 0-387-06726-4, 3-540-06726-4. xi + 156 pp. LCCN QA3L4V.376.

Osteyee:1974:I

[OG74m] David Bridston Osteyee and Irving John Good. Introduction. Lecture Notes in Mathematics, 376:1–2, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064127.

Osteyee:1974:LS

[OG74n] David Bridston Osteyee and Irving John Good. List of symbols. Lecture Notes in Mathematics, 376:3–7, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0064128/.

Osteyee:1974:OEE

[OG74o] David Bridston Osteyee and Irving John Good. Other expressions for expected mutual information. Lecture Notes in Mathematics, 376:81–91, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0064137/.

Osteyee:1974:REW

[OG74p] David Bridston Osteyee and Irving John Good. Rate of expected weight of evidence. Lecture Notes in Mathematics, 376: 112–117, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0064141/.

Osteyee:1974:RBC

[OG74q] David Bridston Osteyee and Irving John Good. Relationships between certain random processes and the singularity between probability measures. Lecture Notes in Mathematics, 376: 54–80, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064136/.

Osteyee:1974:SBT

[OG74r] David Bridston Osteyee and Irving John Good. Singularity between two probability measures. Lecture Notes in Mathematics, 376:18–25, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064131/.

Osteyee:1974:SMR

[OG74s] David Bridston Osteyee and Irving John Good. Summary of the major results in this survey for Gaussian processes, including Gaussian signals and noise. Lecture Notes in Mathematics, 376:123–129, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0064143/.

Osteyee:1976:RBS

[OG76] David B. Osteyee and I. J. Good.

Regeneration of a binary signal in a uniform transmission line. IEEE Transactions on Communications, COM-2(9):1054– 1057, 1976. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Oakley:1984:PSE

[OK84] C. O. Oakley and L. Kuipers. Problems and solution: Elementary problems: E2886. American Mathematical Monthly, 91 (5):312–313, May 1984. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HSO^{+}81].

OConnor:2009:IJG

[OR09] J. J. O’Connor and E. F. Robertson. Irving John Good. Web site., April 2009. URL https://

                                        mathshistory.st-andrews.ac.
                                        uk/Biographies/Good/.

Osofsky:1978:PSSb

[Oso78] Barbara Osofsky. Problems and solutions: Solutions of advanced problems: 6134. American Mathematical Monthly, 85(9):771–772, November 1978. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ECG^{+}77].

Oyved:1926:VJA

[Oyv26] Moysheh Oyved. Visions and Jewels; an Autobiography. H. Holt and Company, New York, NY, USA, 1926. xi + 163 pp. This is an autobiography of I. J. Good’s father, Morris Edward Good, who wrote under the pseudonym Moysheh Oyved after the breakup of his marriage.

Oyved:1952:VJA

[Oyv52] Moysheh Oyved. Visions and Jewels; Autobiographic in Three Parts. Faber and Faber, London, UK, 1952. 320 pp. This is an autobiography of I. J. Good’s father, Morris Edward Good, who wrote under the pseudonym Moysheh Oyved after the breakup of his marriage.

Plante:1971:CEL

[PBB^{+}71] André Plante, G. A. Barnard, D. J. Bartholomew, M. S. Bartlett, H. Rubin, I. J. Good, O. Kempthorne, and Jock MacKay. Counter-examples and likelihood. In Foundations of statistical inference (Proc. Sympos., Univ. Waterloo, Waterloo, Ont., 1970), pages 357– 371. Holt, Rinehart and Winston of Canada, Ltd., Toronto, ON, Canada, 1971. With comments by G. A. Barnard, D. J. Bartholomew, M. S. Bartlett, I. J. Good, O. Kempthorne, J. MacKay and H. Rubin and a reply by the author.

Phillips:1974:LE

[PCM^{+}74] Haynes Phillips, P. Clifford, James H. Matis, Herbert F. Spirer, C. Terrence Ireland, John P. Hoyt, G. A. Barnard, Theodore D. Sterling, Samuel B. Lyerly, I. J. Good, Julian L. Simon, Dan Weidenfeld, V. P. Godambe, Ernest W. Grove, George A. F. Seber, Nathan Mantel, John E. Alman, T. W. Anderson, John A. Gaines, and Michael A. Golberg. Letters to the Editor.

The American Statistician, 28(4): 160–166, November 1974. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2683345.

Percival:2000:CFS

[Per00] J. N. Percival. Correcting the flaw in special relativity that leads to the twin paradox — reflections on a debate with I. J. Good and I. McCausland. Physics Essays, 13 (4):555–567, 2000. CODEN PH-ESEM. ISSN 0836-1398.

Pelz:1976:ALT

[PG76] Wolfgang Pelz and I. J. Good. Approximating the lower tailareas of the Kolmogorov–Smirnov one-sample statistic. Journal of the Royal Statistical Society. Series B (Methodological), 38(2): 152–156, ???? 1976. CO-DEN JSTBAJ. ISSN 0035-9246.

                                        URL https://www.jstor.org/
                                        stable/2985019.

Pelz:1983:CKS

[PG83] Wolfgang Pelz and I. J. Good. Calculating the Kolmogorov– Smirnov one-sample statistic. The American Statistician, 37 (2):173, May 1983. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2685888.

Pelz:1986:EPC

[PG86] Wolfgang Pelz and I. J. Good. Estimating probabilities from contingency tables when the marginal probabilities are known, by using additive objective functions. Journal of the Royal Statistical Society. Series D (The Statistician), 35(1):45–50, ???? 1986. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic).  URL https://www.
jstor.org/stable/2988297.

Petty:1968:PSS

[PJR^{+}68] C. M. Petty, W. E. Johnson, P. V. Subba Rao, B. Ramachandra Rao, and Olga Taussky. Problems and solutions: Solutions of advanced problems: 5464. American Mathematical Monthly, 75(2): 205–206, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Reese:1974:PSS

[RC74] Sylvester Reese and D. G. Cantor. Problems and solutions: Solutions of advanced problems: 5898. American Mathematical Monthly, 81(4):414–415, April 1974. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ber73].

Reeds:2012:BTC

[RDF12] James A. Reeds, Whitfield Diffie, and Judith Veronica Field, editors. Breaking Teleprinter Ciphers at Bletchley Park: an Edition of I. J. Good, D. Michie, and G. Timms, General Report on Tunny with Emphasis on Statistical Methods (1945). Wiley, New York, NY, USA, 2012. ISBN 0-470-46589-1 (hardcover). cxi + 673 pp. LCCN D810.C88 G66 2015.

Rees:1946:NPJ

[Ree46] D. Rees. Note on a paper by I. J. Good. Journal of the London Mathematical Society, 21: 169–172, 1946. CODEN JLM-SAK. ISSN 0024-6107,1469-7750.

Rejewski:1981:HPM

[Rej81] Marian Rejewski. How Polish mathematicians deciphered the Enigma. Annals of the History of Computing, 3(3): 213–234, July/September 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1981/pdf/a3213.pdf;   http:
//www.computer.org/annals/
an1981/a3213abs.htm. After-

words by Cipher A. Deavours and I. J. Good. This article was entitled “Jak matematycy polscy rozszyfrowz Enigmgȩ” in the Annals of the Polish Mathematical Society, Series II, Wiadomości Matematyczne, Volume 23, 1980, 1-28, translated by Joan Stepenske. See minor correction [Ano81].

Rao:1968:PSSa

[RF68] D. R. Rao and William Forman. Problems and solutions: Solutions of elementary problems: E1919. American Mathematical Monthly, 75(2):191–192, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Rubin:1969:QAd

[RG69] Ernest Rubin and I. J. Good.

Questions and answers. The American Statistician, 23(4): 42–45, October 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2681742.

Russell:1995:AIM

[RN95] Stuart J. (Stuart Jonathan) Russell and Peter Norvig. Artificial Intelligence: a Modern Approach. Prentice Hall series in artificial intelligence. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-103805-2, 0-13-360124-2 (paperback). xxviii + 932 pp. LCCN Q335 .R86 1995. Contributing writers: John F. Canny, Jitendra M. Malik, and Douglas D. Edwards.

Russell:1998:AIM

[RN98] Stuart J. Russell and Peter Norvig. Artificial Intelligence: a Modern Approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-790395-2 (hardcover), 0-13-790395-2 (paperback). xxviii + 1081 pp. LCCN ????

Russell:2010:AIM

[RN10] Stuart J. Russell and Peter Norvig. Artificial Intelligence: a Modern Approach. Prentice Hall series in artificial intelligence. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2010. ISBN 0-13-604259-7. xviii + 1132 pp. LCCN Q335 .R86 2010.

Russell:2021:AIM

[RNC^{+}21] Stuart Jonathan Russell, Peter Norvig, Ming-Wei Chang, Jacob Devlin, and Anca Dragan. Artificial Intelligence: a Modern Approach. Pearson series in artificial intelligence. Pearson, Hoboken, NJ, USA, fourth edition, 2021. ISBN 0-13-461099-7, 0-13-695907-5 (e-book), 0-13-467193-7, 0-13-604259-7. xvii + 1115 pp. LCCN Q335 .R86 2021.

Rosenkrantz:1975:TCD

[Ros75] Roger D. Rosenkrantz. Truthlikeness: Comment on David Miller. Synthese, 30(1–2): 193–197, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

link.springer.com/article/
10.1007/BF00485305.  See

[Mil75a] and reply [Mil75b].

Spohn:1968:PSS

[SAG68] W. G. Spohn, Jr., A. S. Adikesavan, and H. W. Gould. Problems and solutions: Solutions of advanced problems: 5463. American Mathematical Monthly, 75(2): 204–205, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Savage:1954:TFS

[Sav54] Leonard J. Savage. The Foundations of Statistics. Wiley, New York, NY, USA, 1954. 294 pp. LCCN HA29 .S28 1954.

Savage:1962:FSIb

[Sav62] Leonard Jimmie Savage, editor.

The Foundations of Statistical Inference; a Discussion Opened by L. J. Savage at a Meeting of the Joint Statistics Seminar, Birkbeck and Imperial Colleges, in the University of London. Methuen’s monographs on applied probability and statistics. Methuen, London, UK, 1962. ISBN 0-416-60810-8. 112 pp. LCCN HA29 .J59 1959.

Savage:1972:FS

[Sav72] Leonard J Savage. The Foundations of Statistics. Dover, New York, NY, USA, second revised edition, 1972. ISBN 0-486-62340-1, 0-486-62349-1. xv + 310 pp. LCCN HA29 .S28

1972. URL http://catdir.
loc.gov/catdir/description/
dover032/79188245.html;   http:
//catdir.loc.gov/catdir/enhancements/
fy1318/79188245-t.html;   http:
//library.mpib-berlin.mpg.
de/toc/ze\_2006\_942.pdf;
http://swbplus.bsz-bw.de/bsz011185627inh.
htm.

Schurle:1974:PSS

[SB74] A. W. Schurle and D. W. Brown.

Problems and solutions: Solutions of advanced problems: 5896. American Mathematical Monthly, 81(4):414, April 1974. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ber73].

Savage:1962:FSIa

[SBC^{+}62] L. J. Savage, George Barnard, Jerome Cornfield, Irwin Bross, George E. P. Box, I. J. Good, D. V. Lindley, C. W. Clunies-Ross, John W. Pratt, Howard Levene, Thomas Goldman, A. P. Dempster, Oscar Kempthorne, and Allan Birnbaum. On the foundations of statistical inference: Discussion. Journal of the American Statistical Association, 57(298):307–326, June 1962. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL https://www.jstor.org/
stable/2281641. See [Bir62].

Schaffner:1974:PPB

[SC74] Kenneth F. Schaffner and Robert S. Cohen, editors. PSA 1972: proceedings of the 1972 Biennial Meeting Philosophy of Science Association (Olds Plaza Hotel, East Lansing, Michigan, October 27–29, 1972), volume 20 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0408-2, 90-277-0409-0, 94-010-2140-6 (e-book). ISSN 0068-0346. LCCN Q175 .B7312. URL https:/

/link.springer.com/book/10.
1007/978-94-010-2140-1.

Subbarao:1968:PSSa

[SG68] M. V. Subbarao and M. G. Greening. Problems and solutions: Solutions of advanced problems: 5467. American Mathematical Monthly, 75(2):207, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Spencer:1974:APS

[SH74] Joel Spencer and M. J. Hoffman. Advanced problems and solutions: Solutions: 5899. American Mathematical Monthly, 81 (4):415–416, April 1974. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Ber73].

Shenoy:1994:DKB

[She94] Prakash P. Shenoy. Discussion of Kyburg’s “Believing on the basis of the evidence”. Computational Intelligence, 10(1):92–93, February 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic).

Shiu:2025:IBP

[Shi25] Daniel Shiu. The influence of Bletchley Park on UK mathematics. Cryptologia, 49(4):372– 387, 2025. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http:

//www.tandfonline.com/doi/
full/10.1080/01611194.2025.
2457086.

Singpurwalla:1986:SSA

[Sin86] Nozer D. Singpurwalla. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):177–178, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013694. See [Goo86y].

Singpurwalla:2008:JGZ

[Sin08] N. Singpurwalla. Jack: The good Zoroastrian. In Good et al. [GBS08], page ?? ISBN 0-89263-006-X (hardcover), 0-89263-007-8 (perfect). LCCN QA99 .G66 2008.

Srirama:1971:LE

[SKR^{+}71] M. Srirama, Scott A. Krane, B. E. Rodden, Robert W. Kennard, Robert A. Agnew, Edward Pollak, I. J. Good, J. A. Hartigan, Edward E. Cureton, D. C. Crocker, C. Philip Cox, E. Ann Blair, David M. Brown, and Al Wilson. Letters to the Editor. The American Statistician, 25(3):52–56, June 1971. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

                                        URL https://www.jstor.org/
                                        stable/2683328.

Smets:1994:BBE

[Sme94] Phillipe Smets. Believing on the basis of evidence: Answer to Kyburg’s paper. Computational Intelligence, 10(1):94–98, February 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Smith:2019:WCS

[Smi19] Michael Smith. WWII Codebreakers and Spies. Andre Deutsch, London, UK, 2019. ISBN 0-233-00602-8 (hardcover). 160 pp. LCCN D810.S7.

Solomon:1986:SSA

[Sol86] Herbert Solomon. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):174–176, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.
ss/1177013692.

Segiguchi:1984:PSS

[SS84] T. Segiguchi and D. Hammond Smith. Problems and solutions: Solutions of elementary problems: E2885. American Mathematical Monthly, 91(4):257–258, April 1984. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HSO^{+}81].

Scott:2008:IJG

[ST08] D. Scott and J. R. Thompson.

The impact of I. J. Good on density estimation. In Good et al. [GBS08], page ?? ISBN 0-89263-006-X (hardcover), 0-89263-007-8 (perfect). LCCN QA99 .G66 2008.

Stone:1993:AWC

[Sto93] Richard Stone. The assumptions on which causal inferences rest. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):455–466, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL https://www.jstor.
org/stable/2346206. See review

[Goo95h].

Sloane:1993:CESa

[SW93] N. J. A. (Neil James Alexander) Sloane and A. D. (Aaron D.) Wyner, editors. Claude Elwood Shannon: Collected Papers.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-470-54424-4, 0-7803-0434-9. xliv + 924 pp. LCCN TK5101 .S448 1993. With a profile of Shannon by Anthony Liversidge (pp. xix–xxxiii).

Salzer:1981:RHH

[SWT^{+}81] Herbert E. Salzer, Eric A. Weiss, Henry S. Tropp, Jane Smith, and Robert W. Rector. Reviews: H. H. Goldstine: A History of Numerical Analysis; Electronics: An Age of Innovation; J. A. N. Lee: Banquet Anecdotes and Conference Excerpts; R. L. Wexelblat: History of Programming Languages: Capsule reviews. Annals of the History of Computing, 3(3):289–302, July/September 1981. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1981/pdf/a3289.pdf;   http:
//www.computer.org/annals/
an1981/a3289abs.htm. See mi-

nor correction [Ano81].

Tan:1966:PSE

[TLM^{+}66] Kaidy Tan, J. E. Lewis, K. O. May, L. J. Warren, Jerry Tice, D. R. Rao, H. S. M. Coxeter, Gregory Wulczyn, I. J. Good, and Mason Henderson. Problems and solutions: Elementary problems: E1915–E1924. American Mathematical Monthly, 73 (8):890–892, October 1966. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CB68, CGGR68, GT68, HG68, LB68, MH70, RF68, TQ68, WTR68, WV68].

Tan:1968:PSS

[TQ68] Kaidy Tan and J. M. Quoniam.

Problems and solutions: Solutions of elementary problems: E1915. American Mathematical Monthly, 75(2):190, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

vanderVat:2009:OJG

[vdV09] Dan van der Vat. Obituary: Jack Good: Bletchley codebreaker, he worked closely with Alan Turing. The Guardian, ??(??):??, April 28, 2009. ISSN 0261-3077 (print), 1756-3224 (elec-

tronic).  URL https://www.
theguardian.com/science/2009/
apr/29/jack-good-codebreaker-
obituary.

Valentine:1969:LE

[VKT^{+}69] Thomas J. Valentine, Salem H.

Khamis, C. J. Thomas, Francis Emeric Binet, R. C. Juola, John H. Smith, I. J. Good, W. Edwards Deming, James A. Merrill, Samuel H. Plotkin, A. L. Wilson, and Jack Tomsky. Letters to the Editor. The American Statistician, 23 (4):51–55, October 1969. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL https://www.jstor.org/
stable/2681747.

Volk:1968:PSSa

[VM68] Benjamin Volk and M. E. Muldoon. Problems and solutions: Solutions of advanced problems: 5466. American Mathematical Monthly, 75(2):206, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Wellman:1994:KAR

[Wel94] Michael P. Wellman. Kyburgian acceptance: a rejection, hedged. Computational Intelligence, 10 (1):103–106, February 1994. CO-DEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). See [Kyb94].

Wetherill:1960:SRB

[Wet60]  G. B. Wetherill.  Some re-

marks on the Bayesian solution of the single sample inspection scheme. Technometrics, 2(3): 341–352, August 1960. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL https://www.jstor.org/
                                        stable/1266248. See discussion

[BLH^{+}60].

Wettersten:1975:GCC

[Wet75] John R. Wettersten. Good’s compromise: Comments on I. J. Good. Synthese, 30(1–2): 79–82, February/March 1975. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-0964 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/BF00485296.  See

[Goo75d].

Williams:1968:APS

[WF68] S. W. Williams and S. P.

Franklin. Advanced problems and solutions: Solutions: 5468. American Mathematical Monthly, 75(2):208, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Wilkes:1982:MRJ

[WG82] Maurice V. Wilkes and I. J.

Good. Meetings in retrospect: J. G. Brainerd on the ENIAC; A Report on T. H. Flowers’s Lecture on Colossus. Annals of the History of Computing, 4(1):53–59, January/ March 1982. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/
an1982/pdf/a1053.pdf;   http:
//www.computer.org/annals/
an1982/a1053abs.htm.

Williams:2013:IJJ

[Wil13] M. R. Williams. Irving John (Jack) Good. Web site, April 19,

2013. URL https://history.
computer.org/pioneers/good.
html;   https://history.computer.
org/pioneers/pdfs/G/Good.pdf.

2013.

Wadycki:1973:LE

[WSG^{+}73] Walter J. Wadycki, B. K.

Shah, P. D. Ghangurde, Edward J. Dudewicz, Nathan Mantel, Charles C. Brown, Harold J. Larson, Donald R. Barr, James W. Frane, Bernard Saperstein, I. J. Good, and Howard L. Jones. Letters to the Editor. The American Statistician, 27(3):123–127, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL https://www.
jstor.org/stable/2683639.

Wells:1990:LE

[WTA^{+}90] Martin T. Wells, Ram C. Tiwari, I. Arizono, H. Ohta, James W. Mergerson, Gabriella M. Belli, Christopher Cox, Murray A. Jorgensen, Jacques Benichou, Mitchell H. Gail, Warren F. Kuhfeld, Brian Dawkins, Walter B. Studdiford, Colin Goodall, W. D. Kaigh, Stephen W. Looney, Robert Kinnison, James A. Gibbons, Joel R. Levin, Ronald C. Serlin, K. J. A. Revfeim, Charles R. McConnell, Robert M. Norton, R. W. Farebrother, I. J. Good, Stanley Lebergott, and Vedula N. Murty. Letters to the editor. The American Statistician, 44(1):56–65, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic).  URL https://www.
jstor.org/stable/2684963.

Warren:1968:PSS

[WTR68] L. J. Warren, Jerry Tice, and Stanley Rabinowitz. Problems and solutions: Solutions of elementary problems: E1918. American Mathematical Monthly, 75 (2):191, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Wulczyn:1968:PSSa

[WV68] Gregory Wulczyn and A. M.

Vaidya. Problems and solutions: Solutions of elementary problems: E1922. American Mathematical Monthly, 75(2):193–194, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TLM^{+}66].

Yudkowsky:2013:IEM

[Yud13] Eliezer Yudkowsky. Intelligence explosion microeconomics. Technical Report 2013-1, Machine Intelligence Research Institute, Berkeley, CA, USA, September 2013. 96 pp. See [Goo65e].

Zabell:2023:SLJ

[Zab23] Sandy Zabell. The secret life of I. J. Good. Statistical Science, 38 (2):285–302, 2023. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Zellner:1980:BAE

[Zel80] Arnold Zellner, editor. Bayesian Analysis in Econometrics and Statistics: Essays in Honor of Harold Jeffreys. North-Holland, Amsterdam, The Netherlands, 1980. ISBN 0-444-85270-0. x + 474 pp. LCCN HB139 .B39.